### v2

# Quantum super-spherical pairs

Abstract
We introduce quantum super-spherical pairs as coideal subalgebras in general linear and orthosymplectic quantum supergroups. These subalgebras play a role of isotropy subgroups for matrices solving $\mathbb{Z}_{2}$-graded reflection equation. They generalize quantum (pseudo)-symmetric pairs of Letzter-Kolb-Regelskis-Vlaar.
2010 AMS Subject Classification: 17B37,17A70.
Key words: Quantum super-spherical pairs, Quantum symmetric pairs, Quantum supergroups, Graded reflection equation
1 Introduction
This paper is devoted to a $\mathbb{Z}_{2}$-graded generalization of quantum symmetric pairs pioneered by Letzter [1] and developed to a deep theory of quantum symmetric spaces, [2, 3, 4]. The classical supergeometry is an established field of mathematics [5], [6] motivated by profound applications of supersymmetry in quantum physics [7, 8]. That concerns the concept of symmetric and, more generally, spherical spaces [9, 10]. Quantum symmetric pairs stem from Reflection Equation, which plays for them a similar fundamental role as the Yang-Baxter Equation for quantum groups. Quantum supergroups have been in research focus from the very birth of quantum groups and are well understood [11]. It seems natural to unify these precursors in a theory of quantum super-symmetric pairs. This was done in [12] for a special case of general linear supergroups following ideas of [13]. The current paper is a step further to a general theory of quantum super-spherical pairs.
We are not aiming at a comprehensive exposition but focus on the special case of classical (basic) Lie superalgebras and confine ourselves with a special symmetric polarization of the root system. Unlike [12] we do not appeal to the bosonization of quantum supergroups and stay within the cateogry of Hopf superalgebras. Our approach is close to [4]. We classify all graded (generalized) Satake diagrams, for the fixed Borel subalgebra, which allow for quantization to the coideal subalgebras in the sense of [1]. We solve the Reflection Equation and relate the coideal subalgebrs to K-matrices as their quantum isotropy subgroups.
Recall that a subalgebra $\mathfrak{k}$ in a simple Lie algebra $\mathfrak{g}$ is called spherical if there is a Borel subalgebra $\mathfrak{b}\subset\mathfrak{g}$ such that $\mathfrak{k}+\mathfrak{b}=\mathfrak{g}$. The pair $(\mathfrak{g},\mathfrak{k})$ is the localization of a classical spherical manifold. Such manifolds generalize symmetric spaces and feature similar nice representation theoretical properties [14]. It turns out that all solutions to non-graded reflection equation for standard $U_{q}(\mathfrak{g})$ are classical points on spherical manifolds, where the Poisson bracket vanishes. Their isotropy Lie algebras form spherical pairs $(\mathfrak{g},\mathfrak{k})$ with the total Lie algebra $\mathfrak{g}$.
We adopt a similar definition for super-spherical pairs assuming that both $\mathfrak{g}$ and $\mathfrak{k}$ are graded. As different Borel subalgebras in $\mathfrak{g}$ are generally not isomorphic, this definition depends on the choice of $\mathfrak{b}$. We consider spherical pairs that are quantizable along the similar lines as non-graded (pseudo)symmetric pairs, [1, 2, 4].
We define a classical pseudo-super-symmetric pair $(\mathfrak{k},\mathfrak{g})$ via an involutive automorphism $\tau$ of the Dynkin diagram $D_{\mathfrak{g}}$ subordinate to a certain subdiagram $D_{\mathfrak{l}}\subset D_{\mathfrak{g}}$. Here $\mathfrak{l}$ is an analog of the semisimple Lie subalgebra in $\mathfrak{k}$ whose roots are fixed by the symmetry in the ordinary symmetric pair $(\mathfrak{g},\mathfrak{k})$. The key novelty here is a substitute for the longest Weyl group element $w_{\mathfrak{l}}$ of $\mathfrak{l}$, which is present in the non-graded case and which plays a crucial role in the theory. We introduce an operator $w_{\mathfrak{l}}$ that features basic properties of the longest Weyl group element. In particular, it is an even involution that preserves the root systems of $\mathfrak{l}$ and $\mathfrak{g}$ and their weight lattices and flips the highest and lowest weights of basic $\mathfrak{l}$-modules generating $\mathfrak{k}$ over $\mathfrak{l}$ and a certain subset of the Cartan subalgebra. Such an operator can be defined if the total grading of $\mathfrak{g}$ induces a special ”symmetric” grading on $\mathfrak{l}$.
We prove that subalgebras $\mathfrak{k}\subset\mathfrak{g}$ constructed this way are spherical and quantizable as supercoideal subalgebras in the quantum super-supergroup $U_{q}(\mathfrak{g})$. They can be conveniently parameterized by decorated Dynkin diagrams associated with $(\mathfrak{g},\Pi_{\mathfrak{l}},\tau)$, which amount to graded (generalized) Satake diagrams, upon filtering through a set of selection rules.
There are two problems arising in such a description: a) if the subalgebra $\mathfrak{k}$ is proper, that is, $\mathfrak{k}\varsubsetneq\mathfrak{g}$, and b) if $\mathfrak{k}$ is related with a solution of the Reflection Equation. We give a classification of graded Satake diagrams within the fixed grading by discarding decorated diagrams with trivial $\mathfrak{k}=\mathfrak{g}$ pairs. We indicate the diagrams related with solutions of the Reflection Equation, which we present explicitly. There are infinite families of ”anomalous” pairs $(\mathfrak{k},\mathfrak{g})$ apparently with no $K$-matrix.
It is interesting to note that contrary to the non-graded case, only one type of reflection equation appears to be relevant: that is the one equivariant with respect to the adjoint $U_{q}(\mathfrak{g})$-action on matrices.
2 Preliminaries
This section contains a general description of quantum supergroups deforming the universal enveloping algebras of general linear and orthosymplectic Lie superalgebras. We explicate their natural representations on the graded vector space $\mathbb{C}^{2\mathbf{m}|N}$ along with the graded R-matrices in a symmetric grading with minimal number of odd simple roots.
2.1 Quantum Supergroup $U_{q}(\mathfrak{g})$
For a textbook on quantum groups, the reader is referred to [15]. In our exposition of quantum supergroups, we follow [16].
An algebra $\mathcal{A}$ is called superalgebra if it is $\mathbb{Z}_{2}$-graded: that is, $\mathcal{A}=\mathcal{A}_{0}\oplus\mathcal{A}_{1}$, and $\mathcal{A}_{i}\mathcal{A}_{j}\subset\mathcal{A}_{i+j\mod 2}$. Elements of $\mathcal{A}_{0}$ are called even and elements of $\mathcal{A}_{1}$ are called odd.
Given two graded associative algebras $\mathcal{A}$ and $\mathcal{B}$ their tensor product $\mathcal{A}\otimes\mathcal{B}$ is a graded algebra too. The multiplication on homogeneous elements is determined by the rule
$(a_{1}\otimes b_{1})(a_{2}\otimes b_{2})=(-1)^{|a_{2}||b_{1}|}a_{1}a_{2}\otimes b_{1}b_{2}.$
where $|a|$ stands for the degree of $a$.
In this section, we recall a general definition of quantum supergroup $U_{q}(\mathfrak{g})$ associated with a graded Lie algebra $\mathfrak{g}$. It is a quasi-triangular Hopf superalgebra with comultiplication ranging in the graded commutative tensor square of $U_{q}(\mathfrak{g})$. The antipode in $U_{q}(\mathfrak{g})$ is a graded anti-automorphism: $\gamma(ab)=(-1)^{|a||b|}\gamma(b)\gamma(a)$ for all homogeneous $a,b\in U_{q}(\mathfrak{g})$.
Let $\mathfrak{g}$ be a finite-dimensional complex Lie superalgebra associated with a root system $\mathrm{R}$ of rank $n$ with a basis $\Pi=\{\alpha_{i}:i\in[1,n]=I\}$ of the simple roots and let $\mathfrak{h}\subset$ denote its even Cartan subalgebra $\mathfrak{h}$. Let $A=(a_{ij})_{1\leq i,j\leq n}$, be the symmetrizable Cartan matrix and (,)(-,-)( - , - ) the corresponding non-degenerate symmetric bilinear form on the dual vector space $\mathfrak{h}^{*}$.
Fix $q\in\mathbb{C}^{\times}$ not a root of unity. Define quantum integers by setting
$[n]_{q}=q^{n-1}+q^{n-3}+\dots+q^{-n+3}+q^{-n+1}=\frac{q^{n}-q^{-n}}{\omega}$
for $n\in\mathbb{Z}$, where the notation $q^{-1}=\bar{q}$ and $\omega=(q-\bar{q})$ is used here and throughout the text. For each simple root $\alpha\in\Pi$ define
$q_{\alpha}=\begin{cases}q,&\mbox{if }(\alpha,\alpha)=0,\\ q^{\frac{(\alpha,\alpha)}{2}},&\mbox{if }(\alpha,\alpha)\neq 0.\end{cases}$
Definition 2.1.
[16] The quantum supergroup $U_{q}(\mathfrak{g})$ is a complex unital associative superalgebra generated by $e_{\pm\alpha_{i}},$ and $q^{\pm h_{\alpha_{i}}}$ with grading
$|e_{\pm{\alpha_{i}}}|=\begin{cases}0,&\mbox{if }i\in\varkappa\subset I\\ 1,&\mbox{if }i\notin\varkappa.\end{cases},\quad|q^{\pm h_{\alpha_{i}}}|=0,\ \forall\ i\in I,$
such that the following relations are satisfied:
  1. (i)
    $q^{h_{\alpha_{i}}}q^{-h_{\alpha_{i}}}=q^{-h_{\alpha_{i}}}q^{h_{\alpha_{i}}}=1,$ $q^{h_{\alpha_{j}}}q^{-h_{\alpha_{i}}}=q^{-h_{\alpha_{i}}}q^{h_{\alpha_{j}}}$,
  2. (ii)
    $q^{h_{\alpha_{i}}}e_{\pm{\alpha_{j}}}q^{-h_{\alpha_{i}}}=q^{\pm(\alpha_{i},\alpha_{j})}e_{\pm\alpha_{j}}$,
  3. (iv)
    $[e_{\alpha_{i}},e_{-\alpha_{j}}]=e_{\alpha_{i}}e_{-\alpha_{j}}-(-1)^{|e_{-\alpha_{j}}||e_{\alpha_{i}}|}e_{-\alpha_{j}}e_{\alpha_{i}}=\delta_{ij}[h_{\alpha_{i}}]_{q_{\alpha_{i}}}$,
  4. (v)
    $(ad_{q^{\prime}}e_{\pm{\alpha_{i}}})^{v_{ij}}e_{\pm{\alpha_{j}}}=0,\quad i\neq j,\ q^{\prime}=q,\bar{q},$ where
    $(ad_{q^{\prime}}e_{\alpha_{i}})x=e_{\alpha_{i}}x-(-1)^{|e_{\alpha_{i}}||x|}(q^{\prime})^{(\alpha_{i},wt(x))}xe_{\alpha_{i}},\quad x\in U_{q}(\mathfrak{g}),$
    $v_{ij}=\begin{cases}1,&\mbox{if }(\alpha_{i},\alpha_{i})=(\alpha_{i},\alpha_{j})=0,\\ 2,&\mbox{if }(\alpha_{i},\alpha_{i})=0,\ (\alpha_{i},\alpha_{j})\neq 0,\\ 1-\frac{2(\alpha_{i},\alpha_{j})}{(\alpha_{i},\alpha_{i})},&\mbox{if }(\alpha_{i},\alpha_{i})\neq 0.\end{cases}$
We will work with $\mathfrak{g}=\mathfrak{g}\mathfrak{l}(n+1)$ too. Therefore, we need to extend the previous definition from $\mathfrak{g}=\mathfrak{s}\mathfrak{l}(n+1)$ to $\mathfrak{g}=\mathfrak{g}\mathfrak{l}(n+1)$. Cartan generators of $\mathfrak{g}=\mathfrak{g}\mathfrak{l}(n+1)$ are $\{q^{h_{\zeta_{i}}}\}_{i=1}^{n+1}$ with commutation relations
$q^{h_{\zeta_{i}}}e_{\pm{\alpha_{j}}}q^{-h_{\zeta_{i}}}=q^{\pm(\zeta_{i},\alpha_{j})}e_{\pm\alpha_{j}}$
we can proceed further as $q^{h_{\alpha_{i}}}=q^{h_{\zeta_{i}}}q^{-h_{\zeta_{i+1}}}$.
A Hopf superalgebra structure is fixed by comultiplication defined on the generators $e_{i}=e_{\alpha_{i}}$, $f_{i}=e_{-\alpha_{i}}$, and $q^{\pm h_{i}}=q^{h_{\alpha_{i}}}$ as
$\Delta(e_{i})=q^{h_{i}}\otimes e_{i}+e_{i}\otimes 1,\quad\Delta(f_{i})=f_{i}\otimes q^{-h_{i}}+1\otimes f_{i},\quad\Delta(q^{\pm h_{i}})=q^{\pm h_{i}}\otimes q^{\pm h_{i}}.$
The counit $\epsilon$ is a homomorphism $U_{q}(\mathfrak{g})\to\mathbb{C}$ that vanishes on all $e_{i}$, $f_{i}$ and returns $1$ on $q^{\pm h_{i}}$. The antipode $\gamma$ can be readily evaluated on the generators as
$\gamma(q^{\pm h_{i}})=q^{\mp h_{i}},\quad\gamma(e_{i})=-q^{-h_{i}}e_{i},\quad\gamma(f_{i})=-f_{i}q^{h_{i}}$
and extended as a graded anti-automorphism to entire $U_{q}(\mathfrak{g})$. This (super) coalgebra structure extends to the general linear algebra in the obvious way.
We denote by $U_{q}(\mathfrak{h}),\ U_{q}(\mathfrak{g}_{+}),\ \mathrm{and}\ U_{q}(\mathfrak{g}_{-})$ the $\mathbb{C}$-subalgebras of $U_{q}(\mathfrak{g})$, generated by $q^{\pm h_{i}},\ e_{i},$ and $f_{i}$ respectively.
Simple root vectors enter the set of generators of $U_{q}(\mathfrak{g})$, while composite root vectors are not given for granted. In the case of quantum groups they can be constructed via the Lusztig braid group action on $U_{q}(\mathfrak{g})$, which is not applicable to the graded case. Still the composite root vectors can be constructed by using a method of Khoroshkin and Tolstoy, based on the concept of normal ordering of positive roots, which is applicable in both cases. As a result, root vectors $e_{\alpha}\in U_{q}(\mathfrak{g}_{+})$ and $f_{\alpha}\in U_{q}(\mathfrak{g}_{-})$ are defined for each positive $\alpha\in\mathrm{R}^{+}$. Along with $q^{\pm h_{\alpha}}$, they generate a quantum supergroup of rank one in $U_{q}(\mathfrak{g})$. Normally ordered monomials in these elements generate a PBW basis in $U_{q}(\mathfrak{g}_{\pm})$ and therefore in the entire $U_{q}(\mathfrak{g})$ (over $U_{q}(\mathfrak{h})$). The universal R-matrix can be explicitly expressed through the root vectors as an ordered product of so-called q-exponentials over positive roots.
2.2 Basic Quantum Supergroups
We specify the root systems of basic (non-exceptional) matrix Lie superalgebras: general linear $\mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$ and ortho-symplectic $\mathfrak{o}\mathfrak{s}\mathfrak{p}(N|2\mathbf{m})$, $\mathfrak{s}\mathfrak{p}\mathfrak{o}(N|2\mathbf{m})$. We fix the following symmetric grading on underlying vector space $V=\mathbb{C}^{2\mathbf{m}+N}=C^{2\mathbf{m}|N}$:
$|i|=\left\{\begin{array}[]{cccc}1,&i\leqslant\mathbf{m}\quad\mbox{or}\quad N+\mathbf{m}<i,\\ 0,&\mathbf{m}<i\leqslant\mathbf{m}+N.\end{array}\right.$
With respect to this grading, there are two odd simple roots in $\mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$ and $\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|2\mathbf{m})$, and one odd simple root for $\mathfrak{o}\mathfrak{s}\mathfrak{p}(N|2\mathbf{m})$, $N\neq 2$. Let $\mathfrak{g}=\mathfrak{g}_{0}\oplus\mathfrak{g}_{1}$ b a Lie superalgebra of these types. Denote the Cartan subalgebra by $\mathfrak{h}\subset\mathfrak{g}_{0}$ equipped with canonical bilinear form on $\mathfrak{h}^{*}$ with a basis
$\{\varepsilon_{i}\}_{i=1}^{k},\quad\{\delta_{i}\}_{i=1}^{2\mathbf{m}},\quad(\varepsilon_{i},\varepsilon_{j})=\delta_{i}^{j}=-(\delta_{i},\delta_{j}),\ (\varepsilon_{i},\delta_{j})=0,$
where $k=N$ for $\mathfrak{g}\mathfrak{l}(N|\mathbf{m})$ and $k=\mathbf{n}$ for $\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m})$, $\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m})$, and $\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m})$. In the Dynkin diagrams below, an odd simple root $\alpha$ is denoted by a black node if $(\alpha,\alpha)\neq 0$, and a grey one if $(\alpha,\alpha)=0$. Even simple roots are represented by white nodes.
We denote the sets of even and odd roots with $\mathrm{R}_{0}$ and $\mathrm{R}_{1}$, respectively, so that $\mathrm{R}=\mathrm{R}_{0}\cup\mathrm{R}_{1}$. Additionally, we divide the root system into positive and negative parts $\mathrm{R}^{\pm}$ with the basis of simple roots $\Pi\subset\mathrm{R}^{+}$.
$\bullet\ \mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$
$\begin{cases}\mathrm{R}^{+}=\{\varepsilon_{i}-\varepsilon_{j}\}_{i<j}\cup\{\delta_{i}-\delta_{j}\}_{i<j}\cup\{\delta_{i}-\varepsilon_{j}\}_{i\leq\mathbf{m}}\cup\{\varepsilon_{j}-\delta_{i}\}_{i\geq\mathbf{m}+1},\\ \Pi=\{\delta_{i}-\delta_{i+1}\}_{i=1,i\neq\mathbf{m}}^{2\mathbf{m}-1}\cup\{\delta_{\mathbf{m}}-\varepsilon_{1}\}\cup\{\varepsilon_{i}-\varepsilon_{i+1}\}_{i=1}^{N-1}\cup\{\varepsilon_{N}-\delta_{\mathbf{m}+1}\}.\\ \end{cases}$
with $\mathrm{R}_{0}=\{\varepsilon_{i}-\varepsilon_{j},\delta_{i}-\delta_{j}\}_{i\neq j}$ and $\mathrm{R}_{1}=\pm\{\varepsilon_{i}-\delta_{j}\}$. The Dynkin diagram, which describes the relations among simple roots and indicates their parity, is
[…]
$\bullet\ \mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m})$. Now we consider separately two cases: $\mathbf{n}\neq 0$ and $\mathbf{n}=0$.
For $\mathbf{n}\neq 0$, we have
$\begin{cases}\mathrm{R}^{+}=\{\varepsilon_{i}\pm\varepsilon_{j}\}_{i<j}\cup\{\delta_{i}\pm\delta_{j}\}_{i<j}\cup\{\delta_{i}\pm\varepsilon_{j},\varepsilon_{i},2\delta_{i},\delta_{i}\},\\ \Pi=\{\delta_{i}-\delta_{i+1}\}_{i=1}^{\mathbf{m}-1}\cup\{\delta_{\mathbf{m}}-\varepsilon_{1}\}\cup\{\varepsilon_{i}-\varepsilon_{i+1}\}_{i=1}^{\mathbf{n}-1}\cup\{\varepsilon_{\mathbf{n}}\}.\\ \end{cases}$
with $\mathrm{R}_{0}=\{\pm\varepsilon_{i}\pm\varepsilon_{j},\pm 2\delta_{i},\pm\delta_{i}\pm\delta_{j},\pm\varepsilon_{i}\}_{i\neq j}$, $\mathrm{R}_{1}=\{\pm\delta_{i},\pm\delta_{i}\pm\varepsilon_{j}\}$, and Dynkin diagram
[…]
In the case of $\mathbf{n}=0$, we have
$\begin{cases}\mathrm{R}^{+}=\{\delta_{i}\pm\delta_{j}\}_{i<j}\cup\{2\delta_{i},\delta_{i}\},\\ \Pi=\{\delta_{i}-\delta_{i+1}\}_{i=1}^{\mathbf{m}-1}\cup\{\delta_{\mathbf{m}}\}.\\ \end{cases}$
with $\mathrm{R}_{0}=\{\pm 2\delta_{i},\pm\delta_{i}\pm\delta_{j}\}_{i\neq j}$, $\mathrm{R}_{1}=\{\pm\delta_{i}\}$. The Dynkin diagram is
$\ldots$$>$$\delta_{\mathbf{m}}$$\delta_{1}-\delta_{2}$$\delta_{2}-\delta_{3}$$\delta_{\mathbf{m}-1}-\delta_{\mathbf{m}}$
$\bullet\ \mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m})$. Here we also distinguish two cases: $\mathbf{n}\neq 1$ and $\mathbf{n}=1$.
For $\mathbf{n}\neq 1$, we have
$\begin{cases}\mathrm{R}^{+}=\{\varepsilon_{i}\pm\varepsilon_{j}\}_{i<j}\cup\{\delta_{i}\pm\delta_{j}\}_{i<j}\cup\{\delta_{j}-\varepsilon_{i},2\delta_{i}\},\\ \Pi=\{\delta_{i}-\delta_{i+1}\}_{i=1}^{\mathbf{m}-1}\cup\{\delta_{\mathbf{m}}-\varepsilon_{1}\}\cup\{\varepsilon_{i}-\varepsilon_{i+1}\}_{i=1}^{\mathbf{n}-1}\cup\{\varepsilon_{\mathbf{n}-1}+\varepsilon_{\mathbf{n}}\}.\\ \end{cases}$
with $\mathrm{R}_{0}=\{\pm\varepsilon_{i}\pm\varepsilon_{j},\pm 2\delta_{i},\pm\delta_{i}\pm\delta_{j},\}_{i\neq j}$, $\mathrm{R}_{1}=\{\pm\delta_{i}\pm\varepsilon_{j}\}$, and the Dynkin diagram.
[…]
When $\mathbf{n}=1$, the root system is
$\begin{cases}\mathrm{R}^{+}=\{\delta_{i}\pm\delta_{j}\}_{i<j}\cup\{\delta_{j}\pm\varepsilon_{1},2\delta_{i}\},\\ \Pi=\{\delta_{i}-\delta_{i+1}\}_{i=1}^{\mathbf{m}-1}\cup\{\delta_{\mathbf{m}}-\varepsilon_{1}\}\cup\{\delta_{\mathbf{m}}+\varepsilon_{1}\}.\end{cases}$
$\mathrm{R}_{0}=\{\pm 2\delta_{i},\pm\delta_{i}\pm\delta_{j},\}_{i\neq j}$ and $\mathrm{R}_{1}=\{\pm\delta_{i}\pm\varepsilon_{1}\}$. Its Dynkin diagram is
[…]
$\bullet\ \mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m})$ has the root system
$\begin{cases}\mathrm{R}^{+}=\{\varepsilon_{i}\pm\varepsilon_{j}\}_{i<j}\cup\{\delta_{i}\pm\delta_{j}\}_{i<j}\cup\{\delta_{j}-\varepsilon_{i},2\varepsilon_{i}\},\\ \Pi=\{\delta_{i}-\delta_{i+1}\}_{i=1}^{\mathbf{m}-1}\cup\{\delta_{\mathbf{m}}-\varepsilon_{1}\}\cup\{\varepsilon_{i}-\varepsilon_{i+1}\}_{i=1}^{\mathbf{n}-1}\cup\{2\varepsilon_{\mathbf{n}}\},\\ \end{cases}$
with $\mathrm{R}_{0}=\{\pm\varepsilon_{i}\pm\varepsilon_{j},\pm 2\varepsilon_{i},\pm\delta_{i}\pm\delta_{j},\}_{i\neq j}$, $\mathrm{R}_{1}=\{\pm\delta_{i}\pm\varepsilon_{j}\}$, and the Dynkin diagram
[…]
2.3 Natural Representations
In this section we describe an irreducible representation of $U_{q}(\mathfrak{g})$ on the graded vector space $\mathrm{End}(\mathbb{C}^{N|2\mathbf{m}})$. Let $\{v_{i}\}\subset\mathbb{C}^{N|2\mathbf{m}}$ the standard homogeneous weight basis. We call the grading symmetric if $|v_{i}|=|v_{i^{\prime}}|$ for all $i$, where $i^{\prime}=N+2\mathbf{m}+1-i$. We choose a symmetric grading with minimal number of odd simple roots:
$\left(\underbrace{1,\ldots,1}_{\mathbf{m}};\underbrace{0,\ldots,0}_{N};\underbrace{1,\ldots,1}_{\mathbf{m}}\right).$
Let $\{e_{ij}\}_{i,j=1}^{N+2m}\subset\mathrm{End}(\mathbb{C}^{N|2\mathbf{m}})$ be the standard matrix basis. An irreducible representation $\pi:U_{q}(\mathfrak{g})\rightarrow\mathrm{End}(\mathbb{C}^{N|2\mathbf{m}})$ is defined by the following assignment:
$q^{h_{\zeta_{i}}}\mapsto\ \sum_{j=1}^{N}q^{(-1)^{|i|}\delta_{j}^{i}}e_{jj},\quad\mathrm{if}\ i\leq n+1,\ \mathfrak{g}=\mathfrak{g}\mathfrak{l}(n+1),\$
$q^{h_{i}}\mapsto$$\begin{cases}\sum_{j=1}^{N+2\mathbf{m}}q^{(-1)^{|i|}(\delta_{j}^{i}-\delta_{j}^{i^{\prime}})+(-1)^{|i+1|}(-\delta_{j}^{(i+1)}+\delta_{j}^{(i+1)^{\prime}})}e_{jj},&\mathrm{if}\ i<\mathbf{n}+\mathbf{m},\ \mathfrak{g}=\begin{cases}\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m}),\\ \mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m}),\\ \mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m}),\end{cases}\\ \sum_{j=1}^{N+2\mathbf{m}}q^{(-1)^{|i|}(\delta_{j}^{i}-\delta_{j}^{i^{\prime}})}e_{jj},&\mathrm{if}\ i=\mathbf{m}+\mathbf{n},\ \mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m}),\\ \sum_{j=1}^{N+2\mathbf{m}}q^{2\delta_{j}^{i}-2\delta_{j}^{i^{\prime}}}e_{jj},&\mathrm{if}\ i=\mathbf{m}+\mathbf{n},\ \mathfrak{g}=\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m}),\\ \sum_{j=1}^{N+2\mathbf{m}}q^{(-1)^{|i-1|}(\delta_{j}^{(i-1)}-\delta_{j}^{(i-1)^{\prime}})}q^{(-1)^{|i|}(\delta_{j}^{i}-\delta_{j}^{i^{\prime}})}e_{jj},&\mathrm{if}\ i=\mathbf{m}+\mathbf{n},\ \mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m}),\\ \end{cases}$
$e_{i}\mapsto$$\begin{cases}e_{i,i+1},&\mathrm{if}\ i\leq N+2\mathbf{m}-1,\ \mathfrak{g}=\mathfrak{g}\mathfrak{l}(N|2\mathbf{m}),\\ q^{-\delta_{i}^{\mathbf{m}}}e_{i,i+1}-(-1)^{(|i|)(|i+1|+1)}e_{(i+1)^{\prime},i^{\prime}},&\mathrm{if}\ i\leq\mathbf{m}+\mathbf{n},\ \mathfrak{g}=\begin{cases}\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m}),\\ \mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m}),\\ \mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m}),\end{cases}\\ q^{-\delta_{i-1}^{\mathbf{m}}}e_{i-1,i+1}-(-1)^{|i-1|}e_{(i+1)^{\prime},(i-1)^{\prime}},&\mathrm{if}\ i=\mathbf{m}+\mathbf{n},\ \mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m}),\\ e_{i,i+1},&\mathrm{if}\ i=\mathbf{m}+\mathbf{n},\ \mathfrak{g}=\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m}),\\ \end{cases}$
$f_{i}\mapsto\begin{cases}-e_{i+1,i},&\mbox{if }\ i=\mathbf{m},\ \mathfrak{g}=\mathfrak{g}\mathfrak{l}(N|2\mathbf{m}),\\ -qe_{i+1,i}+e_{i^{\prime},(i+1)^{\prime}},&\mbox{if }\ i=\mathbf{m},\ \mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(N|2\mathbf{m}),\ \mathrm{or}\ \mathfrak{g}=\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m}),\\ -qe_{i+1,i-1}+e_{(i-1)^{\prime},(i+1)^{\prime}},&\mbox{if }\ i=\mathbf{m}+1,\ \mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|2\mathbf{m}),\end{cases}$
and the similar assignment to the remaining $f_{i}$ as for $e_{i}$ with $e_{ij}$ changed to $e_{ji}$. This representation is compatible with the R-matrices.
$R=\sum_{i,j}q^{(-1)^{|i|}\delta_{ij}}e_{ii}\otimes e_{jj}+\omega\sum_{j<i}(-1)^{|j|}e_{ij}\otimes e_{ji},$ (2.1)
for the general linear Lie [17, 18] and
$R=\sum_{i,j=1}q^{(-1)^{|j|}(\delta_{ij}-\delta_{ij^{\prime}})}e_{ii}\otimes e_{jj}+\omega\sum\limits_{\begin{subarray}{c}j,i=1\\ j<i\end{subarray}}((-1)^{|j|}e_{ij}\otimes e_{ji}-(-1)^{|i|+|j|+|i||j|}\kappa_{i}\kappa_{j}q^{\rho_{i}-\rho_{j}}e_{ij}\otimes e_{i^{\prime}j^{\prime}}),$ (2.2)
for the ortho-symplectic Lie superalgebras [17, 19]. Here
$(\rho_{i})=\begin{cases}(k-\mathbf{m},\dots,k-1;k-1,\dots,\frac{1}{2},0,-\frac{1}{2},\dots,1-k;1-k,\dots,\mathbf{m}-k),\ k=\frac{2\mathbf{n}+1}{2},\ \mathrm{for}\ \mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m}),\\ (\mathbf{n}-\mathbf{m},\dots,\mathbf{n}-1;\mathbf{n}-1,\dots,1,0,0,-1,\dots,1-\mathbf{n};1-\mathbf{n},\dots,\mathbf{m}-\mathbf{n}),\ \mathrm{for}\ \mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m}),\\ (\mathbf{n}-\mathbf{m}+1,\dots,\mathbf{n};\mathbf{n},\dots,1,-1,\dots,1-\mathbf{n};1-\mathbf{n},\dots,\mathbf{m}-\mathbf{n}-1),\ \mathrm{for}\ \mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m}),\end{cases}$
$(\kappa_{i})=\begin{cases}(-1,\dots,-1;1,\dots,1;1,\dots,1),\ \mathrm{for}\ \mathfrak{o}\mathfrak{s}\mathfrak{p}(N|2\mathbf{m}),\\ (-1,\dots,-1;1,\dots,1,-1,\dots,-1;-1,\dots,-1),\ \mathrm{for}\ \mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m}).\end{cases}$
3 Solutions to Graded Reflection Equation
The non-graded version of Reflection Equation appeared in mathematical physics literature [22, 23] and triggered the theory of quantum symmetric pairs in [24, 1]. The graded Reflection Equation has been of interest as well, mostly in the spectral parameter dependent form [25, 26, 27, 28].
In this section, we present solutions to the constant RE for general linear and orthosymplectic quantum supergroups. We are interested in invertible even RE-matrices. Solutions for the general linear case are taken from our recent paper [20].
3.1 $\mathbb{Z}_{2}$-graded Reflection Equation
Suppose that $V$ is a graded vector space and $\mathcal{A}=\mathrm{End}(V)$ is the corresponding graded matrix algebra. An invertible element $R\in\mathcal{A}\otimes\mathcal{A}$ is called an $R$-matrix if it satisfies Yang-Baxter equation
$R_{12}R_{13}R_{23}=R_{23}R_{13}R_{12},$
where the subscripts indicate the tensor factor in the graded tensor cube of $\mathrm{End}(V)$. An even element $P=\sum_{i=1}^{n}(-1)^{|j|}e_{ij}\otimes e_{ji}\in\mathrm{End}(V)\otimes\mathrm{End}(V)$ is called graded permutation. It flips the tensor factors in $V\otimes V$ according to
$P(v\otimes w)=(-1)^{|v||w|}w\otimes v$
for all homogeneous $v,w\in V$. The operator $S=PR\in\mathrm{End}(V)\otimes\mathrm{End}(V)$ satisfies the braid relation
$S_{12}S_{23}S_{12}=S_{23}S_{12}S_{23}.$
A matrix $K\in\mathrm{End}(V)$ is said to satisfy reflection equation if the identity
$SK_{2}SK_{2}=K_{2}SK_{2}S$
holds true in $\mathrm{End}(V)\otimes\mathrm{End}(V)$. We will assume that $K$ is even.
The case of the general linear quantum supergroup has been studied in detail in [20], where the full list of solutions to RE is given relative to an arbitrary grading of $V$. They turn out to be exactly even solutions to the non-graded RE. Invertible solutions occur only for a (arbitrary) symmetric grading (here we use a chance to correct a sloppy remark on page 6 of [20]) and have a form
$K=(\lambda+\mu)\sum_{i=1}^{m}e_{ii}+\lambda\sum_{i=m+1}^{N+2\mathbf{m}-m}e_{ii}+\sum_{i=1}^{m}y_{i}e_{ii^{\prime}}+\sum_{i=1}^{m}y_{i^{\prime}}e_{i^{\prime}i},$ (3.3)
with $y_{i}y_{i^{\prime}}=-\lambda\mu\not=0$. This matrix solves the RE with the R-matrix (2.1). We present solutions for the ortho-symplectic quantum supergroups in the next section.
3.2 Reflection Equation for ortho-symplectic quantum groups
In this section, we find even solutions to RE of the quantised orthosymplectic superalgebra. We introduce a grading the underlying vector space $V=\mathbb{C}^{2\mathbf{m}+N}$ by setting
$|i|=\left\{\begin{array}[]{cccc}1,&i\leqslant\mathbf{m}&\mbox{or}&N+\mathbf{m}<i,\\ 0,&\mathbf{m}<i\leqslant\mathbf{m}+N.\end{array}\right.$
This is a symmetric grading with the smallest number of odd simple roots. We are looking for solutions of the RE of the following three forms:
$A=\sum_{i=1}^{N+2\mathbf{m}}x_{i}e_{ii}+\sum_{i=1}^{N+2\mathbf{m}}y_{i}e_{ii^{\prime}},\quad B=\sum_{i=1}^{N+2\mathbf{m}}x_{i}e_{ii}+\sum_{i=1}^{N+2\mathbf{m}}y_{i}(e_{i,i^{\prime}-1}-e_{i+1,i^{\prime}}),$
$C=\sum\limits_{\begin{subarray}{c}i=1\\ i\neq n\end{subarray}}^{2n-1}x_{i}(e_{i,i^{\prime}-1}-e_{i+1,i^{\prime}})+x_{n}e_{n,n^{\prime}}+x_{n^{\prime}}e_{n^{\prime},n},\quad\mathrm{for}\ \mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|4\mathbf{m}).$
The matrices $A$ and $B$ will depend on an integer parameter $m$ subject to inequality $1\leq m\leq\mathbf{m}$ in the forthcoming theorem.
Theorem 3.1.
The following matrices solve the reflection equation associated with ortho-symplectic quantum groups:
$A=\sum_{i=1}^{m}\lambda(1-\kappa_{m}\kappa_{m^{\prime}}q^{-2\rho_{m}})e_{ii}+\sum_{m<i<m^{\prime}}\lambda e_{ii}+\sum\limits_{\begin{subarray}{c}1\leqslant i,i^{\prime}\leqslant m\end{subarray}}y_{i}e_{i,i^{\prime}},$ (3.4)
$\begin{split}B=\sum_{i=1}^{m}&\lambda(1+\kappa_{m-1}\kappa_{m^{\prime}+1}q^{-2(\rho_{m}+1)})e_{ii}+\sum_{m<i<m^{\prime}}\lambda e_{ii}\hskip 30.0pt\\ &\hskip 30.0pt+\sum\limits_{\begin{subarray}{c}i\leqslant m\\ i\in 2\mathbb{Z}+1\end{subarray}}\bigl{(}z_{i}(e_{i,i^{\prime}-1}-e_{i+1,i^{\prime}})+z_{i^{\prime}-1}(e_{i^{\prime}-1,i}-e_{i^{\prime},i+1})\bigr{)},\end{split}$ (3.5)
$C=\sum\limits_{\begin{subarray}{c}i=1\\ i\in 2\mathbb{N}+1\\ i\neq n\end{subarray}}^{2n-1}x_{i}(e_{i,i^{\prime}-1}-e_{i+1,i^{\prime}})+x_{n}e_{n,n^{\prime}}+x_{n^{\prime}}e_{n^{\prime},n},$ (3.6)
where the parameters $y_{i},z_{i},x_{i}\in\mathbb{C}$ satisfy the conditions
  • $y_{i}y_{i^{\prime}}=\kappa_{m}\kappa_{m^{\prime}}\lambda^{2}q^{-2\rho_{m}}$,
  • $z_{i}z_{i^{\prime}-1}=-\kappa_{m-1}\kappa_{m^{\prime}+1}\lambda^{2}q^{-2(\rho_{m}+1)}$,
  • $x_{i}x_{i^{\prime}-1}=x_{n}x_{n^{\prime}}$.
Proof.
Direct calculation. ∎
These matrices have the following shape:
$A=$$\scriptstyle y_{1}$$\scriptstyle\lambda+\mu$$\scriptstyle\lambda+\mu$$\scriptstyle\lambda$$\scriptstyle\lambda$$\scriptstyle y_{m}$$\scriptstyle y_{m^{\prime}}$$\scriptstyle y_{1^{\prime}}$
,       $B=$
$\scriptstyle z_{1}\nu$
$\scriptstyle\lambda+\mu$$\scriptstyle\lambda+\mu$
$\scriptstyle\lambda$
$\scriptstyle\lambda$
$\scriptstyle z_{m-1}\nu$
$\scriptstyle z_{m^{\prime}}\nu$$\scriptstyle z_{2^{\prime}}\nu$   ,
$C=$$\scriptstyle x_{n}$$\scriptstyle x_{n+1}$$\scriptstyle x_{1}\nu$$\scriptstyle x_{n-2}\nu$$\scriptstyle x_{n+2}\nu$$\scriptstyle x_{2^{\prime}}\nu$
with $\nu=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}$ and $\mu=-\lambda\kappa_{m}\kappa_{m^{\prime}}q^{-2\rho_{m}}$ for $A$, $\mu=\lambda\kappa_{m-1}\kappa_{m^{\prime}+1}q^{-2(\rho_{m}+1)}$ for $B$. In both cases of $A$ and $B$, the blocks with diagonal entries of $\lambda+\mu$ and $\lambda$ are, respectively, of size $m$ and $N+2\mathbf{m}-2m$, with $m\leq\mathbf{m}$ as stipulated. Our next goal is to relate them with coideal subalgebras in $U_{q}(\mathfrak{g})$.
4 Classical super-spherical pairs
Let $\mathfrak{g}$ be a Lie super-algebra and $\mathfrak{b}\subset\mathfrak{g}$ its Borel subalgebra.
Definition 4.1.
A Lie super-algebra $\mathfrak{k}\subset\mathfrak{g}$ is called spherical if $\mathfrak{g}=\mathfrak{k}+\mathfrak{b}$. The pair $(\mathfrak{g},\mathfrak{k})$ is called super-spherical.
Note that, contrary to the non-graded case, different Borel subalgebras in $\mathfrak{g}$ are not isomorphic, so this definition depends upon a choice of $\mathfrak{b}$. From now on we restrict our consideration to the case when $\mathfrak{g}$ is either general linear or ortho-symplectic type.
Let us fix a graded basis of weights $\zeta_{i}$, $i=1,\ldots,N$, in natural $\mathfrak{g}$-module $\mathbb{C}^{N}$. They generate the weight lattice $\Lambda$ of $\mathfrak{g}$. As before we use the notation $i^{\prime}=N+1-i$ for all $i=1,\ldots,N$.
Definition 4.2.
  1. 1.
    A weight lattice map $w_{\mathfrak{g}}\colon\Lambda\mapsto\Lambda$, $\zeta_{i}\mapsto\zeta_{i^{\prime}}$, $i=1,\ldots,N,$ is called Weyl operator.
  2. 2.
    The grading on $\mathbb{C}^{N}$ is called symmetric if the Weyl operator is even.
Clearly, a grading is symmetric if and only if every inversion $\sigma_{i}\colon\zeta_{i}\mapsto\zeta_{i^{\prime}}$ is even. It is also obvious that even $w_{\mathfrak{g}}=\prod_{i=1}^{N}\sigma_{i}$ extends to an involutive orthogonal operator $w_{\mathfrak{g}}\colon\mathfrak{h}^{*}\to\mathfrak{h}^{*}$.
From now on we will consider only the minimal symmetric grading on $\mathfrak{g}$. Consider the Dynkin diagram $D$ of a basic Lie superalgebra $\mathfrak{g}$. If we discard the grading information, then we get a diagram $\tilde{D}$ that we call shape of $D$. In the special case of $\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|2\mathbf{m})$, we understand by $\tilde{D}$ the Dynkin diagram of $\mathfrak{g}=\mathfrak{s}\mathfrak{o}(2+2\mathbf{m})$. Let $\tilde{W}$ denote the group of automorphisms of $\tilde{D}$. Denote by $\tilde{W}_{0}$ its subgroup that preserves the grading.
Lemma 4.3.
The group $\tilde{W}_{0}$ preserves the root system $\mathrm{R}$ and the weight lattice $\Lambda$.
Proof.
For general linear $\mathfrak{g}$, the group $\tilde{W}$ is generated by all permutations of $\zeta_{i}$. For orthosymplectic $\mathfrak{g}$, it is generated by transpositions $\zeta_{i}\leftrightarrow\zeta_{j}$ and inversions $\sigma_{i}\colon\zeta_{i}\leftrightarrow-\zeta_{i}$, $i,j\leqslant\frac{N}{2}$. Now the statement follows from the explicit description of the root systems given in Section 2.2. ∎
It now follows that even Weyl operator $w_{\mathfrak{g}}$ preserves the root system $\mathrm{R}$.
Remark 4.4.
So defined $w_{\mathfrak{g}}$ is analogous to the longest element of the Weyl group of the ordinary simple Lie algebras. In the case of $\mathfrak{g}=\mathfrak{s}\mathfrak{o}(2n)$ of odd rank, the operator $w_{\mathfrak{g}}$ differs from the longest element by the non-trivial automorphism of the Dynkin diagram (the tail flip). This operator can be equally used for construction of non-graded pseudo-symmetric pairs, because it flips the highest and lowest weights of the minimal $\mathfrak{g}$-module, cf. Lemma 4.8. It is easier to define than the ”honest” analog of the longest Weyl group element for $\mathfrak{g}=\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m})$ with odd $\mathbf{m}$. This explains our choice of $w_{\mathfrak{g}}$.
Pick up a subset $\Pi_{\mathfrak{l}}\subset\Pi$, then put $\bar{\Pi}_{\mathfrak{l}}=\Pi\backslash\Pi_{\mathfrak{l}}$, and generate a subalgebra $\mathfrak{l}=\langle e_{\alpha},f_{\alpha}\rangle_{\alpha\in\Pi_{\mathfrak{l}}}\subset\mathfrak{g}$. It is a sum of subalgebras, $\mathfrak{l}=\sum_{i}\mathfrak{l}_{i}$, corresponding to connected components of $\Pi_{\mathfrak{l}}$. If $\mathfrak{l}_{i}$ is of type A, then set $\hat{\mathfrak{l}}_{i}\subset\mathfrak{g}$ to be the natural $\mathfrak{g}\mathfrak{l}$-extension of $\mathfrak{l}_{i}$ and leave $\hat{\mathfrak{l}}_{i}=\mathfrak{l}_{i}$ otherwise. Denote by $\hat{\mathfrak{l}}=\oplus_{i}\hat{\mathfrak{l}}_{i}\subset\mathfrak{g}$ and by $\mathfrak{h}_{\hat{\mathfrak{l}}}^{*}$ its Cartan subaglebra. The restriction of the canonical inner product from $\mathfrak{h}^{*}$ to $\mathfrak{h}^{*}_{\hat{\mathfrak{l}}}$ is non-degenerate.
For the $i$-th connected component of $\Pi_{\mathfrak{l}}$ denote $w_{\mathfrak{l}_{i}}=w_{\hat{\mathfrak{l}}_{i}}$ and define $w_{\mathfrak{l}}=\prod_{i}w_{\mathfrak{l}_{i}}\in\mathrm{End}(\mathfrak{h}^{*}_{\mathfrak{l}}$), the Weyl operator of the subalgebra $\hat{\mathfrak{l}}$.
Definition 4.5.
We call $\Pi_{\mathfrak{l}}$ admissible if the grading on $\mathfrak{g}$ induces a minimal symmetric grading on each connected component of $\hat{\mathfrak{l}}$.
Here is an easy consequence of this definition.
Proposition 4.6.
An admissible set $\Pi_{\mathfrak{l}}$ cannot contain an isolated grey odd root.
Proof.
Indeed, if $\alpha\in\Pi_{\mathfrak{l}}$ is such a root, then the corresponding subalgebra $\hat{\mathfrak{l}}_{k}$ is isomorphic to $\mathfrak{g}\mathfrak{l}(1|1)$. It has no symmetric grading, and the operator $w_{\mathfrak{l}_{k}}$ is not even. ∎
Note that an isolated black odd root is possible: one can take $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(1|2\mathbf{m})$ and $\Pi_{\mathfrak{l}}=\{\alpha_{\mathbf{m}}\}$.
If $\Pi_{\mathfrak{l}}$ is admissible, then the operator $w_{\mathfrak{l}}$ is even.
Lemma 4.7.
For admissible $\Pi_{\mathfrak{l}}$, the operator $w_{\mathfrak{l}}\in\mathrm{End}(\mathfrak{h}^{*}_{\hat{\mathfrak{l}}})$ extends to an even involutive operator on $\mathfrak{h}^{*}$ that preserves the weight lattices and root systems of $\mathfrak{l}$ and $\mathfrak{g}$.
Proof.
Since the inner product on $\mathfrak{h}^{*}_{\hat{\mathfrak{l}}}$ is non-degenerate, we have orthogonal decomposition $\mathfrak{h}^{*}=(\mathfrak{h}_{\hat{\mathfrak{l}}}^{*})^{\perp}\oplus\mathfrak{h}^{*}_{\hat{\mathfrak{l}}}$. The orthogonal projection to $\hat{\mathfrak{l}}$ commutes with the parity operator because connected components of A-type in $\hat{\mathfrak{l}}$ correspond to $\mathfrak{g}\mathfrak{l}$. Then the extension of $w_{\mathfrak{l}}$ to $\mathfrak{h}^{*}$ by the identical operator on $(\mathfrak{h}_{\hat{\mathfrak{l}}}^{*})^{\perp}$ satisfies the requirement.
Finally, $w_{\mathfrak{l}}\in\tilde{W}_{0}$ and therefore preserves $\Lambda$ and $\mathrm{R}$, by Lemma 4.3. ∎
The vector subspaces
$\mathfrak{m}_{+}=\sum_{\alpha\in\mathrm{R}^{+}_{\mathfrak{g}}\>-\>\mathrm{R}^{+}_{\mathfrak{l}}}\mathfrak{g}_{\alpha},\quad\mathfrak{m}_{-}=\sum_{\alpha\in\mathrm{R}^{+}_{\mathfrak{g}}\>-\>\mathrm{R}^{+}_{\mathfrak{l}}}\mathfrak{g}_{-\alpha},$
are graded $\mathfrak{l}$-modules. For $\alpha\in\bar{\Pi}_{\mathfrak{l}}$ let $V^{\pm}_{\alpha}\subset\mathfrak{g}_{\pm}$ denote the $\mathfrak{l}$-submodule generated by $e_{\pm\alpha}\in\mathfrak{g}_{\pm\alpha}$.
Lemma 4.8.
Suppose that $\mathfrak{g}$ is a basic Lie superalgebra and $\Pi_{\mathfrak{l}}\subset\Pi$ is admissible. Then the operator $w_{\mathfrak{l}}$ flips the highest and lowest weights of $V^{\pm}_{\alpha}$ for each $\alpha\in\bar{\Pi}_{\mathfrak{l}}$.
Proof.
The subalgebra $\mathfrak{l}$ is a sum of $\mathfrak{l}=\sum_{i}\mathfrak{l}_{i}$ of basic Lie superalgebras $\mathfrak{l}_{i}$ corresponding to connected components of the Dynkin diagram of $\mathfrak{l}$. In all cases excepting $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m})$, the modules $V^{\pm}_{\alpha}$ are tensor products of the fundamental $\mathfrak{l}_{i}$-modules of minimal dimension, for which $w_{\mathfrak{l}}$ does the job. We are left to consider the case of $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m})$.
The only possible module $V^{\pm}_{\alpha}$ which is not minimal for $\mathfrak{l}_{i}\subset\mathfrak{l}$ occurs when $\alpha_{\mathbf{n}-1}\in\Pi_{\mathfrak{l}_{i}}$, $\alpha_{\mathbf{n}}\in\bar{\Pi}_{\mathfrak{l}}$ (or the other way around) and $\mathrm{rk}\>\mathfrak{l}_{i}\geqslant 3$. Now notice that $w_{\mathfrak{l}_{i}}$ is even if and only if the tail roots and $\mathfrak{l}_{i}$ are even (mind that we work with the minimal symmetric grading on $\mathfrak{g}$), in which case $w_{\mathfrak{l}_{i}}$ is just the longest element of the Weyl group of $\mathfrak{l}_{i}$. Then the statement is true either. ∎
We will consider only admissible $\Pi_{\mathfrak{l}}$. Suppose that $\tau\in\mathrm{Aut}(\Pi)$ is an even permutation that coincides with $-w_{\mathfrak{l}}$ on $\Pi_{\mathfrak{l}}$. Set $\tilde{\alpha}=w_{\mathfrak{l}}\circ\tau(\alpha)\in\mathrm{R}^{+}$ for $\alpha\in\bar{\Pi}_{\mathfrak{l}}$.
Definition 4.9.
The triple $(\mathfrak{g},\mathfrak{l},\tau)$ is called pseudo-super-symmetric if
$(\mu+\tilde{\mu},\alpha)$ $=0,$ $\quad\forall\mu\in\bar{\Pi}_{\mathfrak{l}},\quad\alpha\in\Pi_{\mathfrak{l}},$ (4.7)
$(\mu+\tilde{\mu},\nu-\tilde{\nu})$ $=0,$ $\quad\forall\mu,\nu\in\bar{\Pi}_{\mathfrak{l}}.$ (4.8)
Lemma 4.10.
Condition (4.7) implies that $\tau\circ w_{\mathfrak{l}}=w_{\mathfrak{l}}\circ\tau$.
Proof.
By construction, $\tau$ and $w_{\mathfrak{l}}$ commute when restricted to $\mathfrak{h}^{*}_{\mathfrak{l}}$. We are left to check that on simple roots from $\bar{\Pi}_{\mathfrak{l}}$.
By Lemma 4.8, $w_{\mathfrak{l}}$ takes the highest weight of an irreducible $\mathfrak{l}$-module $V^{\pm}_{\mu}$, $\mu\in\bar{\Pi}_{\mathfrak{l}}$, to the lowest weight and vice versa. For each $\mu\in\bar{\Pi}_{\mathfrak{l}}$ we have $w_{\mathfrak{l}}(\mu)=\mu+\eta$ for some $\eta\in\mathbb{Z}_{+}\Pi_{\mathfrak{l}}$ that satisfies $w_{\mathfrak{l}}(\eta)=-\eta$ because $w_{\mathfrak{l}}^{2}=\mathrm{id}$. The weight $\eta$ depends only on the projection of $\tau(\mu)$ to $\mathfrak{h}^{*}_{\mathfrak{l}}$, therefore $-\tau(\mu)=w_{\mathfrak{l}}(-\tilde{\mu})=-w_{\mathfrak{l}}\bigl{(}\tau(\mu)\bigr{)}+\eta$ or $\tau(\mu)+\eta=w_{\mathfrak{l}}\bigl{(}\tau(\mu)\bigr{)}$ because of the condition (4.7). Then, since $\tau(\eta)=-w_{\mathfrak{l}}(\eta)$ for all $\eta\in\mathfrak{h}^{*}_{\mathfrak{l}}$,
$\tau\bigl{(}w_{\mathfrak{l}}(\mu)\bigr{)}=\tau(\mu+\eta)=\tau(\mu)+\tau(\eta)=\tau(\mu)-w_{\mathfrak{l}}(\eta)=\tau(\mu)+\eta=w_{\mathfrak{l}}\bigl{(}\tau(\mu)\bigr{)},$
as required. ∎
Define an even linear map $\theta=-w_{\mathfrak{l}}\circ\tau\colon\mathfrak{h}^{*}\to\mathfrak{h}^{*}$. It preserves $\mathrm{R}$ because $\theta(\alpha)=-\tilde{\alpha}$ for $\alpha\in\bar{\Pi}_{\mathfrak{l}}$ and $\theta(\alpha)=\alpha$ for $\alpha\in\Pi_{\mathfrak{l}}$. The system of equalities (4.7) and (4.8) is equivalent to
$\bigl{(}\alpha+\theta(\alpha),\beta-\theta(\beta)\bigr{)}=0,\quad\forall\alpha,\beta\in\Pi.$ (4.9)
Proposition 4.11.
Condition (4.9) is fulfilled if and only if the permutation $\tau$ extends to an involutive even orthogonal operator on $\mathfrak{h}^{*}$ coinciding with $-w_{\mathfrak{l}}$ on $\Pi_{\mathfrak{l}}$.
Proof.
First of all remark that a linear operator being orthogonal involutive is the same as symmetric involutive, or orthogonal and symmetric simultaneously. Also, an automorphism of the Dynkin diagram is exactly an orthogonal operator on $\mathfrak{h}^{*}$ that permutes the basis $\Pi$.
Condition (4.9) is bilinear and therefore holds true for any pair of vectors from $\mathfrak{h}^{*}$. Setting $\alpha=\beta$ in (4.9) we find that $\theta$ is an isometry. Then (4.9) translates to
$\bigl{(}\theta(\alpha),\beta\bigr{)}=\bigl{(}\alpha,\theta(\beta)\bigr{)},\quad\forall\alpha,\beta\in\Pi.$
That is, $\theta$ is a symmetric operator. Therefore it is an involutive orthogonal operator. So is $w_{\mathfrak{l}}$ that commutes with $\theta$ by Lemma 4.10, hence $\tau$ is an involutive orthogonal operator on $\mathfrak{h}^{*}$.
Conversely, suppose that $\tau$ is orthogonal, involutive, and coincides with $-w_{\mathfrak{l}}$ on $\Pi_{\mathfrak{l}}$. Then
$(\alpha,\mu)=\bigl{(}\tau(\alpha),\tau(\mu)\bigr{)}=-\bigl{(}w_{\mathfrak{l}}(\alpha),\tau(\mu)\bigr{)}=-\bigl{(}\alpha,w_{\mathfrak{l}}\circ\tau(\mu)\bigr{)}=-(\alpha,\tilde{\mu})$
for all $\alpha\in\Pi_{\mathfrak{l}}$ and $\mu\in\bar{\Pi}_{\mathfrak{l}}$. Thus, the condition (4.7) is fulfilled, and $\tau$ commutes with $w_{\mathfrak{l}}$ by Lemma 4.10. Therefore $\theta=-w_{\mathfrak{l}}\circ\tau$ is orthogonal and involutive, which implies (4.9). ∎
We arrive at a necessary condition for a triple $(\mathfrak{g},\mathfrak{l},\tau)$ to be pseudo-super-symmetric.
Corollary 4.12.
Suppose that $\tau$ satisfies conditions of Proposition 4.11. Then $\mathfrak{l}$-modules $\mathfrak{m}_{+}$ and $\mathfrak{m}_{-}$ are isomorphic.
Proof.
[…]
The isomorphism between $V_{\alpha}^{\pm}$ and $V_{\alpha^{\prime}}^{\mp}$, $\alpha^{\prime}=\tau(\alpha)$, is a consequence of (4.7). If we extend $\mathfrak{l}$ by adding $\mathfrak{t}=\mathrm{Span}\{h_{\tilde{\alpha}}-h_{\alpha}\}$, for all $\alpha\in\bar{\Pi}_{\mathfrak{l}}$, then $V_{\alpha}^{\pm}$ and $V_{\alpha^{\prime}}^{\mp}$ will be isomorphic as $\mathfrak{l}+\mathfrak{t}$-modules, provided (4.8) is fulfilled. The same will be true for $\mathfrak{m}_{+}$ and $\mathfrak{m}_{-}$.
Suppose that $\Pi_{\mathfrak{l}}$ is admissible and $\tau$ satisfies the conditions of Proposition 4.11. For each $\alpha\in\bar{\Pi}_{\mathfrak{l}}$ pick up $c_{\alpha}\in\mathbb{C}$ and $u_{\alpha}\in\mathfrak{c}$ assuming $u_{\alpha}\not=0$ only if $\alpha$ is even, orthogonal to $\Pi_{\mathfrak{l}}$, and $\tilde{\alpha}=\alpha=\tau(\alpha)$. Put
$\begin{array}[]{rcl}y_{\alpha}&=&h_{\alpha}-h_{\tilde{\alpha}},\\ x_{\alpha}&=&e_{\alpha}+c_{\alpha}f_{\tilde{\alpha}}+u_{\alpha},\end{array}$ (4.10)
for all $\alpha\in\bar{\Pi}_{\mathfrak{l}}$. Define a Lie subalgebra $\mathfrak{k}\subset\mathfrak{g}$ as the one generated by $\mathfrak{l}+\mathfrak{t}$ and by $x_{\alpha}$, $y_{\alpha}$ with $\alpha\in\bar{\Pi}_{\mathfrak{l}}$.
Definition 4.13.
The pair of Lie superalgebras $\mathfrak{k}\subset\mathfrak{g}$ corresponding to a pseudo-symmetric triple $(\mathfrak{g},\mathfrak{l},\tau)$ is called pseudo-supersymmetric.
Next we prove that pseudo-supersymmetric pairs are super-spherical. Denote by $\mathfrak{p}^{0}=\mathfrak{h}+\mathfrak{l}$ the Levi subalgebra with commutant $\mathfrak{l}$ and set
$\mathfrak{p}_{\pm}^{1}=\sum_{\mu\in\Pi_{\mathfrak{g}/\mathfrak{l}}}V_{\mu}^{\pm}+\mathfrak{p}^{0},\quad\mathfrak{p}_{\pm}^{m+1}=[\mathfrak{p}_{\pm}^{m},\mathfrak{p}^{1}_{\pm}],\quad m>0.$
We have is an $\mathfrak{l}$-equivariant filtration of parabolic subalgebras $\mathfrak{p}_{\pm}$:
$\mathfrak{p}^{0}\subset\mathfrak{p}_{\pm}^{1}\subset\ldots\subset\mathfrak{p}_{\pm}.$
Lemma 4.14.
For all $m,n\geqslant 1$, there is an inclusion $[\mathfrak{p}^{m}_{+},\mathfrak{p}^{n}_{-}]\subset\mathfrak{p}^{m}_{+}+\mathfrak{p}^{n}_{-}$.
Proof.
For $m=n=1$ this readily follows by root arguments. Suppose this is proved for $m=1$ and $n\geqslant 1$. The (graded) Jacobi identity gives
$[\mathfrak{p}^{1}_{+},\mathfrak{p}^{n+1}_{-}]=[\mathfrak{p}^{1}_{+},[\mathfrak{p}^{1}_{-},\mathfrak{p}^{n}_{-}]]\subset[\mathfrak{p}^{1}_{+}+\mathfrak{p}^{1}_{-},\mathfrak{p}^{n}_{-}]+[\mathfrak{p}^{1}_{-},\mathfrak{p}^{1}_{+}+\mathfrak{p}^{n}_{-}]\subset\mathfrak{p}^{1}_{+}+\mathfrak{p}^{n+1}_{-},$
which proves the formula for $m=1$ and all $n$. Now suppose that it is true for $m\geqslant 1$. Then
$[\mathfrak{p}^{m+1}_{+},\mathfrak{p}^{n}_{-}]=[[\mathfrak{p}^{m}_{+},\mathfrak{p}^{1}_{+}],\mathfrak{p}^{n}_{-}]\subset[\mathfrak{p}^{m}_{+}+\mathfrak{p}^{n}_{-},\mathfrak{p}^{1}_{+}]+[\mathfrak{p}^{m}_{+},\mathfrak{p}^{1}_{+}+\mathfrak{p}^{n}_{-}]\subset\mathfrak{p}^{m+1}_{+}+\mathfrak{p}^{n}_{-},$
which completes the proof. ∎
Corollary 4.15.
Let $m_{+}$ be the number of pluses and $m_{-}$ the number of minuses in a sequence $(\varepsilon_{1},\ldots,\varepsilon_{k})$, where $\varepsilon_{i}\in\{\pm\}$. Then
$[\ldots[\mathfrak{p}^{1}_{\varepsilon_{1}},\mathfrak{p}^{1}_{\varepsilon_{2}}],\ldots,\mathfrak{p}^{1}_{\varepsilon_{k}}]\subset\mathfrak{p}^{m_{+}}_{+}+\mathfrak{p}^{m_{-}}_{-}.$
Proof.
For $k=2$, this follows from the definition of $\mathfrak{p}_{\pm}^{m}$ and from Lemma 4.14. Suppose the statement is proved for $k\geqslant 2$. Then, by induction,
$[\ldots[\mathfrak{p}^{1}_{\varepsilon_{1}},\mathfrak{p}^{1}_{\varepsilon_{2}}],\ldots,\mathfrak{p}^{1}_{\varepsilon_{k+1}}]\subset[\mathfrak{p}^{m_{+}}_{+}+\mathfrak{p}^{m_{-}}_{-},\mathfrak{p}^{1}_{\varepsilon_{k+1}}],$
and the statement follows from Lemma 4.14 for general $k$. ∎
Let $X_{\beta}\subset V^{+}_{\beta}\oplus V^{-}_{\beta^{\prime}}$ with $\beta\in\bar{\Pi}_{\mathfrak{l}}$ be the $\mathfrak{l}$-module generated by $x_{\beta}$. We define $\mathfrak{k}^{0}=\mathfrak{t}+\mathfrak{l}$ and
$\mathfrak{k}^{1}=\sum_{\beta\in\Pi_{\mathfrak{g}/\mathfrak{l}}}X_{\beta}+\mathfrak{k}^{0},\quad\mathfrak{k}^{m+1}=[\mathfrak{k}^{m},\mathfrak{k}^{1}],\quad m>0.$
The $\mathfrak{l}$-modules $\mathfrak{k}^{m}$ form an ascending filtration of $\mathfrak{k}$. Furthermore, condition (4.8) implies
$[y_{\alpha},x_{\beta}]=\bigl{(}(\alpha,\beta)-(\tilde{\alpha},\beta)\bigr{)}e_{\beta}+c_{\beta}\bigl{(}-(\alpha,\tilde{\beta})+(\tilde{\alpha},\tilde{\beta})\bigr{)}f_{\tilde{\beta}}\propto x_{\beta}$
for all $\alpha,\beta\in\bar{\Pi}_{\mathfrak{l}}$. The rightmost implication holds because the right-hand side of the equality simply vanishes if $\beta$ is orthogonal to $\Pi_{\mathfrak{l}}$ and $\beta=\tilde{\beta}$. We conclude that the filtration $\mathfrak{k}^{0}\subset\mathfrak{k}^{1}\subset\ldots\subset\mathfrak{k}$ is $\mathfrak{t}$- and therefore $\mathfrak{k}^{0}$-invariant.
Proposition 4.16.
Pseudo-supersymmetric Lie superalgebras are super-spherical.
Proof.
Observe that $\mathfrak{k}^{0}\subset\mathfrak{p}^{0}$ and $\mathfrak{k}^{1}+\mathfrak{p}^{1}_{-}=\mathfrak{p}_{-}^{1}+\mathfrak{p}^{1}_{+}$ by construction. Suppose that we have proved an equality $\mathfrak{k}^{m}+\mathfrak{p}^{m}_{-}=\mathfrak{p}_{-}^{m}+\mathfrak{p}^{m}_{+}$ for $m\geqslant 1$. For any sequence $(x_{\mu_{i}})_{i=1}^{m+1}$ we have
$[\ldots[x_{\mu_{1}},x_{\mu_{2}}],\ldots,x_{\mu_{m+1}}]=[\ldots[e_{\mu_{1}},e_{\mu_{2}}],\ldots,e_{\mu_{m+1}}]+c[\ldots[f_{\tau(\mu_{1})},f_{\tau(\mu_{2})}],\ldots,f_{\tau(\mu_{m+1})}]+\ldots,$
where $c\in\mathbb{C}$ and the omitted terms are in $\mathfrak{p}^{m}_{+}+\mathfrak{p}^{m}_{-}$, by Corollary 4.15. We continue by induction assumption, then the above expression translates to
$=[\ldots[e_{\mu_{1}},e_{\mu_{2}}],\ldots,e_{\mu_{m+1}}]\mod\mathfrak{k}^{m}+\mathfrak{p}^{m+1}_{-}.$
Since the filtration is $\mathfrak{l}$-invariant, and all $[\ldots[e_{\mu_{1}},e_{\mu_{2}}],\ldots,e_{\mu_{m+1}}]$ generate $\mathfrak{p}^{m+1}_{+}\mod\mathfrak{p}^{m}_{+}$ we conclude that $\mathfrak{k}^{m+1}+\mathfrak{p}^{m+1}_{-}\supset\mathfrak{p}^{m+1}_{+}$ and $\mathfrak{k}^{m+1}+\mathfrak{p}^{m+1}_{-}\supset\mathfrak{p}^{m+1}_{+}+\mathfrak{p}^{m+1}_{-}$. This implies
$\mathfrak{k}+\mathfrak{b}_{-}=(\mathfrak{k}+\mathfrak{l})+\mathfrak{b}_{-}=\mathfrak{k}+(\mathfrak{l}+\mathfrak{b}_{-})=\mathfrak{k}+\mathfrak{p}_{-}\supset\mathfrak{p}_{+}+\mathfrak{p}_{-}=\mathfrak{g}.$
The proof is complete. ∎
In the next section we answer the question when the subalgebra $\mathfrak{k}\subset\mathfrak{g}$ is proper for a given pseudo-supersymmetric triple, in the special symmetric grading we have fixed.
4.1 Decorated diagrams and selection rules
Like in the non-graded case the permutation $\tau$ entering a pseudo-supersymmetric triple $(\mathfrak{g},\mathfrak{l},\tau)$ is an involutive even automorphism of the Dynkin diagram coinciding with $-w_{\mathfrak{l}}$ on $\Pi_{\mathfrak{l}}$. Such triples can be visualized via decorated Dynkin diagrams similar to pseudo-symmetric triples from [4].
Some diagrams lead to trivial subalgebras $\mathfrak{k}$ that exhaust all of $\mathfrak{g}$. In the non-graded case, there is a selection rule that reads as follows. Let $C(\alpha)$ denote the union of connected components of $\Pi_{\mathfrak{l}}$ that are connected to $\{\alpha,\tau(\alpha)\}$. Then diagrams with
[…]
should be ruled out. In this section we work out selection rules that we impose on graded decorated diagrams. In the case of basic algebras, the remaining diagrams describe non-trivial pairs $(\mathfrak{g},\mathfrak{k})$, which correspond to solutions of RE listed in Section 3. We assume throughout this section that all elements $u_{\alpha}$ in the definition (4.10) of the subalgebra $\mathfrak{k}$ are set to zero.
Graded selection rules are more intricate than non-graded. They amount to a set lemmas and are based on the same strategy: assuming that $\alpha\in\bar{\Pi}_{\mathfrak{l}}$ to show that the Lie superalgebra $\mathfrak{g}^{(\alpha)}=\mathrm{Span}\{e_{\alpha},f_{\alpha},h_{\alpha}\}$ is in $\mathfrak{k}$. Let us demonstrate that the only non-graded selection rule fails to be definite in the super case.
Lemma 4.17.
Suppose that the diagram is such that (4.11) holds for some $\alpha\in\bar{\Pi}_{l}$. Then $\mathfrak{g}^{(\alpha)}\subset\mathfrak{k}$ unless $\alpha$ is odd and $\beta$ is even.
Proof.
Suppose that the condition (4.11) is fulfilled with some $\alpha\in\bar{\Pi}_{\mathfrak{l}}$ and $\beta\in\Pi_{\mathfrak{l}}$. First of all observe that $\beta$ cannot be odd, because otherwise $w_{\mathfrak{l}}$ will not be even, by Proposition 4.6 (that will also violate the condition (4.7)).
Furthermore, for even $\beta$, we have $\tilde{\alpha}=\alpha+\beta$. With $X_{\alpha}=e_{\alpha}+c_{\alpha}[f_{\beta},f_{\alpha}]\in\mathfrak{k}$, we define
$X_{\tilde{\alpha}}=[e_{\beta},X_{\alpha}]=[e_{\beta},e_{\alpha}]+c_{\alpha}[h_{\beta},f_{\alpha}]=[e_{\beta},e_{\alpha}]+c_{\alpha}f_{\alpha}\in\mathfrak{k}.$
Then
$[X_{\alpha},X_{\tilde{\alpha}}]=[e_{\alpha}+c_{\alpha}[f_{\beta},f_{\alpha}],[e_{\beta},e_{\alpha}]+c_{\alpha}f_{\alpha}]=c_{\alpha}(h_{\alpha}+(-1)^{|\alpha|}h_{\alpha}+(-1)^{|\alpha|}h_{\beta}).$
Since $h_{\beta}\in\mathfrak{k}$, we conclude that $h_{\alpha}\in\mathfrak{k}$ and therefore $e_{\alpha},f_{\alpha}\in\mathfrak{k}$ only if $\alpha$ is even. In that case, the set $\Pi_{\mathfrak{l}}$ can be extended to comprise $\alpha$. ∎
It follows from the proof of this lemma that the diagram $\alpha$$\beta$ with even $\beta$ and odd $\alpha$ gives rise to a 5-dimensional super-spherical subalgebra $\mathfrak{k}=\mathrm{Span}\{e_{\beta},h_{\beta},f_{\beta},X_{\alpha},X_{\tilde{\alpha}}\}$ in $\mathfrak{g}=\mathfrak{g}\mathfrak{l}(2|1)$. This $\mathfrak{k}$ is beyond the scope of this study because this grading of $\mathfrak{g}\mathfrak{l}(2|1)$ is not symmetric. As a subdiagram, it nevertheless appears, for instance, in $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|\mathbf{m})$ in relation with the RE matrix $C$.
Similarly to shape of graded Dynkin diagrams we consider shapes of decorated diagrams by discarding grading. In other words, a graded decorated diagram is obtained from its shape by adding a grading datum.
Proposition 4.18.
Let $\mathcal{D}$ be a decoration of a Dynkin diagram $D$. Then the shape of $\mathcal{D}$ is a decorated shape of $D$.
Proof.
Notice that the automorphism $\tau$ of a graded diagram is an automorphism of its shape. ∎
Proposition 4.6 states that a grey odd root cannot comprise a connected component of $\Pi_{\mathfrak{l}}$. It turns out that such a node cannot be isolated of other nodes from $\Pi_{\mathfrak{l}}$ in all cases.
Lemma 4.19.
Suppose that a gray node $\alpha\in\bar{\Pi}_{\mathfrak{l}}$ is isolated from $\Pi_{\mathfrak{l}}$ and $\tau(\alpha)=\alpha$. Suppose that $\beta\in\bar{\Pi}_{\mathfrak{l}}$ is such that $(\alpha,\beta)\not=0$. Then $\mathfrak{g}^{(\alpha)}+\mathfrak{g}^{(\beta)}\subset\mathfrak{k}$.
Proof.
Indeed, we conclude that $h_{\alpha}\in\mathfrak{k}$ because
$[X_{\alpha},X_{\alpha}]=X_{\alpha}X_{\alpha}+X_{\alpha}X_{\alpha}=[e_{\alpha},c_{\alpha}f_{\alpha}]+[c_{\alpha}f_{\alpha},e_{\alpha}]=2c_{\alpha}h_{\alpha}.$
Observe that $(\alpha,\tilde{\beta})=\bigl{(}\alpha,\tau(\beta)+\ldots\bigr{)}=\bigl{(}\tau(\alpha),\tau(\beta)\bigr{)}=(\alpha,\beta)$, where we have suppressed the terms from $\mathbb{Z}\Pi_{\mathfrak{l}}$. Commuting $h_{\alpha}$ with $X_{\beta}=e_{\beta}+c_{\beta}f_{\tilde{\beta}}$ we find
$[h_{\alpha},X_{\beta}]=(\alpha,\beta)(e_{\beta}-c_{\beta}f_{\tilde{\beta}})\in\mathfrak{k}.$
Thus we find that $e_{\beta}\in\mathfrak{k}$, $f_{\tilde{\beta}}\in\mathfrak{k}$, and therefore $f_{\beta},h_{\beta}\in\mathfrak{k}$. Finally, $h_{\beta}$ splits $X_{\alpha}$, so $e_{\alpha},f_{\alpha}\in\mathfrak{k}$ either. ∎
Corollary 4.20.
Suppose that $m>\mathbf{m}$. Then the subalgebra $\mathfrak{k}$ corresponding to a diagram of type $A$ exhausts all of ortho-symplectic Lie superalgebra $\mathfrak{g}$.
Lemma 4.21.
Suppose that decorated Dynkin diagram contains a subdiagram
[…]
$\alpha$$\beta$
$\gamma$$\sigma$
where $\alpha$ and $\gamma$ are even, $\beta$ is odd, and $\gamma$ with $\sigma$ form an arbitrary connected even diagram of rank 2. Then $\mathfrak{g}^{(\beta)}+\mathfrak{g}^{(\sigma)}\subset\mathfrak{k}$.
Proof.
By the assumption, $\tau(\beta)=\beta$ and $\tau(\sigma)=\sigma$. Choose the grading such that $(\alpha,\beta)=1=-(\beta,\gamma)$. Evaluating commutators of $e_{\alpha}$ and $e_{\gamma}$ with $X_{\beta}=e_{\beta}+c_{\beta}[[f_{\alpha},f_{\beta}],f_{\gamma}]\in\mathfrak{k}$, we obtain $X_{\tilde{\beta}}=c_{\beta}f_{\beta}-[[e_{\alpha},e_{\beta}],e_{\gamma}]\in\mathfrak{k}$. Then $[X_{\beta},X_{\tilde{\beta}}]=c_{\beta}(2h_{\beta}+h_{\alpha}+h_{\gamma})=c_{\beta}h_{\delta}\in\mathfrak{k}.$ Since
$(\delta,\sigma)=(\gamma,\sigma)\not=-(\delta,\tilde{\sigma})=-(2\beta+\gamma,\sigma-\frac{2(\gamma,\sigma)}{(\gamma,\gamma)}\gamma+\ldots)=-2\frac{2(\gamma,\sigma)}{(\gamma,\gamma)}+(\gamma,\sigma),$
the element $h_{\delta}$ splits $X_{\sigma}=e_{\sigma}+c_{\sigma}f_{\tilde{\sigma}}$ and therefore $X_{\tilde{\sigma}}\propto f_{\sigma}+c_{-\sigma}e_{\tilde{\sigma}}\in\mathfrak{k}$ with some $c_{-\sigma}\in\mathbb{C}^{\times}$. Hence $e_{\sigma},f_{\sigma}$, and $h_{\sigma}\in\mathfrak{k}$. The element $h_{\sigma}$ splits $X_{\beta}$ and $X_{\tilde{\beta}}$ because $(\sigma,\beta)=0$ and $-(\sigma,\tilde{\beta})=-(\sigma,\gamma)\neq 0$. Therefore $e_{\beta},f_{\beta},h_{\beta}\in\mathfrak{k}$ either. ∎
Remark 4.22.
There are also subdiagrams related to $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m})$ of rank $n$, which are
$\alpha$$\beta$$\gamma$$\sigma=\alpha_{n-1}=\alpha_{m}$$\alpha_{n}=\sigma^{\prime}$$\alpha$$\beta$$\gamma$$\sigma=\alpha_{n}=\alpha_{m}$
where $\alpha$, $\gamma$, and $\sigma$ are even, $\beta$ is odd. We can achieve an identical result to Lemma 4.21 through a similar line of reasoning.
Proof.
The proof is similar to Lemma 4.21, so we only indicate the differences. By the assumption, $\tau(\beta)=\beta$ and $\tau(\sigma)=\sigma^{\prime}$ for left diagram and $\tau(\sigma)=\sigma$ for the right one. As before, we find that $h_{\delta}$ with $\delta=2\beta+\alpha+\gamma$ is in $\mathfrak{k}$. Further we find that
$(\delta,\sigma)=(\gamma,\sigma)=-1\not=-(\delta,\tilde{\sigma})=-(2\beta+\gamma,\alpha_{n}+\gamma)=1\ \mathrm{(for\ left\ diagram)},$
$(\delta,\sigma)=(2\beta,\sigma)=-2\not=-(\delta,\tilde{\sigma})=-(2\beta,\alpha_{n})=2\ \mathrm{(for\ right\ diagram)}.$
As a consequence, the element $h_{\delta}$ splits $X_{\sigma}$ and $X_{\tilde{\sigma}}$. The rest of the proof is literally the same as in Lemma 4.21. ∎
Corollary 4.23.
Suppose that $m>\mathbf{m}$. Then the subalgebra $\mathfrak{k}$ corresponding to a diagram of type $B$ exhausts all of ortho-symplectic Lie superalgebra $\mathfrak{g}$.
Consider decorated Dynkin diagrams
of $\mathfrak{g}\mathfrak{l}$-type and
of $\mathfrak{o}\mathfrak{s}\mathfrak{p}$ and $\mathfrak{s}\mathfrak{p}\mathfrak{o}$ types, where the big circles on the right stand for even tail, the black nodes designate roots from $\Pi_{\mathfrak{l}}$, and the square nodes are odd. The diagrams enclosed in the brackets mean the period (including zero occurrence for ()()( ) and at least one for [][][ ]).
Conjecture 1.
The diagrams above give rise to proper subalgebras $\mathfrak{k}\subset\mathfrak{g}$.
It has been checked for several simplest cases.
Remark that diagrams
with two odd grey nodes corresponding to $\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|4)$ yield non-trivial pairs $(\mathfrak{g},\mathfrak{k})$. These diagrams can be extended to the left with alternating white and black nodes.
Examination of possible decorated Dynkin diagrams against the selection rules of this section leads to a classification of proper super-spherical triples.
Proposition 4.24.
All proper super-spherical subalgebras $\mathfrak{k}\subset\mathfrak{g}$ in the minimal symmetric grading except for those from Conjecture 1 are related with solutions to RE of type $A$, $B$, and $C$.
Proof.
A direct examination. ∎
We conjecture that pairs $(\mathfrak{g},\mathfrak{k})$ corresponding to diagrams from Conjecture 1 have no non-trivial K-matrices.
4.2 Relation to K-matrices
Consider matrices $A,B,C$ from Theorem 3.1, and the invertible matrix $A$ from (3.3). Denote their classical limit ($q\rightarrow 1$) by $A_{0}$, $B_{0}$, and $C_{0}$ ($=C$ as it is independent of $q$). As in the non-graded case, we denote by $\theta$ the map $-w_{\mathfrak{l}}\circ\tau$, which takes $\alpha\in\bar{\Pi}_{\mathfrak{l}}$ to $-\tilde{\alpha}\in\mathrm{R}^{-}_{\mathfrak{g}}$. Let $K\subset G$ be the centralizer subgroup of $A_{0}$, $B_{0}$, and $C_{0}$ with Lie superalgebra $\mathfrak{k}$.
We will describe the action of $\theta$ on the basis of $\mathfrak{h}^{*}$, as follows (recall that $n$ stands for the rank of $\mathfrak{g}$). In the case of matrix $A_{0}$ it is
$\theta(\zeta_{i})=\begin{cases}\zeta_{i^{\prime}},&\mathrm{if}\ i\leqslant m,\\ \zeta_{i},&\mathrm{if}\ m<i,\\ \end{cases},\quad\zeta_{i}=\delta_{i}\ \mathrm{or}\ \varepsilon_{i},$
for $\mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$ and
$\theta(\delta_{i})=\begin{cases}-\delta_{i},&\mathrm{if}\ i\leqslant m,\\ \delta_{i},&\mathrm{if}\ i>m,\\ \end{cases},\quad\theta(\varepsilon_{i})=\varepsilon_{i}.$
for ortho-symplectic $\mathfrak{g}$. In the case of matrix $B_{0}$ for ortho-symplectic $\mathfrak{g}$ it is
$\theta(\delta_{i})=\begin{cases}-\delta_{i+1},&\mathrm{if}\ i<m\ \mathrm{is\ odd},\\ -\delta_{i-1},&\mathrm{if}\ i\leqslant m\ \mathrm{is\ even},\\ \delta_{i},&\mathrm{if}\ i>m,\\ \end{cases}\quad\theta(\varepsilon_{i})=\varepsilon_{i}.$
In the case of matrix $C_{0}$ for ortho-symplectic $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|4\mathbf{m})$ it is
$\theta(\delta_{i})=\begin{cases}-\delta_{i+1},&\mathrm{if}\ i<n\ \mathrm{is\ odd},\\ -\delta_{i-1},&\mathrm{if}\ i\leqslant n-1\ \mathrm{is\ even},\\ \end{cases}\quad\theta(\varepsilon_{1})=-\varepsilon_{1}.$
The root basis of the subalgebra $\mathfrak{l}$ is explicitly
$\Pi_{\mathfrak{l}}=\begin{cases}\{\alpha_{i}\}_{i=m+1}^{n-m},&\mbox{for }\mathfrak{g}\mathfrak{l}(N|2\mathbf{m}),\\ \{\alpha_{i}\}_{i=m+1}^{n},&\mbox{for }\mathfrak{o}\mathfrak{s}\mathfrak{p}(N|2\mathbf{m}),\ \mbox{and}\ \mathfrak{s}\mathfrak{p}\mathfrak{o}(N|2\mathbf{m})\ \mbox{related\ to}\ A_{0},\\ \{\alpha_{2i+1}\}_{i=0}^{\frac{m}{2}-1}\cup\{\alpha_{i}\}_{i=m+1}^{n},&\mbox{for }\mathfrak{o}\mathfrak{s}\mathfrak{p}(N|2\mathbf{m}),\ \mbox{and}\ \mathfrak{s}\mathfrak{p}\mathfrak{o}(N|2\mathbf{m})\ \mbox{related\ to}\ B_{0}.\\ \{\alpha_{2i+1}\}_{i=0}^{\frac{n-3}{2}},&\mbox{for }\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|4\mathbf{m}),\ \mbox{related\ to}\ C_{0}.\\ \end{cases}$
For simple roots $\alpha\in\bar{\Pi}_{\mathfrak{l}}$ in the case of $A_{0}$ the roots $\tilde{\alpha}$ are given by
$\tilde{\alpha}_{i}=\alpha_{i^{\prime}-1},\quad i=1,\ldots,m-1,\quad i=m=\frac{N+2\mathbf{m}}{2},\quad\tilde{\alpha}_{i}=\alpha_{i^{\prime}-1},\quad i=m^{\prime}+1,\ldots,n,$
$\tilde{\alpha}_{m}=\sum_{l=m+1}^{n-m+1}\alpha_{l},\quad\tilde{\alpha}_{n-m+1}=\sum_{l=m}^{n-m}\alpha_{l},$
for $\mathfrak{g}=\mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$,
$\tilde{\alpha}_{i}=\alpha_{i},\quad i=1,\ldots,m-1,\quad\tilde{\alpha}_{m}=\alpha_{m}+2\sum_{l=m+1}^{n}\alpha_{l},$
for $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m})$,
$\tilde{\alpha}_{i}=\alpha_{i},\quad i=1,\ldots,m-1,\quad\tilde{\alpha}_{m}=\alpha_{m}+2\sum_{l=m+1}^{n-2}\alpha_{l}+\alpha_{n-1}+\alpha_{n},\quad m<n-1,$
$\tilde{\alpha}_{n-1}=\alpha_{n},\quad\tilde{\alpha}_{n}=\alpha_{n-1},\quad m=n-1,$
for $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m})$,
$\tilde{\alpha}_{i}=\alpha_{i},\quad i=1,\ldots,m-1,\quad\tilde{\alpha}_{m}=\alpha_{m}+2\sum_{l=m+1}^{n-1}\alpha_{l}+\alpha_{n},\quad m<n-1,$
$\tilde{\alpha}_{m}=\alpha_{n-1}+\alpha_{n},\quad m=n-1,$
for $\mathfrak{g}=\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m})$.
For simple roots $\alpha\in\bar{\Pi}_{\mathfrak{l}}$, in the case of $B_{0}$ the roots $\tilde{\alpha}$ are given by
$\tilde{\alpha}_{2i}=\sum_{l=2i-1}^{2i+1}\alpha_{l},\quad i=1,\ldots,\frac{m}{2}-1,\quad\tilde{\alpha}_{m}=\alpha_{m-1}+\alpha_{m}+2\sum_{l=m+1}^{n}\alpha_{l},$
for $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m})$,
$\tilde{\alpha}_{2i}=\sum_{l=2i-1}^{2i+1}\alpha_{l},\quad i=1,\ldots,\frac{m}{2}-1,\quad\tilde{\alpha}_{m}=\alpha_{m-1}+\alpha_{m}+2\sum_{l=m+1}^{n-2}\alpha_{l}+\alpha_{n-1}+\alpha_{n},$
$\tilde{\alpha}_{n-1}=\alpha_{n-2}+\alpha_{n},\quad\tilde{\alpha}_{n}=\alpha_{n-2}+\alpha_{n-1},\quad m=n-1,$
for $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m})$,
$\tilde{\alpha}_{2i}=\sum_{l=2i-1}^{2i+1}\alpha_{l},\quad i=1,\ldots,\frac{m}{2}-1,\quad\tilde{\alpha}_{m}=\alpha_{m-1}+\alpha_{m}+2\sum_{l=m+1}^{n-1}\alpha_{l}+\alpha_{n},$
for $\mathfrak{g}=\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m})$.
For simple roots $\alpha\in\bar{\Pi}_{\mathfrak{l}}$, in the case of $C_{0}$ the roots $\tilde{\alpha}$ are given by
$\tilde{\alpha}_{2i}=\sum_{l=2i-1}^{2i+1}\alpha_{l},\quad i=1,\ldots,\frac{n-3}{2},$
$\tilde{\alpha}_{n-1}=\alpha_{n-2}+\alpha_{n-1},\quad\tilde{\alpha}_{n}=\alpha_{n-2}+\alpha_{n},$
for $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|4\mathbf{m})$.
The Lie superalgebra $\mathfrak{k}$ is generated by $\mathfrak{l}$ and additional elements $X_{\alpha}=e_{\alpha}+c_{\alpha}f_{\tilde{\alpha}}+\grave{c}_{\alpha}h_{\alpha}$, $h_{\tilde{\alpha}}-h_{\alpha}$, where $\alpha\in\bar{\Pi}_{\mathfrak{l}}$, and $c_{\alpha}\in\mathbb{C}^{\times}$, $\grave{c}_{\alpha}\in\mathbb{C}$. The scalar $\grave{c}_{\alpha}\neq 0$ only if $\mathfrak{g}=\mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$, and $\alpha=\alpha_{\frac{N+2\mathbf{m}}{2}}$.
Consider the non-graded shape of the graded Dynkin diagram of $\mathfrak{g}$. It turns out that the pseudo-supersymmetric triples associated with graded RE-matrices correspond to pseudo-symmetric triples of the shape diagrams subject to the condition on the position of the odd node $m\leqslant\mathbf{m}$, for orthosymplectic $\mathfrak{g}$. Diagrams with $m>\mathbf{m}$ are forbidden by Corollaries 4.20 and 4.23. There are no such restrictions for general linear $\mathfrak{g}$.
5 Coideal subalgebras
Recall that a total order on the set of positive roots $\mathrm{R}^{+}\supset\Pi$ of $\mathfrak{g}$ is called normal if every sum $\alpha+\beta\in\mathrm{R}^{+}$ with $\alpha,\beta\in\mathrm{R}^{+}$ is between $\alpha$ and $\beta$.
Choose a normal order on $\mathrm{R}^{+}$ and extend $F_{\beta}$ to all $\beta\in\mathrm{R}^{+}$ as $a_{\beta}e_{-\beta}$ in [21], where $a_{\beta}$ is a non-zero complex number. Denote by $\mathcal{U}_{m}^{-}$ the subalgebra in $\mathcal{U}^{-}$ generated by $F_{\beta^{i}}$ with $i\leqslant m$.
Lemma 5.1.
For each for each $m=1,\ldots,|\mathrm{R}^{+}|$, $\mathcal{U}_{m}^{-}$ is a left coideal subalgebra in $U_{q}(\mathfrak{g})$. Furthermore,
$\Delta(F_{\beta^{m}})=(F_{\beta^{m}}\otimes 1+q^{h_{\beta^{m}}}\otimes F_{\beta^{m}})+U_{q}(\mathfrak{g})\otimes\mathcal{U}^{-}_{m-1}.$
Proof.
This readily follows from [21], Prop. 8.3. upon the assignment $F_{\alpha}\mapsto e_{-\alpha}$ on simple root vectors, extended as an algebra automorphism to $U_{q}(\mathfrak{b}_{-})$. It is then a coalgebra anti-isomorphism. ∎
We conclude, by Lemma 5.1, that for each $\alpha\in\bar{\Pi}_{\mathfrak{l}}$ the elements $F_{\tilde{\alpha}}$ and $F_{\mu}$ with $\mu\in\Pi_{\mathfrak{l}_{\tilde{\alpha}}}$ form a left coideal subalgebra in $U_{q}(\mathfrak{g})$.
A subalgebra $U_{q}(\mathfrak{k})\subset U_{q}(\mathfrak{g})$ is generated over $U_{q}(\mathfrak{l})$ by
$X_{\alpha}=q^{h_{\tilde{\alpha}}-h_{\alpha}}e_{\alpha}+c_{\alpha}F_{\tilde{\alpha}}+\grave{c}_{\alpha}(q^{h_{\alpha}}-1),$
and by $q^{\pm(h_{\tilde{\alpha}}-h_{\alpha})}$ with $\alpha\in\bar{\Pi}_{\mathfrak{l}}$ and $c_{\alpha}\neq 0,\grave{c}_{\alpha}\in\mathbb{C}$. The constant $\grave{c}_{\alpha}$ is distinct from zero only if $\mathfrak{g}=\mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$ with $\alpha=\tilde{\alpha}_{m}=\alpha_{m},\ m=\frac{N+2\mathbf{m}}{2}$, and $\mu\neq-\lambda$.
Theorem 5.2.
The algebra generated by $e_{\alpha},f_{\alpha},q^{\pm h_{\alpha}}$ with $\alpha\in\Pi_{\mathfrak{l}}$, and $X_{\alpha}=q^{h_{\tilde{\alpha}}-h_{\alpha}}e_{\alpha}+c_{\alpha}F_{\tilde{\alpha}}+\grave{c}_{\alpha}(q^{h_{\alpha}}-1)$, $q^{\pm(h_{\tilde{\alpha}}-h_{\alpha})}$ with $\alpha\in\bar{\Pi}_{\mathfrak{l}}$, is a left coideal in $U_{q}(\mathfrak{g})$.
Proof.
First set $\grave{c}_{\alpha}=0$. By Lemma 5.1, one has for each $\alpha\in\bar{\Pi}_{\mathfrak{l}}$:
$\Delta(F_{\tilde{\alpha}})\in(q^{h_{\tilde{\alpha}}}\otimes F_{\tilde{\alpha}}+F_{\tilde{\alpha}}\otimes 1)+U_{q}(\mathfrak{g})\otimes U_{q}(\mathfrak{l}_{\tilde{\alpha}}),$
$\Delta(q^{h_{\tilde{\alpha}}-h_{\alpha}}e_{\alpha})=q^{h_{\tilde{\alpha}}}\otimes q^{h_{\tilde{\alpha}}-h_{\alpha}}e_{\alpha}+q^{h_{\tilde{\alpha}}-h_{\alpha}}e_{\alpha}\otimes q^{h_{\tilde{\alpha}}-h_{\alpha}}.$
Adding the lines together one gets
$\Delta(X_{\tilde{\alpha}})\in q^{h_{\tilde{\alpha}}}\otimes X_{\alpha}+F_{\tilde{\alpha}}\otimes 1+q^{h_{\tilde{\alpha}}-h_{\alpha}}e_{\alpha}\otimes q^{h_{\alpha}-h_{\tilde{\alpha}}}+U_{q}(\mathfrak{g})\otimes U_{q}(\mathfrak{l}_{\tilde{\alpha}}),$
as required. It is straightforward to see that the coproduct of the term with $\grave{c}_{\alpha}\neq 0$ is in $U_{q}(\mathfrak{g})\otimes U_{q}(\mathfrak{k})$ too. This completes the proof. ∎
We apply this theorem to quantize $U(\mathfrak{k})$. Pick up $\alpha\in\bar{\Pi}_{\mathfrak{l}}$, put $\alpha^{\prime}=\tau(\alpha)\in\bar{\Pi}_{\mathfrak{l}}$ and $\tilde{\alpha}=w_{\mathfrak{l}}\circ(\alpha^{\prime})\in\mathrm{R}^{+}$. Consider the root system generated by $\alpha^{\prime}$ and $\Pi_{\mathfrak{l}}$; denote by $\Pi_{\mathfrak{k}_{\tilde{\alpha}}}$ its connected component containing $\alpha^{\prime}$ and by $\mathfrak{k}_{\tilde{\alpha}}$ the subalgebra generated by the corresponding simple root vectors. Denote also $\Pi_{\mathfrak{l}_{\tilde{\alpha}}}=\Pi_{\mathfrak{k}_{\tilde{\alpha}}}\cap\Pi_{\mathfrak{l}}$ and by $\mathfrak{l}_{\tilde{\alpha}}$ the corresponding subalgebra in $\mathfrak{l}$.
Proposition 5.3.
For each $\alpha\in\Pi_{\alpha}$ there is a normal order on $\mathrm{R}^{+}_{\mathfrak{k}_{\tilde{\alpha}}}$ such that
  • $\alpha^{\prime}$ is in the rightmost position and all $\mathrm{R}^{+}_{\mathfrak{l}_{\tilde{\alpha}}}$ are on the left.
  • the root $\tilde{\alpha}$ is next to the right after $\mathrm{R}^{+}_{\mathfrak{l}_{\tilde{\alpha}}}$.
Proof.
In the non-graded case this is derived from the properties of the longest element of the Weyl group. In our case this is checked by a direct examination. ∎
We conclude that the algebra $U(\mathfrak{k})$ can be quantized as a left coideal subalgebra $U_{q}(\mathfrak{k})$. It is generated by $e_{\alpha}$, $f_{\alpha}$, $q^{\pm h_{\alpha}}$ with $\alpha\in\Pi_{\mathfrak{l}}$, and by $X_{\alpha}=q^{h_{\tilde{\alpha}}-h_{\alpha}}e_{\alpha}+c_{\alpha}F_{\tilde{\alpha}}+\grave{c}_{\alpha}(q^{h_{\alpha}}-1)$, $q^{\pm(h_{\tilde{\alpha}}-h_{\alpha})}$, where $\alpha\in\bar{\Pi}_{\mathfrak{l}}$. The elements $F_{\alpha}=q^{h_{\alpha}}f_{\alpha}$ for $\alpha\in\Pi$ are introduced above.
To enhance the readability of this presentation, super-generalized Satake-like diagrams are further used. To avoid a conflict with the conventional notation of graded Dynkin diagrams, we do not use coloured nodes for odd roots assuming that the number of such nodes and their nature have been fixed for each type of $\mathfrak{g}$ by the minimal symmetric grading of the underlying vector space. Instead, we consider families of decorated Dynkin diagrams of the same shape, where possible positions of odd roots are indicated with square nodes.
There are two symmetrically allocated odd simple roots for $\mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$ and $\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|2\mathbf{m})$, and one such root otherwise. An arc connects the root $\alpha^{\prime}$ and $\alpha$ from $\bar{\Pi}_{\mathfrak{l}}$ if $\alpha^{\prime}$ is distinct from $\alpha$. Black nodes depict simple roots from $\Pi_{\mathfrak{l}}$, whereas white nodes label simple roots from $\bar{\Pi}_{\mathfrak{l}}$.
It can be seen that the matrix $A$ (respectively $B$ and $C$) commutes with $\pi\bigl{(}U_{q}(\mathfrak{l})\bigr{)}$ and all $\pi(q^{h_{\tilde{\alpha}}-h_{\alpha}})$, where $\alpha\in\bar{\Pi}_{\mathfrak{l}}$, while the values of $c_{\alpha}$ and $\grave{c}_{\alpha}$ are to be determined. This will be done by requiring that $A$ (respectively $B$ and $C$) commutes with all $\pi(X_{\alpha})$, $\alpha\in\bar{\Pi}_{\mathfrak{l}}$. Expressions for $F_{\tilde{\alpha}}$ with non-simple $\tilde{\alpha}$ will be provided explicitly in terms of q-commutator defined as
$[F_{\alpha},F_{\beta}]_{q^{\pm i}}=F_{\alpha}F_{\beta}-(-1)^{(|F_{\alpha}||F_{\beta}|)}q^{\mp\lceil\frac{i}{2}\rceil(\alpha,\beta)}F_{\beta}F_{\alpha},\quad\forall\alpha,\beta\in\mathrm{R}^{+}.$
Here $i=0,1,2$ and $i\mapsto\lceil\frac{i}{2}\rceil$ sends it to $0,1,1$.
5.1 Diagrams of type $A$
It is found that $\tilde{\alpha}_{i}=\alpha_{i}$ for all $i<m$, with $c_{\alpha_{i}}=-\frac{y_{i+1}}{qy_{i}}$, for ortho-symplectic $\mathfrak{g}$. The four types of $\mathfrak{g}$ will be further treated separately.
$\bullet\quad\mathfrak{g}=\mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$
Its conjugacy classes are expressed by two diagrams
[…]
with the mixture parameters
$c_{\alpha_{i}}=(-1)^{\delta_{i}^{\mathbf{m}^{\prime}}}\frac{y_{i+1}}{y_{i}},\quad\mathrm{if}\quad i<m\quad\mathrm{or}\quad i>N+2\mathbf{m}-m,$
the roots $\tilde{\alpha}_{m}$ and $\tilde{\alpha}_{N+2\mathbf{m}-m}$ are not simple when $m<\frac{N+2\mathbf{m}}{2}$. The corresponding root vectors can be defined as
$F_{\tilde{\alpha}_{m}}=[\dots[F_{m+1},F_{m+2}]_{q},\dots F_{n-m+1}]_{q},\quad F_{\tilde{\alpha}_{n-m+1}}=[\dots[F_{m},F_{m+1}]_{\bar{q}},\dots F_{n-m}]_{\bar{q}}.$
The coefficients are expressed by
$c_{\alpha_{m}}=\frac{(-1)^{N}\mu}{y_{m}},\quad c_{\alpha_{N+2\mathbf{m}-m}}=\begin{cases}\frac{(-1)^{N+1+\delta_{m}^{\mathbf{m}}}q^{2(N+2\mathbf{m})-4m-3}\lambda}{y_{m}},&\mbox{if }\ \mathbf{m}\leq m,\\ \frac{(-1)^{N+1}q^{2(N-2\mathbf{m})+4m+3}\lambda}{y_{m}},&\mbox{if }\ m<\mathbf{m}.\end{cases}$
The case $m=\frac{N+2\mathbf{m}}{2}$ with $\mathbf{m}<m$ is described by the diagram on the right. The simple root $\alpha_{m}=\alpha$ corresponds to the following mixed vector
$X_{\alpha}=e_{\alpha}+c_{\alpha}F_{\alpha}+\grave{c}_{\alpha}(q^{h_{\alpha}}-1),$
where
$c_{\alpha}=\frac{-\lambda\mu q}{y_{m}^{2}},\quad\grave{c}_{\alpha}=\frac{(\mu+\lambda)q}{(q^{2}-1)y_{m}}.$
$\bullet\quad\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m})$
The corresponding diagrams are:
The root vectors are
$F_{\tilde{\alpha}_{m}}=[\dots[[[\dots[[F_{m},F_{m+1}]_{q},F_{m+2}]_{q},\dots F_{n}]_{q},F_{n}],F_{n-1}]_{q},\dots F_{m+1}]_{q},\ m<n-1,$
$F_{\tilde{\alpha}_{m}}=[[F_{n-1},F_{n}]_{q},F_{n}],\quad m=n-1,$
with mixture coefficient
$c_{\alpha_{m}}=\frac{(-1)^{(n-m-\delta_{m}^{\mathbf{m}})}\bar{q}^{2\delta_{m}^{\mathbf{m}}}\lambda}{y_{m}}.$
In the scenario where $\mathbf{m}=m=n$, as depicted in the second diagram, the constant equals
$c_{\alpha_{m}}=-\frac{\lambda}{y_{m}q}.$
$\bullet\quad\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m})$
The case $m\leqslant n-2$ includes two diagrams
$\ldots$ $\ldots$ $\alpha_{1}$$\alpha_{m}$$\alpha_{n-1}$$\alpha_{n}$$\mathbf{n}\neq 1$
$\ldots$ $\ldots$   $\alpha_{1}$$\alpha_{m}$$\alpha_{n-1}$$\alpha_{n}$$\mathbf{n}=1$
The root vectors are
$F_{\tilde{\alpha}_{m}}=[\dots[[\dots[[F_{m},F_{m+1}]_{q},F_{m+2}]_{q},\dots F_{n}]_{q},F_{n-2}]_{q},\dots F_{m+1}]_{q},\quad\mbox{for}\quad m<n-2,$
$F_{\tilde{\alpha}_{m}}=[[F_{n-2},F_{n-1}]_{q},F_{n}]_{q}\quad\mbox{for}\quad m=n-2.$
In both cases, the parameters are
$c_{\alpha_{m}}=\frac{(-1)^{(n-m+1+\delta_{m}^{\mathbf{m}})}\bar{q}^{2\delta_{m}^{\mathbf{m}}}\lambda}{y_{m}}.$
The case $m=n-1$ corresponds to the diagram
[…]
$\bullet\quad\mathfrak{g}=\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m})$
The diagram is
$\ldots$ $\ldots$ $<$$\alpha_{1}$$\alpha_{m}$$\alpha_{n}$
with coefficients
$c_{\alpha_{m}}=(-1)^{n-m+\delta_{m}^{\mathbf{m}}}\frac{\bar{q}^{2\delta_{m}^{\mathbf{m}}}\lambda}{y_{m}},$
and root vectors
$F_{\tilde{\alpha}_{m}}=[\dots[F_{m},F_{m+1}]_{q},F_{m+2}]_{q},\dots F_{n}]_{q^{2}},F_{n-1}]_{q},\dots F_{m+1}]_{q},\quad m<n-1,$
$F_{\tilde{\alpha}_{m}}=[F_{n-1},F_{n}]_{q^{2}},\quad m=n-1.$
The non-common description of the ortho-symplectic type will be provided below.
5.2 Diagrams of type $B$
For all cases, we find
$F_{\tilde{\alpha}_{2i}}=[[F_{2i},F_{2i+1}]_{q},F_{2i-1}]_{q}\ \mathrm{with}\ i=1,\ldots,\frac{m}{2}-1,\quad\mbox{and }$
[…]
$\bullet\quad\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m})$
The Satake diagram is
$\ldots$ $\ldots$ $>$ $\alpha_{1}$$\alpha_{m}$$\alpha_{n}$
Regarding $2i=m$, they are
$F_{\tilde{\alpha}_{m}}=[\dots[[[\dots[[F_{m},F_{m+1}]_{q},F_{m+2}]_{q},\dots F_{n}]_{q},F_{n}],F_{n-1}]_{q},\dots F_{m+1}]_{q},F_{m-1}]_{q},\quad m<n,$
with
$c_{\alpha_{m}}=(-1)^{n+\delta_{m}^{\mathbf{m}}}\frac{q^{-2\delta_{m}^{\mathbf{m}}-1}\lambda}{z_{m-1}},$
and
$F_{\tilde{\alpha}_{m}}=[F_{n-1},F_{n}]_{q^{-1}},\quad m=n,$
with
$c_{\alpha_{m}}=\frac{\lambda}{z_{n-1}q^{3}}.$
$\bullet\quad\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m})$
For $m<n-1$, and $\mathbf{n}\neq 1$, we have the diagram
$\ldots$  $\alpha_{1}$$\alpha_{m}$$\alpha_{n-1}$$\alpha_{n}$$\ldots$
When $m<n-1$, and $\mathbf{n}=1$, the diagram is
$\ldots$    $\alpha_{1}$$\alpha_{m}$$\alpha_{n-1}$$\alpha_{n}$$\ldots$
The mixture parameter is
$c_{\alpha_{m}}=(-1)^{n+1+\delta_{m}^{\mathbf{m}}}\frac{q^{-2\delta_{m}^{\mathbf{m}}-1}\lambda}{z_{m-1}},\quad m<n-1$
and root vector
$F_{\tilde{\alpha}_{m}}=[[\dots[[\dots[[F_{m},F_{m+1}]_{q},F_{m+2}]_{q},\dots F_{n}]_{q},F_{n-2}]_{q},\dots F_{m+1}]_{q},F_{m-1}]_{q}.$
In the case of $m=n-1$, it corresponds to the diagram
[…]
Composite root vectors are defined as
$F_{\tilde{\alpha}_{n-1}}=[F_{n},F_{n-2}]_{q},\quad F_{\tilde{\alpha}_{n}}=[F_{n-1},F_{n-2}]_{q},\quad\mbox{and}\quad c_{\alpha_{n-1}}=c_{\alpha_{n}}=-\frac{\lambda}{q^{3}z_{n-1}}.$
$\bullet\quad\mathfrak{g}=\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m})$
The diagram is
$\ldots$ $\ldots$ $<$$\alpha_{1}$$\alpha_{m}$$\alpha_{n}$
Depending on the value of $m$ we define the root vectors
$F_{\tilde{\alpha}_{m}}=[[\dots[[\dots[[F_{m},F_{m+1}]_{q},F_{m+2}]_{q},\dots F_{n}]_{q^{2}},F_{n-1}]_{q},\dots F_{m+1}]_{q},F_{m-1}]_{q},\quad\mathrm{if}\ m<n-1,$
$F_{\tilde{\alpha}_{m}}=[[F_{n-1},F_{n}]_{q^{2}},F_{n-2}]_{q},\quad\mathrm{if}\ m=n-1\in 2\mathbb{Z},$
with the mixture parameter
$c_{\alpha_{m}}=(-1)^{n+\delta_{m}^{\mathbf{m}}}\frac{q^{-2\delta_{m}^{\mathbf{m}}-1}\lambda}{z_{m-1}},\quad m\leq\mathbf{m}.$
5.3 Diagrams of type $C$
The matrix $C$ is related with just one type of graded Satake diagram of the algebra $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|4\mathbf{m})$, which is
[…]
The composit root vectors of roots $\tilde{\alpha}$, $\alpha\in\bar{\Pi}_{\mathfrak{l}}$ are
$F_{\tilde{\alpha}_{2i}}=[[F_{2i},F_{2i+1}]_{q},F_{2i-1}]_{q}\ \mathrm{with}\ c_{\alpha_{2i}}=-\frac{x_{2i+1}}{qx_{2i-1}},\ \mathrm{where}\ i=1,\ldots,\frac{n-3}{2},$
for even $\alpha$. Also, there are two odd root vectors
$F_{\tilde{\alpha}_{n}}=[F_{n},F_{n-2}]_{q},\quad F_{\tilde{\alpha}_{n-1}}=[F_{n-1},F_{n-2}]_{q},$
with mixture parameters
$c_{\alpha_{n}}=-\frac{x_{n+2}}{x_{n}q},\quad c_{\alpha_{n-1}}=-\frac{x_{n}}{x_{n-2}q}.$
Acknowledgement.
This work is done at the Center of Pure Mathematics MIPT under the project FSMG-2023-0013. It is financially supported by Russian Science Foundation grant 23-21-00282.
The first author (D.A.) is thankful to the Deanship of Graduate Studies and Scientific Research at University of Bisha for the financial support through the Scholars Research Support Program of the University.