### v5

# Quantum super-spherical pairs

Abstract
We introduce quantum super-spherical pairs as coideal subalgebras in general linear and orthosymplectic quantum supergroups. These subalgebras play a role of isotropy subgroups for matrices solving the $\mathbb{Z}_{2}$-graded reflection equation. They generalize quantum (pseudo)-symmetric pairs of Letzter-Kolb-Regelskis-Vlaar.
2010 AMS Subject Classification: 17B37,17A70.
Key words: Quantum super-spherical pairs, Quantum symmetric pairs, Quantum supergroups, Graded reflection equation
1 Introduction
This paper is devoted to a $\mathbb{Z}_{2}$-graded generalization of quantum symmetric pairs pioneered by Letzter [1] and developed to a deep theory of quantum symmetric spaces, [2, 3, 4, 5]. The classical supergeometry is an established field of mathematics [6], [7] motivated by profound applications of supersymmetry in quantum physics [8, 9]. That refers to the concept of symmetric and, more generally, spherical spaces [10, 11]. Quantum symmetric pairs stem from the Reflection Equation (RE), which plays for them a similar fundamental role as the Yang-Baxter Equation for quantum groups. Quantum supergroups have been in research focus from the very birth of quantum groups and are now well understood [12]. It seems natural to unify these precursors in a theory of quantum super-symmetric pairs. Such an attempt was made in [13] for a special case of general linear supergroups following ideas of [14]. The current paper is a step further to a general theory of quantum super-spherical pairs. We have also to mention [15] addressing a similar topic that appeared shortly after our arXiv preprint.
We are not aiming at a comprehensive exposition of this broad area but focus on the special case of classical (basic) matrix Lie superalgebras and confine ourselves with a special symmetric polarization of the root system. Unlike [13, 15] we do not appeal to the Radford-Majid bosonization of quantum supergroups [16] but stay within the category of Hopf superalgebras. Our approach is close to [4] and based on classification of graded (generalized) Satake diagrams, for a fixed Borel subalgebra in $\mathfrak{g}$. However our logic is rather inverse comparing to [4]: we start with transparent quasiclassical conditions on a subalgebra $\mathfrak{k}$ that allows for Letzter’s quantization and arrive at an involutive automorphism of the Cartan matrix, which is a starting point of [4]. Such an automorphism is subordinate to a subset of simple roots of $\mathfrak{g}$. Together they define a generalized Satake diagram: a combinatorial datum underlying $\mathfrak{k}$. We then solve the RE for a special class of $\mathfrak{k}$ and relate the K-matrices to the corresponding coideal subalgebras as their quantum isotropy subgroups.
Recall that a subalgebra $\mathfrak{k}$ in a simple Lie algebra $\mathfrak{g}$ is called spherical if there is a Borel subalgebra $\mathfrak{b}\subset\mathfrak{g}$ such that $\mathfrak{k}+\mathfrak{b}=\mathfrak{g}$. The pair $(\mathfrak{g},\mathfrak{k})$ is the localization of a classical homogeneous spherical manifold. Such manifolds generalize symmetric spaces and feature similar nice representation theoretical properties [17]. Solutions to non-graded RE for standard $U_{q}(\mathfrak{g})$ deliver classical points on spherical manifolds, where the Poisson bracket vanishes [18]. Their isotropy Lie algebras enter spherical pairs $(\mathfrak{g},\mathfrak{k})$ with the total Lie algebra $\mathfrak{g}$.
We adopt a similar definition for super-spherical pairs assuming that both $\mathfrak{g}$ and $\mathfrak{k}$ are graded. As different Borel subalgebras in $\mathfrak{g}$ are generally not conjugated and even not isomorphic (it is an easy exercise already for $\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|2)$), this definition depends on the choice of $\mathfrak{b}$. We study spherical pairs that are quantizable along the similar lines as their non-graded (pseudo) symmetric analogs [1, 2, 4].
1.1 Current work
We define a classical pseudo-symmetric pair $(\mathfrak{g},\mathfrak{k})$ via an involutive automorphism $\tau$ of the Dynkin diagram $D_{\mathfrak{g}}$ subordinate to a certain subdiagram $D_{\mathfrak{l}}\subset D_{\mathfrak{g}}$, cf. Definition 4.9. Here $\mathfrak{l}$ is an analog of the semisimple Lie subalgebra in $\mathfrak{k}$ whose roots are fixed by the symmetry in the ordinary symmetric pair $(\mathfrak{g},\mathfrak{k})$. The key novelty here is a substitute for the longest Weyl group element $w_{\mathfrak{l}}$ of $\mathfrak{l}$, which is present in the non-graded case and which plays a crucial role in the non-graded theory. For an admissible subset $\Pi_{\mathfrak{l}}\subset\Pi$ of simple roots specified by Definition 4.5, we introduce an operator $w_{\mathfrak{l}}$ that features basic properties of the longest Weyl group element. In particular, it is an even involution that preserves the root systems of $\mathfrak{l}$ and $\mathfrak{g}$ with their weight lattices and flips the highest and lowest weights of basic $\mathfrak{l}$-submodules in $\mathfrak{g}\ominus\mathfrak{l}$ generating $\mathfrak{k}$ over $\mathfrak{l}$ and a certain subset $\mathfrak{t}$ of the Cartan subalgebra $\mathfrak{h}\subset\mathfrak{b}$. Such an operator can be defined if the total grading of $\mathfrak{g}$ induces a special ”symmetric” grading on $\mathfrak{l}$. The relation of $\tau$ with $\mathfrak{l}$ consists in the requirement $\tau|_{\Pi_{\mathfrak{l}}}=-w_{\mathfrak{l}}$.
We prove that subalgebras $\mathfrak{k}\subset\mathfrak{g}$ constructed this way are spherical (Proposition 4.16) and quantizable as coideal subalgebras in the quantum supergroup $U_{q}(\mathfrak{g})$ (Theorem 5.3). They can be conveniently parameterized by decorated Dynkin diagrams associated with the triples $(\mathfrak{g},\mathfrak{l},\tau)$, which amount to graded (generalized) Satake diagrams upon filtering through a set of selection rules in Section 4.2.
For each decorated Dynkin diagram, $\mathfrak{k}$ is generated over $\mathfrak{l}+\mathfrak{t}$ by $x_{\alpha}=e_{\alpha}+c_{\alpha}f_{\tilde{\alpha}}+\grave{c}_{\alpha}u_{\alpha}$, $\alpha\in\bar{\Pi}_{\mathfrak{l}}=\Pi\backslash\Pi_{\mathfrak{l}}$, where the positive root $\tilde{\alpha}$ is determined by $\tau$, see (4.13). Here $u_{\alpha}\in\mathfrak{h}$ is an element centralizing $\mathfrak{l}$ for even $\alpha\in\bar{\Pi}_{\mathfrak{l}}$ orthogonal to $\Pi_{\mathfrak{l}}$, and scalars $c_{\alpha}\in\mathbb{C}^{\times}$, $\grave{c}_{\alpha}\in\mathbb{C}$ are called mixture parameters. Thus every diagram gives rise to a whole family of spherical subalgebras $\mathfrak{k}\subset\mathfrak{g}$.
There are two problems arising in connection with such a combinatorial description of $\mathfrak{k}$: a) it needs to be checked if the subalgebra $\mathfrak{k}$ is proper, that is, $\mathfrak{k}\varsubsetneq\mathfrak{g}$, and b) one has to check if $\mathfrak{k}$ is related with a non-trivial K-matrix. The case $\mathfrak{k}=\mathfrak{g}$ should be regarded as trivial and not interesting. Note that, while the presence of $K$-matrix is of prime interest by itself, it is a sufficient condition for a) meaning that $\mathfrak{k}$ has a bigger supply of matrix invariants than $\mathfrak{g}$. We give a classification of graded Satake diagrams within the fixed grading by discarding decorated diagrams leading to trivial pairs with $\mathfrak{k}=\mathfrak{g}$ at any choice of mixture parameters. Our selection rules reduce to forbidding subdiagrams (4.14-4.17) to appear in a decorated Dynkin diagram. They turn out to be more intricate than the non-graded selection rules, which involve only one subdiagram (4.14), see [4] for details.
Graded generalized Satake diagrams are listed in Section 4.3. Their initial classification may be conducted by the shape of the diagram. By this we mean the non-graded decorated diagram obtained by throwing the grading away. Diagrams (4.21)–(4.26) whose shape is an admissible non-graded generalized Satake diagrams are said to be of type I. The remaining diagrams of type II are ”essentially graded” and comprise (4.18), (4.19) and (4.20). All they have odd roots in $\bar{\Pi}_{\mathfrak{l}}$.
We present K-matrices for type I in Theorem 3.1 thus proving that the corresponding pairs $(\mathfrak{g},\mathfrak{k})$ are proper. With regard to type II we conjecture that they produce proper $\mathfrak{k}$ at least for some values of mixture parameters defining $\mathfrak{k}$. For the special case of (4.20) that holds due to the K-matrix presented in Theorem 3.1. We expect that diagrams (4.18) and (4.19) lead to solutions of a twisted version (3.4) of RE relative to the matrix super-transposition. Their K-matrices are believed to be (4.32) and (4.31). Those diagrams from (4.19) associated with non-trivial flip of the Satake diagram are related with RE twisted by the outer automorphism (3.5). Their conjectured K-matrices are explicated in (4.30). With regard to the remaining part of diagrams (4.19), we guess their K-matrices (4.27), (4.28), and (4.29) by studying the simplest examples. This makes us believe that all graded Satake diagrams give rise to non-trivial spherical pairs associated with an appropriate version of RE. This discussion will be a matter of a forthcoming study.
Solutions to the RE for the spherical pairs of type I are given in Section 3, Theorem 3.1. They turn out to have a similar shape as in the non-graded case, with a certain restriction on the location of the odd root for the ortho-symplectic $\mathfrak{g}$, see Corollary 4.25.
For each Satake diagram, we quantize the universal enveloping algebra $U(\mathfrak{k})$ of the relative spherical subalgebra $\mathfrak{k}\subset\mathfrak{g}$ as a coideal subalgebra $U_{q}(\mathfrak{k})\subset U_{q}(\mathfrak{g})$, cf. Theorem 5.3. Therein we relate the K-matrices found in Section 3 to corresponding $U_{q}(\mathfrak{k})$.
2 Preliminaries
This section contains a general description of quantum supergroups deforming the universal enveloping algebras of general linear and orthosymplectic Lie superalgebras. We explicate their natural representations on the graded vector space $\mathbb{C}^{N|2\mathbf{m}}$ along with the graded R-matrices in a symmetric grading with minimal number of odd simple roots.
2.1 Quantum Supergroup $U_{q}(\mathfrak{g})$
For a textbook on quantum groups, the reader is referred to [19]. In our exposition of quantum supergroups, we follow [20].
An algebra $\mathcal{A}$ is called superalgebra if it is $\mathbb{Z}_{2}$-graded: that is, $\mathcal{A}=\mathcal{A}_{0}\oplus\mathcal{A}_{1}$, and $\mathcal{A}_{i}\mathcal{A}_{j}\subset\mathcal{A}_{i+j\!\!\mod 2}$. Elements of $\mathcal{A}_{0}$ are called even and elements of $\mathcal{A}_{1}$ are called odd.
Given two graded associative algebras $\mathcal{A}$ and $\mathcal{B}$ their tensor product $\mathcal{A}\otimes\mathcal{B}$ is a graded algebra too. The multiplication on homogeneous elements is determined by the rule
$(a_{1}\otimes b_{1})(a_{2}\otimes b_{2})=(-1)^{|a_{2}||b_{1}|}a_{1}a_{2}\otimes b_{1}b_{2}.$
where $|a|\in\{0,1\}$ stands for the degree of $a$.
In this section, we recall a general definition of quantum supergroup $U_{q}(\mathfrak{g})$ associated with a graded Lie superalgebra $\mathfrak{g}$ possessing a set of Chevalley-like generators. It is a quasi-triangular Hopf superalgebra with comultiplication ranging in the graded commutative tensor square of $U_{q}(\mathfrak{g})$. The antipode in $U_{q}(\mathfrak{g})$ is a graded anti-automorphism: $\gamma(ab)=(-1)^{|a||b|}\gamma(b)\gamma(a)$ for all homogeneous $a,b\in U_{q}(\mathfrak{g})$.
Let $\mathfrak{g}$ be a finite-dimensional complex Lie superalgebra associated with a polarized root system $\mathrm{R}=\mathrm{R}^{-}\cup\mathrm{R}^{+}$ of rank $n$ with a basis $\Pi=\{\alpha_{i}:i\in[1,n]=I\}\subset\mathrm{R}^{+}$ of the simple roots and even Cartan subalgebra $\mathfrak{h}\subset\mathfrak{g}$. Let $A=(a_{ij})_{1\leq i,j\leq n}$, denote the symmetrizable Cartan matrix and (,)(-,-)( - , - ) the corresponding non-degenerate symmetric bilinear form on the dual vector space $\mathfrak{h}^{*}$.
Fix $q\in\mathbb{C}^{\times}$ to be not a root of unity. Define quantum integers by setting
$[n]_{q}=q^{n-1}+q^{n-3}+\dots+q^{-n+3}+q^{-n+1}=\frac{q^{n}-q^{-n}}{\omega}$
for $n\in\mathbb{Z}$, with the notation, $\omega=q-\bar{q}$, $\bar{q}=q^{-1}$, used throughout the text. For each simple root $\alpha\in\Pi$ define
$q_{\alpha}=\begin{cases}q,&\mbox{if }(\alpha,\alpha)=0,\\ q^{\frac{(\alpha,\alpha)}{2}},&\mbox{if }(\alpha,\alpha)\neq 0.\end{cases}$
Denote by $\varkappa\subset I$ the subset indexing even simple roots of $\mathfrak{g}$.
Definition 2.1.
[20] The quantum supergroup $U_{q}(\mathfrak{g})$ is a complex unital associative superalgebra generated by $e_{\pm\alpha_{i}},$ and $q^{\pm h_{\alpha_{i}}}$ with grading
$|e_{\pm{\alpha_{i}}}|=\begin{cases}0,&\mbox{if }i\in\varkappa\subset I\\ 1,&\mbox{if }i\notin\varkappa.\end{cases},\quad|q^{\pm h_{\alpha_{i}}}|=0,\ \forall\ i\in I,$
such that the following relations are satisfied:
  1. (i)
    $q^{h_{\alpha_{i}}}q^{-h_{\alpha_{i}}}=q^{-h_{\alpha_{i}}}q^{h_{\alpha_{i}}}=1,$ $q^{h_{\alpha_{j}}}q^{h_{\alpha_{i}}}=q^{h_{\alpha_{i}}}q^{h_{\alpha_{j}}}$,
  2. (ii)
    $q^{h_{\alpha_{i}}}e_{\pm{\alpha_{j}}}q^{-h_{\alpha_{i}}}=q^{\pm(\alpha_{i},\alpha_{j})}e_{\pm\alpha_{j}}$,
  3. (iv)
    $[e_{\alpha_{i}},e_{-\alpha_{j}}]=e_{\alpha_{i}}e_{-\alpha_{j}}-(-1)^{|e_{-\alpha_{j}}||e_{\alpha_{i}}|}e_{-\alpha_{j}}e_{\alpha_{i}}=\delta_{ij}[h_{\alpha_{i}}]_{q_{\alpha_{i}}}$,
  4. (v)
    $(ad_{q^{\prime}}e_{\pm{\alpha_{i}}})^{v_{ij}}e_{\pm{\alpha_{j}}}=0,\quad i\neq j,\ q^{\prime}=q,\bar{q},$ where
    $(ad_{q^{\prime}}e_{\alpha_{i}})x=e_{\alpha_{i}}x-(-1)^{|e_{\alpha_{i}}||x|}(q^{\prime})^{(\alpha_{i},wt(x))}xe_{\alpha_{i}},\quad x\in U_{q}(\mathfrak{g}),$
    $v_{ij}=\begin{cases}1,&\mbox{if }(\alpha_{i},\alpha_{i})=(\alpha_{i},\alpha_{j})=0,\\ 2,&\mbox{if }(\alpha_{i},\alpha_{i})=0,\ (\alpha_{i},\alpha_{j})\neq 0,\\ 1-\frac{2(\alpha_{i},\alpha_{j})}{(\alpha_{i},\alpha_{i})},&\mbox{if }(\alpha_{i},\alpha_{i})\neq 0.\end{cases}$
The left equality in (iv) reminds that the commutator is understood in the graded sense, while the right one imposes a relation of $U_{q}(\mathfrak{g})$.
We will work with general linear Lie superalgebras too. To that end, we need to extend the previous definition of special linear $\mathfrak{g}$ by adding Cartan generators $\{q^{h_{\zeta_{i}}}\}_{i=1}^{n+1}$ with commutation relations
$q^{h_{\zeta_{i}}}e_{\pm{\alpha_{j}}}q^{-h_{\zeta_{i}}}=q^{\pm(\zeta_{i},\alpha_{j})}e_{\pm\alpha_{j}}.$
We then proceed further by setting $q^{h_{\alpha_{i}}}=q^{h_{\zeta_{i}}}q^{-h_{\zeta_{i+1}}}$.
A Hopf superalgebra structure is fixed by comultiplication defined on the generators $e_{i}=e_{\alpha_{i}}$, $f_{i}=e_{-\alpha_{i}}$, and $q^{\pm h_{i}}=q^{\pm h_{\alpha_{i}}}$ as
$\Delta(e_{i})=q^{h_{i}}\otimes e_{i}+e_{i}\otimes 1,\quad\Delta(f_{i})=f_{i}\otimes q^{-h_{i}}+1\otimes f_{i},\quad\Delta(q^{\pm h_{i}})=q^{\pm h_{i}}\otimes q^{\pm h_{i}}.$
Mind that the tensor product is supercommutative: $a\otimes b=(-1)^{|a||b|}b\otimes a$ for all homogeneous $a$ and $b$.
The counit $\epsilon$ is defined as the homomorphism $U_{q}(\mathfrak{g})\to\mathbb{C}$ that vanishes on all $e_{i}$, $f_{i}$ and returns $1$ on $q^{\pm h_{i}}$. The antipode $\gamma$ can be readily evaluated on the generators as
$\gamma(q^{\pm h_{i}})=q^{\mp h_{i}},\quad\gamma(e_{i})=-q^{-h_{i}}e_{i},\quad\gamma(f_{i})=-f_{i}q^{h_{i}}$
and extended as a graded anti-automorphism to entire $U_{q}(\mathfrak{g})$. This supercoalgebra structure extends to the Cartan generators of the quantum general linear quantum supergroup in the obvious way.
We denote by $U_{q}(\mathfrak{h}),\ U_{q}(\mathfrak{g}_{+}),\ \mathrm{and}\ U_{q}(\mathfrak{g}_{-})$ the $\mathbb{Z}_{2}$-graded $\mathbb{C}$-subalgebras of $U_{q}(\mathfrak{g})$ generated by $q^{\pm h_{i}},\ e_{i},$ and $f_{i}$ respectively.
Simple root vectors enter the set of generators of $U_{q}(\mathfrak{g})$, while composite root vectors are not given for granted. For quantum groups they can be constructed via the Lusztig braid group action on $U_{q}(\mathfrak{g})$ [19], which is not applicable to quantum supergroups. Still the composite root vectors can be obtained by a method of Khoroshkin and Tolstoy [25] based on the concept of normal ordering of positive roots, which works in both cases. As a result, root vectors $e_{\alpha}\in U_{q}(\mathfrak{g}_{+})$ and $f_{\alpha}\in U_{q}(\mathfrak{g}_{-})$ can be defined for each positive $\alpha\in\mathrm{R}^{+}$. They are used for a description of the universal R-matrix of $U_{q}(\mathfrak{g})$ in [25]. We will need them in Section 5 for construction of coideal subalgebras in $U_{q}(\mathfrak{g})$.
2.2 Basic Quantum Supergroups
Next we specify the root systems of basic (non-exceptional) matrix Lie superalgebras: general linear $\mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$ and ortho-symplectic $\mathfrak{o}\mathfrak{s}\mathfrak{p}(N|2\mathbf{m})$, $\mathfrak{s}\mathfrak{p}\mathfrak{o}(N|2\mathbf{m})$, with even $N$ in the latter case. Although $\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{m}|2\mathbf{n})$ is isomorphic to $\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m})$ as Lie superalgebras, they are considered as different triangular polarizations leading to different quantum supergroups. In particular, they have different graded Dynkin diagrams and different K-matrices, see Section 4.3.
We fix the following symmetric grading on underlying vector space $V=\mathbb{C}^{\mathbf{m}+N+\mathbf{m}}=\mathbb{C}^{N|2\mathbf{m}}$:
$|i|=\left\{\begin{array}[]{cccc}1,&i\leqslant\mathbf{m}\quad\mbox{or}\quad N+\mathbf{m}<i,\\ 0,&\mathbf{m}<i\leqslant\mathbf{m}+N.\end{array}\right.$
With respect to this grading, there are two odd simple roots in $\mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$ and $\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|2\mathbf{m})$, and one odd simple root for $\mathfrak{o}\mathfrak{s}\mathfrak{p}(N|2\mathbf{m})$, $N\neq 2$, and $\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m})$. Let $\mathfrak{g}=\mathfrak{g}_{0}\oplus\mathfrak{g}_{1}$ be a Lie superalgebra of these types. Denote the Cartan subalgebra by $\mathfrak{h}\subset\mathfrak{g}_{0}$. In the dual vector space $\mathfrak{h}^{*}$, choose a basis $\{\varepsilon_{i}\}_{i=1}^{k}\cup\{\delta_{i}\}_{i=1}^{2\mathbf{m}}$, where $k=N$ for $\mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$ and $k=\mathbf{n}$ for $\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m})$, $\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m})$, and $\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m})$. Introduce a symmetric bilinear form on $\mathfrak{h}^{*}$ by
$(\varepsilon_{i},\varepsilon_{j})=\delta_{i}^{j}=-(\delta_{i},\delta_{j}),\ (\varepsilon_{i},\delta_{j})=0.$
In the Dynkin diagrams below, an odd simple root $\alpha$ is denoted by a black node if $(\alpha,\alpha)\neq 0$, and a grey one if $(\alpha,\alpha)=0$. Even simple roots are represented by white nodes.
We denote the sets of even and odd roots with $\mathrm{R}_{0}$ and $\mathrm{R}_{1}$, respectively, so that $\mathrm{R}=\mathrm{R}_{0}\cup\mathrm{R}_{1}$. Additionally, we divide the root system into positive and negative parts $\mathrm{R}^{\pm}$ with the basis of simple roots $\Pi\subset\mathrm{R}^{+}$.
$\bullet\ \mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$
$\begin{cases}\mathrm{R}^{+}=\{\varepsilon_{i}-\varepsilon_{j}\}_{i<j}\cup\{\delta_{i}-\delta_{j}\}_{i<j}\cup\{\delta_{i}-\varepsilon_{j}\}_{i\leq\mathbf{m}}\cup\{\varepsilon_{j}-\delta_{i}\}_{i\geq\mathbf{m}+1},\\ \Pi=\{\delta_{i}-\delta_{i+1}\}_{i=1,i\neq\mathbf{m}}^{2\mathbf{m}-1}\cup\{\delta_{\mathbf{m}}-\varepsilon_{1}\}\cup\{\varepsilon_{i}-\varepsilon_{i+1}\}_{i=1}^{N-1}\cup\{\varepsilon_{N}-\delta_{\mathbf{m}+1}\}.\\ \end{cases}$
with $\mathrm{R}_{0}=\{\varepsilon_{i}-\varepsilon_{j},\delta_{i}-\delta_{j}\}_{i\neq j}$ and $\mathrm{R}_{1}=\pm\{\varepsilon_{i}-\delta_{j}\}$. The Dynkin diagram, which describes the relations among simple roots and indicates their parity, is
[…]
$\bullet\ \mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m})$. We consider separately two cases: $\mathbf{n}\neq 0$ and $\mathbf{n}=0$.
For $\mathbf{n}\neq 0$, we have
$\begin{cases}\mathrm{R}^{+}=\{\varepsilon_{i}\pm\varepsilon_{j}\}_{i<j}\cup\{\delta_{i}\pm\delta_{j}\}_{i<j}\cup\{\delta_{i}\pm\varepsilon_{j},\varepsilon_{i},2\delta_{i},\delta_{i}\},\\ \Pi=\{\delta_{i}-\delta_{i+1}\}_{i=1}^{\mathbf{m}-1}\cup\{\delta_{\mathbf{m}}-\varepsilon_{1}\}\cup\{\varepsilon_{i}-\varepsilon_{i+1}\}_{i=1}^{\mathbf{n}-1}\cup\{\varepsilon_{\mathbf{n}}\}.\\ \end{cases}$
with $\mathrm{R}_{0}=\{\pm\varepsilon_{i}\pm\varepsilon_{j},\pm 2\delta_{i},\pm\delta_{i}\pm\delta_{j},\pm\varepsilon_{i}\}_{i\neq j}$, $\mathrm{R}_{1}=\{\pm\delta_{i},\pm\delta_{i}\pm\varepsilon_{j}\}$, and Dynkin diagram
[…]
In the case of $\mathbf{n}=0$, we have
$\begin{cases}\mathrm{R}^{+}=\{\delta_{i}\pm\delta_{j}\}_{i<j}\cup\{2\delta_{i},\delta_{i}\},\\ \Pi=\{\delta_{i}-\delta_{i+1}\}_{i=1}^{\mathbf{m}-1}\cup\{\delta_{\mathbf{m}}\}.\\ \end{cases}$
with $\mathrm{R}_{0}=\{\pm 2\delta_{i},\pm\delta_{i}\pm\delta_{j}\}_{i\neq j}$, $\mathrm{R}_{1}=\{\pm\delta_{i}\}$. The Dynkin diagram is
$\ldots$$>$$\delta_{\mathbf{m}}$$\delta_{1}-\delta_{2}$$\delta_{2}-\delta_{3}$$\delta_{\mathbf{m}-1}-\delta_{\mathbf{m}}$
$\bullet\ \mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m})$. Here we also distinguish two cases: $\mathbf{n}\neq 1$ and $\mathbf{n}=1$.
For $\mathbf{n}\neq 1$, we have
$\begin{cases}\mathrm{R}^{+}=\{\varepsilon_{i}\pm\varepsilon_{j}\}_{i<j}\cup\{\delta_{i}\pm\delta_{j}\}_{i<j}\cup\{\delta_{j}\pm\varepsilon_{i},2\delta_{i}\},\\ \Pi=\{\delta_{i}-\delta_{i+1}\}_{i=1}^{\mathbf{m}-1}\cup\{\delta_{\mathbf{m}}-\varepsilon_{1}\}\cup\{\varepsilon_{i}-\varepsilon_{i+1}\}_{i=1}^{\mathbf{n}-1}\cup\{\varepsilon_{\mathbf{n}-1}+\varepsilon_{\mathbf{n}}\}.\\ \end{cases}$
with $\mathrm{R}_{0}=\{\pm\varepsilon_{i}\pm\varepsilon_{j},\pm 2\delta_{i},\pm\delta_{i}\pm\delta_{j},\}_{i\neq j}$, $\mathrm{R}_{1}=\{\pm\delta_{i}\pm\varepsilon_{j}\}$, and the Dynkin diagram
[…]
When $\mathbf{n}=1$, the root system is
$\begin{cases}\mathrm{R}^{+}=\{\delta_{i}\pm\delta_{j}\}_{i<j}\cup\{\delta_{j}\pm\varepsilon_{1},2\delta_{i}\},\\ \Pi=\{\delta_{i}-\delta_{i+1}\}_{i=1}^{\mathbf{m}-1}\cup\{\delta_{\mathbf{m}}-\varepsilon_{1}\}\cup\{\delta_{\mathbf{m}}+\varepsilon_{1}\},\end{cases}$
with $\mathrm{R}_{0}=\{\pm 2\delta_{i},\pm\delta_{i}\pm\delta_{j},\}_{i\neq j}$ and $\mathrm{R}_{1}=\{\pm\delta_{i}\pm\varepsilon_{1}\}$. Its Dynkin diagram is
[…]
$\bullet\ \mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m})$ has the root system
$\begin{cases}\mathrm{R}^{+}=\{\varepsilon_{i}\pm\varepsilon_{j}\}_{i<j}\cup\{\delta_{i}\pm\delta_{j}\}_{i<j}\cup\{\delta_{j}\pm\varepsilon_{i},2\varepsilon_{i}\},\\ \Pi=\{\delta_{i}-\delta_{i+1}\}_{i=1}^{\mathbf{m}-1}\cup\{\delta_{\mathbf{m}}-\varepsilon_{1}\}\cup\{\varepsilon_{i}-\varepsilon_{i+1}\}_{i=1}^{\mathbf{n}-1}\cup\{2\varepsilon_{\mathbf{n}}\},\\ \end{cases}$
with $\mathrm{R}_{0}=\{\pm\varepsilon_{i}\pm\varepsilon_{j},\pm 2\varepsilon_{i},\pm\delta_{i}\pm\delta_{j},\}_{i\neq j}$, $\mathrm{R}_{1}=\{\pm\delta_{i}\pm\varepsilon_{j}\}$, and the Dynkin diagram
[…]
2.3 Natural Representations
In this section we describe an irreducible representation of $U_{q}(\mathfrak{g})$ on the graded vector space $\mathrm{End}(\mathbb{C}^{N|2\mathbf{m}})$. Let $\{v_{i}\}\subset\mathbb{C}^{N|2\mathbf{m}}$ be the standard homogeneous weight basis. We call the grading symmetric (see Definition 4.2 below) if $|v_{i}|=|v_{i^{\prime}}|$ for all $i$, where
$i^{\prime}=N+2\mathbf{m}+1-i.$
We choose a symmetric grading with minimal number of odd simple roots:
$\left(\underbrace{1,\ldots,1}_{\mathbf{m}};\underbrace{0,\ldots,0}_{N};\underbrace{1,\ldots,1}_{\mathbf{m}}\right).$
We will use notation $|i|=|v_{i}|$ for all basis vectors $v_{i}$.
Let $\{e_{ij}\}_{i,j=1}^{N+2m}\subset\mathrm{End}(\mathbb{C}^{N|2\mathbf{m}})$ be the standard matrix basis with multiplication $e_{ij}e_{mn}=\delta_{jm}e_{in}$. An irreducible representation $\pi:U_{q}(\mathfrak{g})\rightarrow\mathrm{End}(\mathbb{C}^{N|2\mathbf{m}})$ is defined by the following assignment:
$q^{h_{\zeta_{i}}}\mapsto\ \sum_{j=1}^{N}q^{(-1)^{|i|}\delta_{j}^{i}}e_{jj},\quad\mathrm{if}\ i\leq n+1,\ \mathfrak{g}=\mathfrak{g}\mathfrak{l}(n+1),$
$q^{h_{i}}\mapsto$$\begin{cases}\sum_{j=1}^{N+2\mathbf{m}}q^{(-1)^{|i|}(\delta_{j}^{i}-\delta_{j}^{i^{\prime}})+(-1)^{|i+1|}(-\delta_{j}^{(i+1)}+\delta_{j}^{(i+1)^{\prime}})}e_{jj},&\mathrm{if}\ i<\mathbf{n}+\mathbf{m},\ \mathfrak{g}=\begin{cases}\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m}),\\ \mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m}),\\ \mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m}),\end{cases}\\ \sum_{j=1}^{N+2\mathbf{m}}q^{(-1)^{|i|}(\delta_{j}^{i}-\delta_{j}^{i^{\prime}})}e_{jj},&\mathrm{if}\ i=\mathbf{m}+\mathbf{n},\ \mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m}),\\ \sum_{j=1}^{N+2\mathbf{m}}q^{2\delta_{j}^{i}-2\delta_{j}^{i^{\prime}}}e_{jj},&\mathrm{if}\ i=\mathbf{m}+\mathbf{n},\ \mathfrak{g}=\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m}),\\ \sum_{j=1}^{N+2\mathbf{m}}q^{(-1)^{|i-1|}(\delta_{j}^{(i-1)}-\delta_{j}^{(i-1)^{\prime}})}q^{(-1)^{|i|}(\delta_{j}^{i}-\delta_{j}^{i^{\prime}})}e_{jj},&\mathrm{if}\ i=\mathbf{m}+\mathbf{n},\ \mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m}),\\ \end{cases}$
$e_{i}\mapsto$$\begin{cases}e_{i,i+1},&\mathrm{if}\ i\leq N+2\mathbf{m}-1,\ \mathfrak{g}=\mathfrak{g}\mathfrak{l}(N|2\mathbf{m}),\\ q^{-\delta_{i}^{\mathbf{m}}}e_{i,i+1}-(-1)^{(|i|)(|i+1|+1)}e_{(i+1)^{\prime},i^{\prime}},&\mathrm{if}\ i\leq\mathbf{m}+\mathbf{n},\ \mathfrak{g}=\begin{cases}\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m}),\\ \mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m}),\\ \mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m}),\end{cases}\\ q^{-\delta_{i-1}^{\mathbf{m}}}e_{i-1,i+1}-(-1)^{|i-1|}e_{(i+1)^{\prime},(i-1)^{\prime}},&\mathrm{if}\ i=\mathbf{m}+\mathbf{n},\ \mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m}),\\ e_{i,i+1},&\mathrm{if}\ i=\mathbf{m}+\mathbf{n},\ \mathfrak{g}=\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m}),\\ \end{cases}$
$f_{i}\mapsto\begin{cases}-e_{i+1,i},&\mbox{if }\ i=\mathbf{m},\ \mathfrak{g}=\mathfrak{g}\mathfrak{l}(N|2\mathbf{m}),\\ -qe_{i+1,i}+e_{i^{\prime},(i+1)^{\prime}},&\mbox{if }\ i=\mathbf{m},\ \mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(N|2\mathbf{m}),\ \mathrm{or}\ \mathfrak{g}=\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m}),\\ -qe_{i+1,i-1}+e_{(i-1)^{\prime},(i+1)^{\prime}},&\mbox{if }\ i=\mathbf{m}+1,\ \mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|2\mathbf{m}),\end{cases}$
and the similar assignment to the remaining $f_{i}$ as for $e_{i}$ with $e_{ij}$ changed to $e_{ji}$. In the tensor square of this representation, the universal R-matrix mentioned in Section 2.1 has the following expression, up to a scalar multiple:
$R=\sum_{i,j}q^{(-1)^{|i|}\delta_{ij}}e_{ii}\otimes e_{jj}+\omega\sum_{j<i}(-1)^{|j|}e_{ij}\otimes e_{ji},$ (2.1)
for general linear $\mathfrak{g}$ [21, 22] and
$R=\sum_{i,j=1}q^{(-1)^{|j|}(\delta_{ij}-\delta_{ij^{\prime}})}e_{ii}\otimes e_{jj}+\omega\sum\limits_{\begin{subarray}{c}j,i=1\\ j<i\end{subarray}}((-1)^{|j|}e_{ij}\otimes e_{ji}-(-1)^{|i|+|j|+|i||j|}\kappa_{i}\kappa_{j}q^{\rho_{i}-\rho_{j}}e_{ij}\otimes e_{i^{\prime}j^{\prime}}),$ (2.2)
for ortho-symplectic $\mathfrak{g}$ [21, 23]. Here
$(\rho_{i})=\begin{cases}(k-\mathbf{m},\dots,k-1;k-1,\dots,\frac{1}{2},0,-\frac{1}{2},\dots,1-k;1-k,\dots,\mathbf{m}-k),\ k=\frac{2\mathbf{n}+1}{2},\ \mathrm{for}\ \mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m}),\\ (\mathbf{n}-\mathbf{m},\dots,\mathbf{n}-1;\mathbf{n}-1,\dots,1,0,0,-1,\dots,1-\mathbf{n};1-\mathbf{n},\dots,\mathbf{m}-\mathbf{n}),\ \mathrm{for}\ \mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m}),\\ (\mathbf{n}-\mathbf{m}+1,\dots,\mathbf{n};\mathbf{n},\dots,1,-1,\dots,1-\mathbf{n};1-\mathbf{n},\dots,\mathbf{m}-\mathbf{n}-1),\ \mathrm{for}\ \mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m}),\end{cases}$
$(\kappa_{i})=\begin{cases}(-1,\dots,-1;1,\dots,1;1,\dots,1),\ \mathrm{for}\ \mathfrak{o}\mathfrak{s}\mathfrak{p}(N|2\mathbf{m}),\\ (-1,\dots,-1;1,\dots,1,-1,\dots,-1;-1,\dots,-1),\ \mathrm{for}\ \mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m}).\end{cases}$
Let us emphasise that the tensor product in (2.1) and (2.2) is graded. The identity $e_{ij}\otimes e_{lk}=(-1)^{(i+j)(l+k)}e_{lk}\otimes e_{ij}$ is satisfied for all matrix units.
3 Graded Reflection Equation
The non-graded version of RE appeared in mathematical physics literature [26, 27] and triggered the theory of quantum symmetric pairs in [28, 29, 1]. The graded RE has been of interest as well, mostly in the spectral parameter dependent form [30, 31, 32, 33].
In this section, we present a class of solutions to the constant RE for the general linear and orthosymplectic quantum supergroups. We are interested in invertible even RE-matrices. Solutions for the general linear supergroups are taken from our recent paper [24].
3.1 $\mathbb{Z}_{2}$-graded Reflection Equation
Suppose that $V$ is a graded vector space and $\mathcal{A}=\mathrm{End}(V)$ is the corresponding graded matrix algebra. An invertible element $R\in\mathcal{A}\otimes\mathcal{A}$ is called an $R$-matrix if it satisfies Yang-Baxter equation
$R_{12}R_{13}R_{23}=R_{23}R_{13}R_{12},$
where the subscripts indicate the tensor factor in the graded tensor cube of $\mathrm{End}(V)$. An even element $P=\sum_{i=1}^{n}(-1)^{|j|}e_{ij}\otimes e_{ji}\in\mathrm{End}(V)\otimes\mathrm{End}(V)$ is called graded permutation. It flips the tensor factors in $V\otimes V$ by the rule
$P(v\otimes w)=(-1)^{|v||w|}w\otimes v$
for all homogeneous $v,w\in V$. The operator $S=PR\in\mathrm{End}(V)\otimes\mathrm{End}(V)$ satisfies the braid relation
$S_{12}S_{23}S_{12}=S_{23}S_{12}S_{23}.$
A matrix $K\in\mathrm{End}(V)$ is said to satisfy RE if the identity
$SK_{2}SK_{2}=K_{2}SK_{2}S$ (3.3)
holds true in $\mathrm{End}(V)\otimes\mathrm{End}(V)$. In particular, a scalar matrix satisfies this equation. This solution is not interesting and should be considered as trivial. We will assume that $K$ is even. Note that the particular form (3.3) is related to left coideal subalgebras, cf. Section 5.
K-matrices amount to cylindric braiding in a suitable category of $U_{q}(\mathfrak{g})$-modules, [36], via a sort of fusion procedure similar to R-matrices. Cylindric brading is a collection of intertwiners $K_{M}\in\mathrm{End}_{U_{q}(\mathfrak{k})}(M)$ parameterized by a $U_{q}(\mathfrak{g})$-module $M$ that satisfy Reflection Equation
$S_{M,M}(\mathrm{id}\otimes K_{M})S_{M,M}(\mathrm{id}\otimes K_{M})=(\mathrm{id}\otimes K_{M})S_{M,M}(\mathrm{id}\otimes K_{M})S_{M,M},$
where $S_{M,M}$ is the product $P_{M,M}R_{M,M}$ of the permutation $P_{M,M}$ and the image $R_{M,M}$ of the universal matrix $\mathcal{R}$ in $\mathrm{End}(M\otimes M)$. In the non-graded case, $K_{M}$ can be obtained as the representation of an element $\mathcal{K}$ (universal K-matrix) from a completion of $U_{q}(\mathfrak{g})$ that satisfies
$\Delta(\mathcal{K})=\mathcal{R}_{21}\mathcal{K}_{1}\mathcal{R}_{12}\mathcal{K}_{2}.$
The universal K-matrix is constructed in [3] for symmetric pairs $(\mathfrak{g},\mathfrak{k})$ with symmetrizable Kac-Moody algebra $\mathfrak{g}$ of finite type, in development of [34] for the general linear case. Its construction is extended for pseudo-symmetric pairs in [4] and to all symmetrizable Kac-Moody algebras in [5]. It is natural to expect a generalization of the universal K-matrix for quantum super-spherical pairs.
Equation (3.3) admits generalizations with applications to integrable systems [35]. We present here two versions which are of relevance to left coideal subalgebras. Suppose that $\sigma\colon U_{q}(\mathfrak{g})\to U_{q}(\mathfrak{g})$ is a superinvolutive superalgebra graded anti-automorphism and supercoalgebra automorphism (Chevalley superinvolution) such that $\pi\circ\sigma(x)=\pi^{t}(x)$ for all $x\in U_{q}(\mathfrak{g})$, where $\pi$ is the representation of $U_{q}(\mathfrak{g})$ on $V$ and $t$ is the matrix super-transposition. If $R^{t_{1}t_{2}}=R_{21}$, then the identity
$R_{21}K_{1}R_{21}^{t_{1}}K_{2}=K_{2}R_{12}^{t_{2}}K_{1}R_{21}$ (3.4)
is called twisted RE. Here $t_{i}$, $i=1,2$ designate the super-transposition applied to the $i$-th tensor factor.
Now suppose that $\vartheta\colon U_{q}(\mathfrak{g})\to U_{q}(\mathfrak{g})$ is an involutive Hopf superalgebra automorphism and $\theta\colon\mathrm{End}(V)\to\mathrm{End}(V)$ a matrix super-algebra automorphism (conjugation with a fixed invertible even matrix) such that $\pi\circ\vartheta(x)=\theta\circ\pi(x)$ for all $x\in U_{q}(\mathfrak{g})$. Suppose also that $R^{\theta_{1}\theta_{2}}=(\theta\otimes\theta)(R)=R$. Then one can consider a twisted RE in the form
$R_{21}K_{1}R^{\theta_{1}}_{12}K_{2}=K_{2}R_{21}^{\theta_{2}}K_{1}R_{12},$ (3.5)
In the special case $\vartheta=\mathrm{id}$ we return to the equation (3.3). We will focus on (3.3) in the current paper.
In what follows, we mean by RE only the form (3.3) unless otherwise is explicitly stated. The case of the general linear quantum supergroup has been studied in detail in [24], where the full list of solutions to RE is given relative to an arbitrary grading of $V$. They turn out to be exactly the even matrices solving the non-graded RE. Invertible solutions occur only for a (arbitrary) symmetric grading (here we use a chance to correct a sloppy remark on page 6 of [24]) and have the form
$K=(\lambda+\mu)\sum_{i=1}^{m}e_{ii}+\lambda\sum_{i=m+1}^{N+2\mathbf{m}-m}e_{ii}+\sum_{i=1}^{m}y_{i}e_{ii^{\prime}}+\sum_{i=1}^{m}y_{i^{\prime}}e_{i^{\prime}i},$ (3.6)
where $y_{i}$ are complex numbers subject to $y_{i}y_{i^{\prime}}=-\lambda\mu\not=0$. This matrix satisfies the RE with the R-matrix (2.1). We present a class of solutions for the ortho-symplectic quantum supergroups in the next section.
3.2 K-matrices for ortho-symplectic quantum groups
In this section, we find even K-matrices for the ortho-symplectic quantum supergroups. We introduce a grading on the underlying vector space $V=\mathbb{C}^{N|2\mathbf{m}}$ by setting
$|i|=\left\{\begin{array}[]{cccc}1,&i\leqslant\mathbf{m}&\mbox{or}&N+\mathbf{m}<i,\\ 0,&\mathbf{m}<i\leqslant\mathbf{m}+N.\end{array}\right.$
This is a symmetric grading with the least number of odd simple roots. We are looking for solutions of the RE in the following three forms:
$A=\sum_{i=1}^{N+2\mathbf{m}}x_{i}e_{ii}+\sum_{i=1}^{N+2\mathbf{m}}y_{i}e_{ii^{\prime}},\quad B=\sum_{i=1}^{N+2\mathbf{m}}x_{i}e_{ii}+\sum_{i=1}^{N+2\mathbf{m}}y_{i}(e_{i,i^{\prime}-1}-e_{i+1,i^{\prime}}),$
$C=\sum\limits_{\begin{subarray}{c}i=1\\ i\neq n\end{subarray}}^{2n-1}x_{i}(e_{i,i^{\prime}-1}-e_{i+1,i^{\prime}})+x_{n}e_{n,n^{\prime}}+x_{n^{\prime}}e_{n^{\prime},n},\quad\mathrm{for}\ \mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|4\mathbf{m}).$
The matrices $A$ and $B$ will depend on an integer parameter $m$ subject to inequality $1\leq m\leq\mathbf{m}$ in the forthcoming theorem.
Theorem 3.1.
The following matrices satisfy the Reflection Equation associated with ortho-symplectic quantum groups:
$A=\sum_{i=1}^{m}\lambda(1-\kappa_{m}\kappa_{m^{\prime}}q^{-2\rho_{m}})e_{ii}+\sum_{m<i<m^{\prime}}\lambda e_{ii}+\sum\limits_{\begin{subarray}{c}i\leqslant m\end{subarray}}(y_{i}e_{i,i^{\prime}}+y_{i^{\prime}}e_{i^{\prime},i^{\prime}}),$ (3.7)
$\begin{split}B=\sum_{i=1}^{m}&\lambda(1+\kappa_{m-1}\kappa_{m^{\prime}+1}q^{-2(\rho_{m}+1)})e_{ii}+\sum_{m<i<m^{\prime}}\lambda e_{ii}\hskip 30.0pt\\ &\hskip 30.0pt+\sum\limits_{\begin{subarray}{c}i\leqslant m\\ i=1\!\!\!\!\mod 2\end{subarray}}\bigl{(}z_{i}(e_{i,i^{\prime}-1}-e_{i+1,i^{\prime}})+z_{i^{\prime}-1}(e_{i^{\prime}-1,i}-e_{i^{\prime},i+1})\bigr{)},\end{split}$ (3.8)
$C=\sum\limits_{\begin{subarray}{c}i=1\\ i=1\!\!\!\!\mod 2\\ i\neq n\end{subarray}}^{2n-1}x_{i}(e_{i,i^{\prime}-1}-e_{i+1,i^{\prime}})+x_{n}e_{n,n^{\prime}}+x_{n^{\prime}}e_{n^{\prime},n},$ (3.9)
where the parameters $y_{i},z_{i},x_{i}\in\mathbb{C}$ satisfy the conditions
  • $y_{i}y_{i^{\prime}}=\kappa_{m}\kappa_{m^{\prime}}\lambda^{2}q^{-2\rho_{m}}$,
  • $z_{i}z_{i^{\prime}-1}=-\kappa_{m-1}\kappa_{m^{\prime}+1}\lambda^{2}q^{-2(\rho_{m}+1)}$,
  • $x_{i}x_{i^{\prime}-1}=x_{n}x_{n^{\prime}}$.
Proof.
Direct calculation. ∎
The presented K-matrices exhaust all of invertible solutions to the RE for general linear $\mathfrak{g}$, and they all are necessarily even [24]. A full classification of K-matrices for orthogonal and symplectic $\mathfrak{g}$ is unknown even in the non-graded case. Theorem 3.1 gives the ”simplest” examples, see the discussion in Section 4.3. They have the following shape:
$A=$$\scriptstyle y_{1}$$\scriptstyle\lambda+\mu$$\scriptstyle\lambda+\mu$$\scriptstyle\lambda$$\scriptstyle\lambda$$\scriptstyle y_{m}$$\scriptstyle y_{m^{\prime}}$$\scriptstyle y_{1^{\prime}}$
,       $B=$
$\scriptstyle z_{1}\nu$
$\scriptstyle\lambda+\mu$$\scriptstyle\lambda+\mu$
$\scriptstyle\lambda$
$\scriptstyle\lambda$
$\scriptstyle z_{m-1}\nu$
$\scriptstyle z_{m^{\prime}}\nu$$\scriptstyle z_{2^{\prime}}\nu$   ,
$C=$$\scriptstyle x_{n}$$\scriptstyle x_{n+1}$$\scriptstyle x_{1}\nu$$\scriptstyle x_{n-2}\nu$$\scriptstyle x_{n+2}\nu$$\scriptstyle x_{2^{\prime}}\nu$
with $\nu=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}$ and $\mu=-\lambda\kappa_{m}\kappa_{m^{\prime}}q^{-2\rho_{m}}$ for $A$, $\mu=\lambda\kappa_{m-1}\kappa_{m^{\prime}+1}q^{-2(\rho_{m}+1)}$ for $B$. In both cases of $A$ and $B$, the blocks with diagonal entries of $\lambda+\mu$ and $\lambda$ are, respectively, of size $m$ and $N+2\mathbf{m}-2m$, with $m\leq\mathbf{m}$ as stipulated. We relate them with coideal subalgebras in $U_{q}(\mathfrak{g})$ in Section 5.
4 Classical super-spherical pairs
In this section we define Lie superalgebras $\mathfrak{k}\subset\mathfrak{g}$ that give rise to coideal subalgebras in generalization of the Letzter theory for non-graded quantum groups.
4.1 Weyl operator and spherical data
Let $\mathfrak{g}$ be a Lie super-algebra that features a triangular decomposition with Cartan subalgebra $\mathfrak{h}$, and let $\mathfrak{b}\subset\mathfrak{g}$ be one its Borel subalgebras containing $\mathfrak{h}$.
Definition 4.1.
A Lie super-algebra $\mathfrak{k}\subset\mathfrak{g}$ is called spherical if $\mathfrak{g}=\mathfrak{k}+\mathfrak{b}$. Then the pair $(\mathfrak{g},\mathfrak{k})$ is called spherical.
It is known that, depending on the polarization and the corresponding choice of simple root basis, Borel subalgebras in $\mathfrak{g}$ are generally not isomorphic. Thus, contrary to the non-graded case, this definition of sphericity depends upon a choice of $\mathfrak{b}$. From now on we restrict our consideration to the case when $\mathfrak{g}$ is either general linear or ortho-symplectic. The choice of $\mathfrak{b}$ is determined by a grading of the underlying natural module.
Let us fix a graded basis of weights $\zeta_{i}$, $i=1,\ldots,N$, of the natural $\mathfrak{g}$-module $\mathbb{C}^{N}$. They generate the weight lattice $\Lambda$ of $\mathfrak{g}$. As before we use the notation $i^{\prime}=N+1-i$ for all $i=1,\ldots,N$. An element $w\in SL(\Lambda)$ is said to be even if $w(\mathrm{R}_{i})=\mathrm{R}_{i}$, $i=0,1$. A grading on $\mathbb{C}^{N}$ splits $\{\zeta_{i}\}_{i=1}^{N}$ to even and odd subsets. An element $w\in SL(\Lambda)$ is clearly even if it preserves this decomposition.
Definition 4.2.
  1. 1.
    A unique $\mathbb{Z}$-linear map $w_{\mathfrak{g}}\colon\Lambda\mapsto\Lambda$ defined by the assignment $\zeta_{i}\mapsto\zeta_{i^{\prime}}$, $i=1,\ldots,N,$ is called Weyl operator.
  2. 2.
    The grading on $\mathbb{C}^{N}$ is called symmetric if the Weyl operator is even.
Clearly, a grading is symmetric if and only if every inversion $\sigma_{i}\colon\zeta_{i}\mapsto\zeta_{i^{\prime}}$, $\zeta_{j}\mapsto\zeta_{j}$, $j\not=i,i^{\prime}$, which extends to an element of $SL(\Lambda)$, is even. It is also obvious that even $w_{\mathfrak{g}}=\prod_{i=1}^{N}\sigma_{i}$ extends to an involutive orthogonal operator $w_{\mathfrak{g}}\colon\mathfrak{h}^{*}\to\mathfrak{h}^{*}$.
From now on we will work only with the minimal symmetric grading on $\mathfrak{g}$. Consider the Dynkin diagram $D$ of a basic Lie superalgebra $\mathfrak{g}$. If we discard the grading information, then we get a diagram $\tilde{D}$ that we call shape of $D$. In the special case of $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|2\mathbf{m})$, we understand by $\tilde{D}$ the Dynkin diagram of $\mathfrak{s}\mathfrak{o}(2+2\mathbf{m})$. Let $\tilde{W}$ denote the group of automorphisms of $\tilde{D}$. Denote by $\tilde{W}_{0}$ its subgroup that preserves the grading.
Lemma 4.3.
The group $\tilde{W}_{0}$ preserves the root system $\mathrm{R}$ and the weight lattice $\Lambda$.
Proof.
For general linear $\mathfrak{g}$, the group $\tilde{W}$ is generated by all permutations of $\zeta_{i}$. For ortho-symplectic $\mathfrak{g}$, it is generated by transpositions $\zeta_{i}\leftrightarrow\zeta_{j}$ and inversions $\sigma_{i}$ taking $\zeta_{i}$ to $\zeta_{i^{\prime}}=-\zeta_{i}$, $i,j\leqslant\frac{N}{2}$. Now the statement follows from the explicit description of the root systems given in Section 2.2. ∎
Remark 4.4.
Thus defined $w_{\mathfrak{g}}$ is analogous to the longest element of the Weyl group of the ordinary simple Lie algebras. In the case of $\mathfrak{g}=\mathfrak{s}\mathfrak{o}(2n)$ of odd $n$, the operator $w_{\mathfrak{g}}$ differs from the longest element by the non-trivial automorphism of the Dynkin diagram (the tail flip). This operator can be equally used for construction of non-graded pseudo-symmetric pairs, because it flips the highest and lowest weights of the minimal $\mathfrak{g}$-module, cf. Lemma 4.8. It is easier to define than the ”honest” analog of the longest Weyl group element for $\mathfrak{g}=\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m})$ with odd $\mathbf{m}$. This explains our choice of $w_{\mathfrak{g}}$.
Pick a subset $\Pi_{\mathfrak{l}}\subset\Pi$, then put $\bar{\Pi}_{\mathfrak{l}}=\Pi\backslash\Pi_{\mathfrak{l}}$, and generate a subalgebra $\mathfrak{l}=\langle e_{\alpha},f_{\alpha}\rangle_{\alpha\in\Pi_{\mathfrak{l}}}\subset\mathfrak{g}$. It is a direct sum of subalgebras, $\mathfrak{l}=\sum_{i}\mathfrak{l}_{i}$, corresponding to connected components of $\Pi_{\mathfrak{l}}$. If $\mathfrak{l}_{i}$ is of type A, then set $\hat{\mathfrak{l}}_{i}\subset\mathfrak{g}$ to be the natural $\mathfrak{g}\mathfrak{l}$-extension of $\mathfrak{l}_{i}$ and leave $\hat{\mathfrak{l}}_{i}=\mathfrak{l}_{i}$ otherwise. Denote by $\hat{\mathfrak{l}}=\oplus_{i}\hat{\mathfrak{l}}_{i}\subset\mathfrak{g}$ and by $\mathfrak{h}_{\hat{\mathfrak{l}}}^{*}$ its Cartan subalgebra. The restriction of the canonical inner product from $\mathfrak{h}^{*}$ to $\mathfrak{h}^{*}_{\hat{\mathfrak{l}}}$ is non-degenerate.
For the $i$-th connected component of $\Pi_{\mathfrak{l}}$ denote $w_{\mathfrak{l}_{i}}=w_{\hat{\mathfrak{l}}_{i}}$ and define $w_{\mathfrak{l}}=\prod_{i}w_{\mathfrak{l}_{i}}\in\mathrm{End}(\mathfrak{h}^{*}_{\mathfrak{l}}$), the Weyl operator of the subalgebra $\hat{\mathfrak{l}}$.
Definition 4.5.
We call $\Pi_{\mathfrak{l}}$ admissible if the grading on $\mathfrak{g}$ induces a minimal symmetric grading on each connected component of $\hat{\mathfrak{l}}$.
Here is an easy consequence of this definition.
Proposition 4.6.
An admissible set $\Pi_{\mathfrak{l}}$ cannot contain an isolated grey odd root.
Proof.
Indeed, if $\alpha\in\Pi_{\mathfrak{l}}$ is such a root, then the corresponding subalgebra $\hat{\mathfrak{l}}_{k}$ is isomorphic to $\mathfrak{g}\mathfrak{l}(1|1)$. It has no symmetric grading, and the operator $w_{\mathfrak{l}_{k}}$ is not even. ∎
Note that an isolated black odd root is possible: one can take $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(1|2\mathbf{m})$ and $\Pi_{\mathfrak{l}}=\{\alpha_{\mathbf{m}}\}$.
If $\Pi_{\mathfrak{l}}$ is admissible, then the operator $w_{\mathfrak{l}}$ is even.
Lemma 4.7.
For admissible $\Pi_{\mathfrak{l}}$, the operator $w_{\mathfrak{l}}\in\mathrm{End}(\mathfrak{h}^{*}_{\hat{\mathfrak{l}}})$ extends to an involutive operator on $\mathfrak{h}^{*}$ that is even on the weight lattices of $\mathfrak{l}$ and $\mathfrak{g}$.
Proof.
The set of weights $\Lambda_{\mathfrak{g}}(V)=\{\zeta_{i}\}$ of the basic module splits into a disjoint union $\Lambda_{\hat{\mathfrak{l}}}(V)\cup\overline{\Lambda_{\hat{\mathfrak{l}}}(V)}$. The operator $w_{\mathfrak{l}}$ preserves the weight lattice of $\hat{\mathfrak{l}}$ and its root system by construction. It acts as a permutation on $\Lambda_{\mathfrak{g}}(V)$ identical on $\overline{\Lambda_{\hat{\mathfrak{l}}}(V)}$, therefore it preserves the weight lattice of $\mathfrak{g}$. Direct examination shows $w_{\mathfrak{l}}(\mathrm{R})=\mathrm{R}$ for the root system of $\mathfrak{g}$. Furthermore, it preserves the parity of all weights from $\Lambda(V)$, therefore it is even. Since the inner product on $\mathfrak{h}^{*}_{\hat{\mathfrak{l}}}$ is non-degenerate, we have orthogonal decomposition $\mathfrak{h}^{*}=(\mathfrak{h}_{\hat{\mathfrak{l}}}^{*})^{\perp}\oplus\mathfrak{h}^{*}_{\hat{\mathfrak{l}}}$, where $\mathfrak{h}^{*}_{\hat{\mathfrak{l}}}$ is spanned by $\Lambda_{\hat{\mathfrak{l}}}(V)$ whereas $(\mathfrak{h}_{\hat{\mathfrak{l}}}^{*})^{\perp}$ by $\overline{\Lambda_{\hat{\mathfrak{l}}}(V)}$. Finally, $w_{\mathfrak{l}}$ extends as the identity operator on $(\mathfrak{h}_{\hat{\mathfrak{l}}}^{*})^{\perp}$, therefore it is orthogonal and involutive. ∎
The vector subspaces
$\mathfrak{m}_{+}=\sum_{\alpha\in\mathrm{R}^{+}_{\mathfrak{g}}\>-\>\mathrm{R}^{+}_{\mathfrak{l}}}\mathfrak{g}_{\alpha},\quad\mathfrak{m}_{-}=\sum_{\alpha\in\mathrm{R}^{+}_{\mathfrak{g}}\>-\>\mathrm{R}^{+}_{\mathfrak{l}}}\mathfrak{g}_{-\alpha},$
are graded $\mathfrak{l}$-modules. For $\alpha\in\bar{\Pi}_{\mathfrak{l}}$ let $V^{\pm}_{\alpha}\subset\mathfrak{g}_{\pm}$ denote the $\mathfrak{l}$-submodule generated by $e_{\pm\alpha}\in\mathfrak{g}_{\pm\alpha}$.
Lemma 4.8.
Suppose that $\mathfrak{g}$ is a basic matrix Lie superalgebra and $\Pi_{\mathfrak{l}}\subset\Pi$ is admissible. Then the operator $w_{\mathfrak{l}}$ flips the highest and lowest weights of $V^{\pm}_{\alpha}$ for each $\alpha\in\bar{\Pi}_{\mathfrak{l}}$.
Proof.
The subalgebra $\mathfrak{l}$ is a sum of $\mathfrak{l}=\sum_{i}\mathfrak{l}_{i}$ of basic Lie superalgebras $\mathfrak{l}_{i}$ corresponding to connected components of the Dynkin diagram of $\mathfrak{l}$. In all cases excepting $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m})$, the modules $V^{\pm}_{\alpha}$ are tensor products of smallest fundamental $\mathfrak{l}_{i}$-modules, for which $w_{\mathfrak{l}}$ does the job.
[…]
From now on we will consider only admissible $\Pi_{\mathfrak{l}}$. Suppose that $\tau\in\mathrm{Aut}(\Pi)$ is an even permutation that coincides with $-w_{\mathfrak{l}}$ on $\Pi_{\mathfrak{l}}$. Set $\tilde{\alpha}=w_{\mathfrak{l}}\circ\tau(\alpha)\in\mathrm{R}^{+}$ for all $\alpha\in\bar{\Pi}_{\mathfrak{l}}$.
Definition 4.9.
The triple $(\mathfrak{g},\mathfrak{l},\tau)$ is called pseudo-symmetric if
$(\mu+\tilde{\mu},\alpha)$ $=0,$ $\quad\forall\mu\in\bar{\Pi}_{\mathfrak{l}},\quad\alpha\in\Pi_{\mathfrak{l}},$ (4.10)
$(\mu+\tilde{\mu},\nu-\tilde{\nu})$ $=0,$ $\quad\forall\mu,\nu\in\bar{\Pi}_{\mathfrak{l}}.$ (4.11)
Condition (4.10) has the following algebraic interpretation.
Lemma 4.10.
As a $\mathbb{Z}$-linear endomorphism of the root lattice, $\tau$ commutes with $w_{\mathfrak{l}}$.
Proof.
By construction, $\tau$ and $w_{\mathfrak{l}}$ commute when restricted to $\mathfrak{h}^{*}_{\mathfrak{l}}$. It suffices to check that for simple roots from $\bar{\Pi}_{\mathfrak{l}}$.
By Lemma 4.8, $w_{\mathfrak{l}}$ takes the highest weight of an irreducible $\mathfrak{l}$-module $V^{\pm}_{\mu}$, $\mu\in\bar{\Pi}_{\mathfrak{l}}$, to the lowest weight and vice versa. For each $\mu\in\bar{\Pi}_{\mathfrak{l}}$ we have $w_{\mathfrak{l}}(\mu)=\mu+\eta$ for some weight $\eta\in\mathbb{Z}_{+}\Pi_{\mathfrak{l}}$ that satisfies $w_{\mathfrak{l}}(\eta)=-\eta$ because $w_{\mathfrak{l}}^{2}=\mathrm{id}$. The weight $\eta$ depends only on the projection of $\tau(\mu)$ to $\mathfrak{h}^{*}_{\mathfrak{l}}$, therefore $-\tau(\mu)=w_{\mathfrak{l}}(-\tilde{\mu})=-w_{\mathfrak{l}}\bigl{(}\tau(\mu)\bigr{)}+\eta$ or $\tau(\mu)+\eta=w_{\mathfrak{l}}\bigl{(}\tau(\mu)\bigr{)}$ due to the condition (4.10). Then, since $\tau(\eta)=-w_{\mathfrak{l}}(\eta)$ for all $\eta\in\mathfrak{h}^{*}_{\mathfrak{l}}$,
$\tau\bigl{(}w_{\mathfrak{l}}(\mu)\bigr{)}=\tau(\mu+\eta)=\tau(\mu)+\tau(\eta)=\tau(\mu)-w_{\mathfrak{l}}(\eta)=\tau(\mu)+\eta=w_{\mathfrak{l}}\bigl{(}\tau(\mu)\bigr{)},$
as required. ∎
Define a linear map $\theta=-w_{\mathfrak{l}}\circ\tau\colon\mathfrak{h}^{*}\to\mathfrak{h}^{*}$. It preserves $\mathrm{R}$ because so do $\tau$ and $w_{\mathfrak{l}}$. Furthermore, $\theta(\alpha)=-\tilde{\alpha}$ for $\alpha\in\bar{\Pi}_{\mathfrak{l}}$ and $\theta(\alpha)=\alpha$ for $\alpha\in\Pi_{\mathfrak{l}}$. As a product of two even operators on $\Lambda$, it is even. The system of equalities (4.10) and (4.11) is equivalent to
$\bigl{(}\alpha+\theta(\alpha),\beta-\theta(\beta)\bigr{)}=0,\quad\forall\alpha,\beta\in\Pi.$ (4.12)
Proposition 4.11.
Condition (4.12) is fulfilled if and only if the permutation $\tau$ is even and extends to an involutive orthogonal operator on $\mathfrak{h}^{*}$ coinciding with $-w_{\mathfrak{l}}$ on $\Pi_{\mathfrak{l}}$.
Proof.
First of all remark that a linear operator being orthogonal and involutive is the same as symmetric and involutive, or orthogonal and symmetric simultaneously. Also, an automorphism of the Dynkin diagram is exactly an orthogonal operator on $\mathfrak{h}^{*}$ that permutes the basis $\Pi$.
Condition (4.12) is bilinear and therefore holds true for any pair of vectors from $\mathfrak{h}^{*}$. Setting $\alpha=\beta$ in (4.12) we find that $\theta$ is an isometry. Then (4.12) translates to
$\bigl{(}\theta(\alpha),\beta\bigr{)}=\bigl{(}\alpha,\theta(\beta)\bigr{)},\quad\forall\alpha,\beta\in\Pi.$
It means that $\theta$ is a symmetric operator. Therefore it is an involutive orthogonal operator. So is $w_{\mathfrak{l}}$, that commutes with $\theta$ by Lemma 4.10. Hence $\tau$ is an involutive orthogonal operator on $\mathfrak{h}^{*}$.
Conversely, suppose that $\tau$ is orthogonal, involutive, and coincides with $-w_{\mathfrak{l}}$ on $\Pi_{\mathfrak{l}}$. Then
$(\alpha,\mu)=\bigl{(}\tau(\alpha),\tau(\mu)\bigr{)}=-\bigl{(}w_{\mathfrak{l}}(\alpha),\tau(\mu)\bigr{)}=-\bigl{(}\alpha,w_{\mathfrak{l}}\circ\tau(\mu)\bigr{)}=-(\alpha,\tilde{\mu})$
for all $\alpha\in\Pi_{\mathfrak{l}}$ and $\mu\in\bar{\Pi}_{\mathfrak{l}}$. Thus, the condition (4.10) is fulfilled, and $\tau$ commutes with $w_{\mathfrak{l}}$ by Lemma 4.10. Therefore $\theta=-w_{\mathfrak{l}}\circ\tau$ is orthogonal and involutive, which implies (4.12). ∎
We arrive at a necessary condition for a triple $(\mathfrak{g},\mathfrak{l},\tau)$ to be pseudo-symmetric.
Corollary 4.12.
Suppose that $\tau$ satisfies conditions of Proposition 4.11. Then $\mathfrak{l}$-modules $\mathfrak{m}_{+}$ and $\mathfrak{m}_{-}$ are isomorphic.
Proof.
[…]
The isomorphism between $V_{\alpha}^{\pm}$ and $V_{\alpha^{\prime}}^{\mp}$, $\alpha^{\prime}=\tau(\alpha)$, is a consequence of (4.10). We extend $\mathfrak{l}$ by adding $\mathfrak{t}=\mathrm{Span}\{h_{\tilde{\alpha}}-h_{\alpha}\}$, for all $\alpha\in\bar{\Pi}_{\mathfrak{l}}$, then $V_{\alpha}^{\pm}$ and $V_{\alpha^{\prime}}^{\mp}$ will be isomorphic as $\mathfrak{l}+\mathfrak{t}$-modules, provided (4.11) is fulfilled. The same will be true for $\mathfrak{m}_{+}$ and $\mathfrak{m}_{-}$.
Suppose that $\Pi_{\mathfrak{l}}$ is admissible and $\tau$ satisfies the conditions of Proposition 4.11. Let $\mathfrak{c}\subset\mathfrak{h}$ denote the centralizer of $\mathfrak{l}$ in $\mathfrak{h}$. For each $\alpha\in\bar{\Pi}_{\mathfrak{l}}$ pick $c_{\alpha}\in\mathbb{C}^{\times}$, $\grave{c}_{\alpha}\in\mathbb{C}$, and $u_{\alpha}\in\mathfrak{c}$ assuming $u_{\alpha}\not=0$ only if $\alpha$ is even, orthogonal to $\Pi_{\mathfrak{l}}$, and $\tilde{\alpha}=\alpha=\tau(\alpha)$. Put
$\begin{array}[]{rcl}y_{\alpha}&=&h_{\alpha}-h_{\tilde{\alpha}},\\ x_{\alpha}&=&e_{\alpha}+c_{\alpha}f_{\tilde{\alpha}}+\grave{c}_{\alpha}u_{\alpha},\end{array}$ (4.13)
for all $\alpha\in\bar{\Pi}_{\mathfrak{l}}$. Define a Lie subalgebra $\mathfrak{k}\subset\mathfrak{g}$ as the one generated by $\mathfrak{l}+\mathfrak{t}$ and by $x_{\alpha}$ with $\alpha\in\bar{\Pi}_{\mathfrak{l}}$.
Definition 4.13.
The pair of Lie superalgebras $\mathfrak{k}\subset\mathfrak{g}$ corresponding to a pseudo-symmetric triple $(\mathfrak{g},\mathfrak{l},\tau)$ is called pseudo-symmetric.
The complex numbers $c_{\alpha},\grave{c}_{\alpha}$ in (4.13) are called mixture parameters.
Next we prove that pseudo-symmetric pairs are spherical. Denote by $\mathfrak{p}^{0}=\mathfrak{h}+\mathfrak{l}$ the Levi subalgebra with commutant $\mathfrak{l}$ and set
$\mathfrak{p}_{\pm}^{1}=\sum_{\mu\in\Pi_{\mathfrak{g}/\mathfrak{l}}}V_{\mu}^{\pm}+\mathfrak{p}^{0},\quad\mathfrak{p}_{\pm}^{m+1}=[\mathfrak{p}_{\pm}^{m},\mathfrak{p}^{1}_{\pm}],\quad m>0.$
We have is an $\mathfrak{l}$-equivariant filtration of parabolic subalgebras $\mathfrak{p}_{\pm}$:
$\mathfrak{p}^{0}\subset\mathfrak{p}_{\pm}^{1}\subset\ldots\subset\mathfrak{p}_{\pm}.$
Lemma 4.14.
For all $m,n\geqslant 1$, there is an inclusion $[\mathfrak{p}^{m}_{+},\mathfrak{p}^{n}_{-}]\subset\mathfrak{p}^{m}_{+}+\mathfrak{p}^{n}_{-}$.
Proof.
For $m=n=1$ this readily follows by root arguments. Suppose this is proved for $m=1$ and $n\geqslant 1$. The (graded) Jacobi identity gives
$[\mathfrak{p}^{1}_{+},\mathfrak{p}^{n+1}_{-}]=[\mathfrak{p}^{1}_{+},[\mathfrak{p}^{1}_{-},\mathfrak{p}^{n}_{-}]]\subset[\mathfrak{p}^{1}_{+}+\mathfrak{p}^{1}_{-},\mathfrak{p}^{n}_{-}]+[\mathfrak{p}^{1}_{-},\mathfrak{p}^{1}_{+}+\mathfrak{p}^{n}_{-}]\subset\mathfrak{p}^{1}_{+}+\mathfrak{p}^{n+1}_{-},$
which proves the formula for $m=1$ and all $n$. Now suppose that it is true for $m\geqslant 1$. Then
$[\mathfrak{p}^{m+1}_{+},\mathfrak{p}^{n}_{-}]=[[\mathfrak{p}^{m}_{+},\mathfrak{p}^{1}_{+}],\mathfrak{p}^{n}_{-}]\subset[\mathfrak{p}^{m}_{+}+\mathfrak{p}^{n}_{-},\mathfrak{p}^{1}_{+}]+[\mathfrak{p}^{m}_{+},\mathfrak{p}^{1}_{+}+\mathfrak{p}^{n}_{-}]\subset\mathfrak{p}^{m+1}_{+}+\mathfrak{p}^{n}_{-},$
which completes the proof. ∎
Corollary 4.15.
Let $m_{+}$ be the number of pluses and $m_{-}$ the number of minuses in a sequence $(\varepsilon_{1},\ldots,\varepsilon_{k})$, where $\varepsilon_{i}\in\{\pm\}$. Then
$[\ldots[\mathfrak{p}^{1}_{\varepsilon_{1}},\mathfrak{p}^{1}_{\varepsilon_{2}}],\ldots,\mathfrak{p}^{1}_{\varepsilon_{k}}]\subset\mathfrak{p}^{m_{+}}_{+}+\mathfrak{p}^{m_{-}}_{-}.$
Proof.
For $k=2$, this follows from the definition of $\mathfrak{p}_{\pm}^{m}$ and from Lemma 4.14. Suppose the statement is proved for $k\geqslant 2$. Then, by induction,
$[\ldots[\mathfrak{p}^{1}_{\varepsilon_{1}},\mathfrak{p}^{1}_{\varepsilon_{2}}],\ldots,\mathfrak{p}^{1}_{\varepsilon_{k+1}}]\subset[\mathfrak{p}^{m_{+}}_{+}+\mathfrak{p}^{m_{-}}_{-},\mathfrak{p}^{1}_{\varepsilon_{k+1}}],$
and the statement follows from Lemma 4.14 for general $k$. ∎
Let $V_{\beta}\subset V^{+}_{\beta}\oplus V^{-}_{\beta^{\prime}}\oplus\mathbb{C}u_{\beta}$ with $\beta\in\bar{\Pi}_{\mathfrak{l}}$ be the $\mathfrak{l}$-module generated by $x_{\beta}$ (the element $u_{\beta}\in\mathfrak{c}$ can be non-zero only if $V^{+}_{\beta}$ and $V^{-}_{\beta^{\prime}}$ are trivial $\mathfrak{l}$-modules). We define $\mathfrak{k}^{0}=\mathfrak{t}+\mathfrak{l}$ and
$\mathfrak{k}^{1}=\sum_{\beta\in\Pi_{\mathfrak{g}/\mathfrak{l}}}V_{\beta}+\mathfrak{k}^{0},\quad\mathfrak{k}^{m+1}=[\mathfrak{k}^{m},\mathfrak{k}^{1}],\quad m>0.$
The $\mathfrak{l}$-modules $\mathfrak{k}^{m}$ form an ascending filtration of $\mathfrak{k}$. Furthermore, condition (4.11) implies
$[y_{\alpha},x_{\beta}]=\bigl{(}(\alpha,\beta)-(\tilde{\alpha},\beta)\bigr{)}e_{\beta}+c_{\beta}\bigl{(}-(\alpha,\tilde{\beta})+(\tilde{\alpha},\tilde{\beta})\bigr{)}f_{\tilde{\beta}}\propto x_{\beta}$
for all $\alpha,\beta\in\bar{\Pi}_{\mathfrak{l}}$. The rightmost implication holds because the right-hand side of the equality simply vanishes if $\beta$ is orthogonal to $\Pi_{\mathfrak{l}}$ and $\beta=\tilde{\beta}$. We conclude that the filtration $\mathfrak{k}^{0}\subset\mathfrak{k}^{1}\subset\ldots\subset\mathfrak{k}$ is $\mathfrak{t}$- and therefore $\mathfrak{k}^{0}$-invariant.
Proposition 4.16.
Pseudo-symmetric Lie superalgebras are spherical.
Proof.
Observe that $\mathfrak{k}^{0}\subset\mathfrak{p}^{0}$ and $\mathfrak{k}^{1}+\mathfrak{p}^{1}_{-}=\mathfrak{p}_{-}^{1}+\mathfrak{p}^{1}_{+}$ by construction. Suppose that we have proved an equality $\mathfrak{k}^{m}+\mathfrak{p}^{m}_{-}=\mathfrak{p}_{-}^{m}+\mathfrak{p}^{m}_{+}$ for $m\geqslant 1$. For any sequence $(x_{\mu_{i}})_{i=1}^{m+1}$ we have
$[\ldots[x_{\mu_{1}},x_{\mu_{2}}],\ldots,x_{\mu_{m+1}}]=[\ldots[e_{\mu_{1}},e_{\mu_{2}}],\ldots,e_{\mu_{m+1}}]+c[\ldots[f_{\tau(\mu_{1})},f_{\tau(\mu_{2})}],\ldots,f_{\tau(\mu_{m+1})}]+\ldots,$
where $c\in\mathbb{C}$ and the omitted terms are in $\mathfrak{p}^{m}_{+}+\mathfrak{p}^{m}_{-}$, by Corollary 4.15. By the induction assumption, the above expression simplifies to
$=[\ldots[e_{\mu_{1}},e_{\mu_{2}}],\ldots,e_{\mu_{m+1}}]\mod\mathfrak{k}^{m}+\mathfrak{p}^{m+1}_{-}.$
Since the filtration is $\mathfrak{l}$-invariant, and all $[\ldots[e_{\mu_{1}},e_{\mu_{2}}],\ldots,e_{\mu_{m+1}}]$ generate $\mathfrak{p}^{m+1}_{+}\mod\mathfrak{p}^{m}_{+}$ we conclude that $\mathfrak{k}^{m+1}+\mathfrak{p}^{m+1}_{-}\supset\mathfrak{p}^{m+1}_{+}$ and $\mathfrak{k}^{m+1}+\mathfrak{p}^{m+1}_{-}\supset\mathfrak{p}^{m+1}_{+}+\mathfrak{p}^{m+1}_{-}$. This implies
$\mathfrak{k}+\mathfrak{b}_{-}=(\mathfrak{k}+\mathfrak{l})+\mathfrak{b}_{-}=\mathfrak{k}+(\mathfrak{l}+\mathfrak{b}_{-})=\mathfrak{k}+\mathfrak{p}_{-}\supset\mathfrak{p}_{+}+\mathfrak{p}_{-}=\mathfrak{g}.$
The proof is complete. ∎
In the next section we address the question when the subalgebra $\mathfrak{k}\subset\mathfrak{g}$ is proper for a given pseudo-symmetric triple, in the special symmetric grading that we have fixed.
4.2 Decorated diagrams and selection rules
Like in the non-graded case the permutation $\tau$ entering a pseudo-symmetric triple $(\mathfrak{g},\mathfrak{l},\tau)$ is an involutive even automorphism of the Dynkin diagram coinciding with $-w_{\mathfrak{l}}$ on $\Pi_{\mathfrak{l}}$. Such triples can be visualized via decorated Dynkin diagrams similar to non-graded pseudo-symmetric triples considered in [4]. In the graded setting, we use black colour for nodes in $\Pi_{\mathfrak{l}}$ and white from $\bar{\Pi}_{\mathfrak{l}}$ regardless of their parity.
In this section, we work out criteria for a decorated diagram when the subalgebra $\mathfrak{k}$ equals $\mathfrak{g}$ for all values of the mixture parameters. Recall that in the non-graded case there is a selection rule that is formulated as follows. Let $C(\beta)$ denote the union of connected components of $\Pi_{\mathfrak{l}}$ that are connected to $\{\beta,\tau(\beta)\}$. Then diagrams with
[…]
should be ruled out since $\mathfrak{k}=\mathfrak{g}$ in this case [4].
Graded selection rules are more intricate. They amount to a set of lemmas which guarantee that under certain conditions on $\beta\in\bar{\Pi}_{\mathfrak{l}}$ the Lie superalgebra $\mathfrak{g}^{(\beta)}=\mathrm{Span}\{e_{\beta},f_{\beta},h_{\beta}\}$ along with $\mathfrak{l}$-modules it generates lies in $\mathfrak{k}$. That will be further used to prove the equality $\mathfrak{k}=\mathfrak{g}$ in Proposition 4.24. Now let us demonstrate that the only non-graded selection rule fails to be definite in the super case.
Lemma 4.17.
Suppose that a decorated Dynkin diagram is such that (4.14) holds for some $\beta\in\bar{\Pi}_{l}$. Then $\mathfrak{g}^{(\beta)}\subset\mathfrak{k}$ unless $\beta$ is odd and $\alpha$ is even.
Proof.
Suppose that (4.14) is fulfilled with some $\beta\in\bar{\Pi}_{\mathfrak{l}}$ and $\alpha\in\Pi_{\mathfrak{l}}$ First of all observe that $\alpha$ cannot be odd, because otherwise $w_{\mathfrak{l}}$ will not be even, by Proposition 4.6 (that will also violate the condition (4.10)).
Furthermore, for even $\alpha$, we have $\tilde{\beta}=\beta+\alpha$. With $x_{\beta}=e_{\beta}+c_{\beta}[f_{\alpha},f_{\beta}]\in\mathfrak{k}$, we define
$x_{\tilde{\beta}}=[e_{\alpha},x_{\beta}]=[e_{\alpha},e_{\beta}]+c_{\beta}[h_{\alpha},f_{\beta}]=[e_{\alpha},e_{\beta}]+c_{\beta}f_{\beta}\in\mathfrak{k}.$
Since $(\alpha,\beta)\not=0$, there is no term $\grave{c}_{\beta}u_{\beta}$ in $x_{\beta}$. Then
$[x_{\beta},x_{\tilde{\beta}}]=[e_{\beta}+c_{\beta}[f_{\alpha},f_{\beta}],[e_{\alpha},e_{\beta}]+c_{\beta}f_{\beta}]=c_{\beta}\bigl{(}h_{\beta}+(-1)^{|\beta|}h_{\beta}+(-1)^{|\beta|}h_{\alpha}\bigr{)}\in\mathfrak{k}.$
Since $h_{\alpha}\in\mathfrak{k}$, we conclude that $h_{\beta}\in\mathfrak{k}$ and therefore $e_{\beta},f_{\beta}\in\mathfrak{k}$ if and only if $\beta$ is even. ∎
It follows from the proof of this lemma that the diagram $\alpha$$\beta$ with even $\alpha$ and odd $\beta$ gives rise to a 5-dimensional spherical subalgebra $\mathfrak{k}=\mathrm{Span}\{e_{\alpha},h_{\alpha},f_{\alpha},x_{\beta},x_{\tilde{\beta}}\}$ in $\mathfrak{g}=\mathfrak{g}\mathfrak{l}(2|1)$. This $\mathfrak{k}$ is beyond the scope of the current study because this grading of $\mathfrak{g}\mathfrak{l}(2|1)$ is not symmetric. As a subdiagram, it nevertheless appears, for instance, in
with two grey odd tail nodes. These diagrams yield non-trivial pairs $(\mathfrak{g},\mathfrak{k})$ with $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|4)$. They can be extended further to the left with alternating white and black nodes, cf. (4.20) and (4.26).
Similarly to shapes of graded Dynkin diagrams we consider shapes of decorated diagrams by discarding grading. In other words, a graded decorated diagram is obtained from its shape by adding the grading datum.
Proposition 4.18.
Let $\mathcal{D}$ be a decoration of a Dynkin diagram $D$. Then the shape of $\mathcal{D}$ is a decorated shape of $D$.
Proof.
We can define the Weyl operator in the non-graded case as we did in in the graded, cf. Remark 4.4. Then for each $\Pi_{\mathfrak{l}}$, the operator $w_{\mathfrak{l}}$ is the same as in the non-graded shape. Now notice that the automorphism $\tau$ of a graded diagram is an automorphism of its shape that coincides with $-w_{\mathfrak{l}}$ on $\Pi_{\mathfrak{l}}$. ∎
Proposition 4.6 states that a grey odd root cannot constitute a connected component of $\Pi_{\mathfrak{l}}$. It turns out that such a node cannot be isolated from $\Pi_{\mathfrak{l}}$ in all cases.
Lemma 4.19.
Suppose that a decorated Dynkin diagram contains a subdiagram
with $\alpha,\beta\in\bar{\Pi}_{\mathfrak{l}}$, where $\alpha$ is of any parity, a grey odd node $\beta=\tau(\beta)$ is isolated from $\Pi_{\mathfrak{l}}$ and $(\beta,\alpha)\not=0$. Then $\mathfrak{g}^{(\alpha)}+\mathfrak{g}^{(\beta)}\subset\mathfrak{k}$.
Proof.
Indeed, we conclude that $h_{\beta}\in\mathfrak{k}$ because
$[x_{\beta},x_{\beta}]=x_{\beta}x_{\beta}+x_{\beta}x_{\beta}=[e_{\beta},c_{\beta}f_{\beta}]+[c_{\beta}f_{\beta},e_{\beta}]=2c_{\beta}h_{\beta}.$
Observe that $(\beta,\tilde{\alpha})=\bigl{(}\beta,\tau(\alpha)+\ldots\bigr{)}=\bigl{(}\tau(\beta),\tau(\alpha)\bigr{)}=(\beta,\alpha)$, where we have suppressed terms from $\mathbb{Z}\Pi_{\mathfrak{l}}$. Commuting $h_{\beta}$ with $x_{\alpha}=e_{\alpha}+c_{\alpha}f_{\tilde{\alpha}}+\grave{c}_{\alpha}u_{\alpha}$ we find
$[h_{\beta},x_{\alpha}]=(\beta,\alpha)(e_{\alpha}-c_{\alpha}f_{\tilde{\alpha}})\in\mathfrak{k}.$
Thus we conclude that $e_{\alpha}\in\mathfrak{k}$, $f_{\tilde{\alpha}}\in\mathfrak{k}$, and therefore $f_{\alpha},h_{\alpha}\in\mathfrak{k}$, if $\grave{c}_{\alpha}u_{\alpha}=0$. Finally, $h_{\alpha}$ splits $x_{\beta}$, so $e_{\beta},f_{\beta}\in\mathfrak{k}$ either.
Now suppose that $\grave{c}_{\alpha}u_{\alpha}\not=0$. Then $\tilde{\alpha}=\alpha$, and, since $h_{\beta}\in\mathfrak{k}$ and $(\alpha,\beta)\neq 0$, we have a system
$\left\{\begin{array}[]{ccc}e_{\alpha}+c_{\alpha}f_{\alpha}+\grave{c}_{\alpha}u_{\alpha}&=&0\mod\mathfrak{k},\\ e_{\alpha}-c_{\alpha}f_{\alpha}&=&0\mod\mathfrak{k},\end{array}\right.\quad\Leftrightarrow\quad\left\{\begin{array}[]{cccl}e_{\alpha}+c_{+}u_{\alpha}&=0&\mod\mathfrak{k},&c_{+}=\grave{c}_{\alpha}/2,\\ f_{\alpha}+c_{-}u_{\alpha}&=0&\mod\mathfrak{k},&c_{-}=\grave{c}_{\alpha}/2c_{\alpha}.\end{array}\right.$
Commuting these elements with $h_{\beta}\in\mathfrak{k}$, we obtain $e_{\alpha},f_{\alpha}\in\mathfrak{k}$ and hence $h_{\alpha}\in\mathfrak{k}$. ∎
Lemma 4.20.
Suppose that decorated Dynkin diagram contains a subdiagram
where $\alpha$ and $\gamma$ are even, $\beta$ is odd, and $\gamma$ with $\sigma$ form an arbitrary connected even diagram of rank 2. Then $\mathfrak{g}^{(\beta)}+\mathfrak{g}^{(\sigma)}\subset\mathfrak{k}$.
Proof.
By the assumption, $\tau(\beta)=\beta$ and $\tau(\sigma)=\sigma$. Choose the grading such that $(\alpha,\beta)=1=-(\beta,\gamma)$. We can set $\grave{c}_{\beta}=0$ and $\grave{c}_{\sigma}=0$ because $\beta$ and $\sigma$ are not isolated from $\Pi_{\mathfrak{l}}$. Evaluating commutators of $e_{\alpha}$ and $e_{\gamma}$ with $x_{\beta}=e_{\beta}+c_{\beta}[[f_{\alpha},f_{\beta}],f_{\gamma}]\in\mathfrak{k}$, we obtain $x_{\tilde{\beta}}=c_{\beta}f_{\beta}-[[e_{\alpha},e_{\beta}],e_{\gamma}]\in\mathfrak{k}$. Then $[x_{\beta},x_{\tilde{\beta}}]=c_{\beta}(2h_{\beta}+h_{\alpha}+h_{\gamma})=c_{\beta}h_{\delta}\in\mathfrak{k}.$ Since
$(\delta,\sigma)=(\gamma,\sigma)\not=-(\delta,\tilde{\sigma})=-\Bigl{(}2\beta+\gamma,\sigma-\frac{2(\gamma,\sigma)}{(\gamma,\gamma)}\gamma+\ldots\Bigl{)}=-4\frac{(\gamma,\sigma)}{(\gamma,\gamma)}+(\gamma,\sigma),$
the element $h_{\delta}$ splits $x_{\sigma}=e_{\sigma}+c_{\sigma}f_{\tilde{\sigma}}$ and therefore $x_{\tilde{\sigma}}\propto f_{\sigma}+c_{-\sigma}e_{\tilde{\sigma}}\in\mathfrak{k}$ with some $c_{-\sigma}\in\mathbb{C}^{\times}$. Hence $e_{\sigma},f_{\sigma}$, and $h_{\sigma}\in\mathfrak{k}$. The element $h_{\sigma}$ splits $x_{\beta}$ and $x_{\tilde{\beta}}$ because $(\sigma,\beta)=0$ and $-(\sigma,\tilde{\beta})=-(\sigma,\gamma)\neq 0$. Therefore $e_{\beta},f_{\beta},h_{\beta}\in\mathfrak{k}$ either. ∎
The next lemma addresses diagrams with even orthogonal shape.
Lemma 4.21.
Suppose that a decorated Dynkin diagram of type $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m})$ with rank $n$ contains one of the subdiagrams
where $\alpha$, $\gamma$, and $\sigma$ are even and $\beta$ is odd. Then $\mathfrak{g}^{(\beta)}+\mathfrak{g}^{(\sigma)}+\mathfrak{g}^{(\sigma^{\prime})}\subset\mathfrak{k}$.
Proof.
The reasoning is similar to the proof of Lemma 4.20. We will only indicate the differences. First we assume that $\grave{c}_{\sigma}=0$ for the diagram on the right. By the assumption, $\tau(\beta)=\beta$ and $\tau(\sigma)=\sigma^{\prime}$ for left diagram and $\tau(\sigma)=\sigma$ for the right one. As before, we argue that $h_{\delta}$ with $\delta=2\beta+\alpha+\gamma$ is in $\mathfrak{k}$. Further we find that
$(\delta,\sigma)=(\gamma,\sigma)=-1\not=-(\delta,\tilde{\sigma})=-(2\beta+\gamma,\alpha_{n}+\gamma)=1\ \mathrm{(for\ left\ diagram)},$
$(\delta,\sigma)=(2\beta,\sigma)=-2\not=-(\delta,\tilde{\sigma})=-(2\beta,\alpha_{n})=2\ \mathrm{(for\ right\ diagram)}.$
As a consequence, the element $h_{\delta}$ splits $x_{\sigma}$, $x_{\tilde{\sigma}}$, and $x_{\beta}$. For the left diagram, the rest of the proof is essentially the same as in Lemma 4.20. For the right diagram, it can be proved that $x_{\beta}$ and $x_{\tilde{\beta}}$ are separated by the element $h_{\sigma}\in\mathfrak{k}$ because
$(\sigma,\beta)=1\neq-(\sigma,\tilde{\beta})=-(\sigma,\alpha+\beta+\gamma)=-(\sigma,\beta)=-1.$
In both cases, $e_{\beta},f_{\beta},h_{\beta}\in\mathfrak{k}$.
Now suppose that the term $\grave{c}_{\sigma}u_{\sigma}$ is present in $x_{\sigma}$ for the diagram on the right. We proceed as in the proof of Lemma 4.19 and demonstrate that $\mathfrak{g}^{(\sigma)}\in\mathfrak{k}$. Then we complete the proof as before. ∎
Definition 4.22.
We call a triple $(\mathfrak{g},\mathfrak{l},\tau_{\mathfrak{l}})$ and the corresponding decorated Dynkin diagram trivial if the subalgebra $\mathfrak{k}$ they generate coincides with $\mathfrak{g}$ for all values of mixture parameters $c_{\alpha}\in\mathbb{C}^{\times}$, $\grave{c}_{\alpha}\in\mathbb{C}$, $\alpha\in\bar{\Pi}_{\mathfrak{l}}$.
We will say that a decorated diagram $D$ violates selection rules if either $D$ contains an even subdiagram (4.14) or one of the subdiagrams (4.15), (4.16), (4.17). It is important to note that we allow for arbitrary orientation of these subdiagrams while the orientation of the total diagram is fixed as in Section 2.2.
For a diagram $D$, we define a subdiagram $D_{0}\subset D$ of $\mathfrak{g}\mathfrak{l}$-type by throwing away the tail. Specifically, we remove from $D$ the two last (rightmost) nodes if the shape $\tilde{D}$ of $D$ is of even orthogonal type, and the last node if $\tilde{D}$ is either odd orthogonal or symplectic. For general linear $D$, we set $D_{0}=D$.
Lemma 4.23.
If $D$ violates the selection rule, then $\tau|_{D_{0}}=\mathrm{id}$ in all cases.
Proof.
We need to consider only the case of general linear $\mathfrak{g}$. Then only subdiagrams (4.14), (4.15), or (4.16) may occur in $D$. For the subdiagrams (4.14) and (4.16) the statement is obvious. In the case of (4.15), the symmetric grading under consideration requires the presence of two odd nodes. Therefore $\beta$ cannot be $\tau$-fixed if $\tau$ reverts $D$. Thus $\tau$ is identical in all cases. ∎
Proposition 4.24.
If a decorated Dynkin diagram violates selection rules, then it is trivial.
Proof.
Denote by $L\subset D$ a subdiagram that violates the selection rules. Split the set of nodes of total diagram to a disjoint union $D=D^{l}\cup D^{r}$, where $D^{l}$ comprises the nodes on the left of $L$. In the case when the shape $\tilde{D}$ of $D$ is even orthogonal, and $L=\{\alpha,\beta\}$ is in the tail of $D$, we include all the three tail nodes in $D^{r}$. The further proof will be done in two steps. First, moving from $L$ rightward, we demonstrate that $\mathfrak{g}^{(\mu)}\subset\mathfrak{k}$ for all $\mu\in D^{r}$. Then we proceed leftward and prove that for the nodes from $D^{l}$.
It follows that $\tau$ is identical on $\{\alpha,\beta\}\subset L$. If $\tilde{D}$ is even orthogonal and $L=\{\alpha,\beta\}$ is in the tail, then $g^{(\mu)}\subset\mathfrak{k}$ for all $\mu\in D^{r}$ by Lemmas 4.17 and 4.19. We arrive to the same conclusion if $L$ is (4.17), by Lemma 4.21: then $L=D^{r}$.
Consider the case when $L\subset D_{0}$ assuming that $L$ is either (4.14) or (4.15). Put $L_{1}=L$ and denote by $\mathfrak{l}_{1}$ the subalgebra with roots in $L_{1}$. We have an inclusion $\mathfrak{l}_{1}\subset\mathfrak{k}$. Suppose that we have constructed $L_{k}\subset D_{0}^{r}$ and $\mathfrak{l}_{k}\subset\mathfrak{k}$ for $k\geqslant 1$. If there is a node $\mu\in D_{0}^{r}\cap\bar{\Pi}_{\mathfrak{l}}$ on the right of $L_{k}$ that is connected to $L_{k}$, then $\mathfrak{g}^{(\mu)}\subset\mathfrak{k}$, because the $\mathfrak{l}_{k}$-module generated by the root vector $e_{\mu}$ is not self-dual (mind that $\tau$ is identical on $D_{0}$ by Lemma 4.23). Let $C_{\mu}\subset D_{0}^{r}\cap\Pi_{\mathfrak{l}}$ be the connected component that is connected to $\mu$ from the right, that is, $C_{\mu}\cap L_{k}=\varnothing$. Then we set $L_{k+1}=L_{k}\cup\{\mu\}\cup C_{\mu}$ and define $\mathfrak{l}_{k+1}=\mathfrak{l}_{k}+\mathfrak{g}^{(\mu)}+\sum_{\nu\in C_{\mu}}\mathfrak{g}^{(\nu)}$.
It is clear that $L_{k}=D_{0}^{r}$ and for sufficiently large $k$. Finally, by considering a tail node $\mu$ of the initial diagram $D^{r}$ we conclude that $\mathfrak{g}^{(\mu)}\subset\mathfrak{k}$ for all $\mu\in D^{r}$.
Now we start moving leftward to process nodes from $D^{l}$. This time we include in consideration the diagram $L$ as in (4.17). Suppose there is a node $\mu\in D^{l}\cap\bar{\Pi}_{\mathfrak{l}}$. We can assume that it is the rightmost among them. Then the mixed generator $x_{\mu}=f_{\mu}+c_{\mu}e_{\tilde{\mu}}+\grave{c}_{\mu}u_{\alpha}$ is split by the subalgebra $\mathfrak{m}$ comprising the roots on the right of $\mu$, because $f_{\mu}$, $e_{\tilde{\mu}}$, and $u_{\mu}$ transform differently under $\mathfrak{m}$. ∎
As a consequence, we derive the following restriction on the position of the odd root in some ortho-symplectic diagrams.
Corollary 4.25.
Suppose that $\mathbf{m}<m$. Then the subalgebra $\mathfrak{k}$ corresponding to a diagram (4.23)-(4.25) exhausts all of ortho-symplectic Lie superalgebra $\mathfrak{g}$.
Proof.
These diagrams correspond to K-matrices of the form $A$ and $B$ from Theorem 3.1. For a K-matrix of type $A$ such a diagram is ruled out by Lemma 4.19, while of type $B$ by Lemma 4.20. ∎
4.3 Graded Satake diagrams
To avoid conflict with the conventional notation of graded Dynkin diagrams, we do not use coloured nodes to indicate their parity. Black nodes will designate roots from $\Pi_{\mathfrak{l}}$, white the nodes from $\bar{\Pi}_{\mathfrak{l}}$, and the odd nodes will be denoted with square.
Our selection rules allow for the following decorated graded Dynkin diagrams with odd nodes in $\bar{\Pi}_{\mathfrak{l}}$:
of $\mathfrak{g}\mathfrak{l}$-type and
of $\mathfrak{o}\mathfrak{s}\mathfrak{p}$- and $\mathfrak{s}\mathfrak{p}\mathfrak{o}$-types, where the big circles on the right stand for even tail. The subdiagrams enclosed in the brackets mean the period including zero occurrence for ()(\>\>)( ) and at least once for [][\>\>][ ].
Observe that shapes of (4.18) and (4.19) are not admissible generalized Satake diagrams from [4]. Here is yet another diagram that shares this property:
It may be therefore attributed to the family with white tail depicted on the right in (4.19).
All other admissible diagrams have non-graded generalized Satake shape; they are listed below. We arrange them in families of diagrams of the same shape, indicating possible position of odd roots with square nodes. Recall that the number of such nodes and their type have been fixed for each $\mathfrak{g}$ by the minimal symmetric grading of the underlying vector space. In particular, there are two symmetrically allocated odd simple roots for $\mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$ and $\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|2\mathbf{m})$, and exactly one such root otherwise. An arc connects the root $\tau(\alpha)$ with $\alpha\in\bar{\Pi}_{\mathfrak{l}}$ if they are distinct, in the usual way. Black nodes depict simple roots from $\Pi_{\mathfrak{l}}$, whereas white nodes label elements of $\bar{\Pi}_{\mathfrak{l}}$.
Definition 4.26.
Diagrams (4.21)-(4.26) are said to be of type I. Diagrams (4.18) and (4.19) along with (4.20) will be referred to as of type II. They are all called $\mathbb{Z}_{2}$-graded Satake diagrams.
We also extend the above classification to the corresponding spherical pairs. Graded Satake diagrams are the only decorated Dynkin diagrams that comply with the selection rules of the previous section.
Conjecture 1.
Graded Satake diagrams are non-trivial.
Diagrams of type I yield non-trivial $\mathfrak{k}$ for certain values of mixture parameters because they correspond to K-matrices of type $A$ and $B$ from Theorem 3.1. That implies inequality $\dim\mathrm{End}_{\mathfrak{k}}(V)>\dim\mathrm{End}_{\mathfrak{g}}(V)$ for the basic module $V$, and therefore such $\mathfrak{k}$ are proper in $\mathfrak{g}$.
Further we present arguments in support that diagrams of type II are non-trivial either. In the special case of (4.20) the coideal subalgebra centralizes a K-matrix of type $C$, cf. Section 5.3. We guess the following K-matrices for black tailed diagrams in (4.19):
$K$ $=$ $\sum_{i=1}^{m}(\lambda+\mu)e_{ii}+\sum_{m<i<m^{\prime}}\lambda e_{ii}+\sum_{i\leqslant\mathbf{m}}(y_{i}e_{i,i^{\prime}}+y_{i^{\prime}}e_{i^{\prime},i})$ (4.27)
$+$ $\sum_{\mathbf{m}<i<m\atop i-\mathbf{m}=1\!\!\!\!\mod 2}z_{i}(e_{i,i^{\prime}-1}-e_{i+1,i^{\prime}})+z_{i^{\prime}-1}(e_{i^{\prime}-1,i}-e_{i^{\prime},i+1}),\quad\mbox{where}$
$y_{i}y_{i^{\prime}}=-\lambda\mu,\quad i\leqslant\mathbf{m},\quad z_{i}z_{i^{\prime}-1}=-\lambda\mu,\quad\mathbf{m}<i<m,\quad\mu=\kappa_{m}\kappa_{m^{\prime}}q^{-2\rho_{m+1}}\lambda.$
The parameter $m$ is the position of the rightmost white node in the Satake diagram while $\mathbf{m}$ is that of the odd root. One can see that the shape of matrix $B$ from Theorem 3.1 may be viewed as a degeneration of (4.27).
The K-matrix for the white tailed diagrams in (4.19) is conjectured to be
$K$ $=$ $\sum_{i=1}^{n}(\mu+\lambda)e_{ii}+\lambda e_{n+1,n+1}+\sum_{\mathbf{m}<i\leqslant n}(y_{i}e_{i,i^{\prime}}+y_{i^{\prime}}e_{i^{\prime},i})$ (4.28)
$+$ $\sum_{i<\mathbf{m}\atop i=1\!\!\!\!\mod 2}z_{i}(e_{i,i^{\prime}-1}-e_{i+1,i^{\prime}})+z_{i^{\prime}-1}(e_{i^{\prime}-1,i}-e_{i^{\prime},i+1}).$
The term $\lambda e_{n+1,n+1}$ is present only for $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m})$ (recall that $n$ stands for the rank $\mathbf{m}+\mathbf{n}$ of $\mathfrak{g}$). The other parameters are subject to the conditions
$\left\{\begin{array}[]{ccl}y_{i}y_{i^{\prime}}&=&-\lambda\mu,\quad\mbox{for}\quad\mathbf{m}<i\leqslant n,\\ z_{i}z_{i^{\prime}-1}&=&-\lambda\mu,\quad\mbox{for}\quad i<\mathbf{m},\end{array}\right.\quad\mu=\left\{\begin{array}[]{ccl}-\lambda,&\mathfrak{g}=&\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m}),\\ -q^{-1}\lambda,&\mathfrak{g}=&\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m}),\end{array}\right.$
The parameters $\lambda$ and $\mu$ are eigenvalues of $K$. The expression (4.28) is a generalization of the matrix $C$, which is yet another justification for relating the diagram (4.20) to type II.
The K-matrix for the middle diagram in (4.19) is believed to be
$K$ $=$ $\sum_{i=1}^{n}(\mu+\lambda)e_{ii}+\sum_{i=1}^{n-2}(y_{i}e_{i,i^{\prime}}+y_{i^{\prime}}e_{i^{\prime},i})$ (4.29)
$+$ $z_{n}(e_{n,n^{\prime}-1}-e_{n+1,n^{\prime}})+z_{n^{\prime}-1}(e_{n^{\prime}-1,n}-e_{n^{\prime},n+1}).$
with arbitrary eigenvalues $\lambda$, $\mu$, and the relations $y_{i}y_{i^{\prime}}=-\lambda\mu$, $z_{n}z_{n^{\prime}-1}=-\lambda\mu$. Conjugation with $\sum_{i\not=n,n^{\prime}}e_{ii}+e_{n,n^{\prime}}+e_{n^{\prime},n}$ produces another K-matrix that corresponds to the tail flipped Satake diagram.
The above matrices cover all diagrams in (4.19) apart from white tailed with $\tau\not=\mathrm{id}$ for $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{m}|2\mathbf{n})$. The representation $\pi$ intertwines the automorphism of $U_{q}(\mathfrak{g})$ that flips the tail root vectors with conjugation by the matrix $\sum_{i\not=n,n^{\prime}}e_{ii}+e_{n,n^{\prime}}+e_{n^{\prime},n}$. The K-matrix for such diagrams is conjectured to satisfy (3.5) and equal
$K$ $=$ $\sum_{i=1}^{n-1}(\mu+\lambda)e_{ii}+\sum_{\mathbf{m}<i\leqslant n}(y_{i}e_{i,i^{\prime}}+y_{i^{\prime}}e_{i^{\prime},i})$ (4.30)
$+$ $\sum_{i<\mathbf{m}\atop i=1\!\!\!\!\mod 2}z_{i}(e_{i,i^{\prime}-1}-e_{i+1,i^{\prime}})+z_{i^{\prime}-1}(e_{i^{\prime}-1,i}-e_{i^{\prime},i+1}),$
where $\mu=-q^{-2}\lambda$, $y_{n}=y_{n^{\prime}}=\lambda$, $y_{i^{\prime}}y_{i}=-\lambda\mu$, for $i=\mathbf{m}+1,\ldots n-1$, and $z_{i}z_{i^{\prime}-1}=-\lambda\mu,\quad\mbox{for}\quad i<\mathbf{m}$.
It is interesting to note that eigenvalues of the K-matrix (4.28) depend on two independent parameters for $\mathfrak{g}=\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m})$ and only one parameter for $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m})$ although $\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m})\simeq\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{m}|2\mathbf{n})$. This may manifest that their quantum supergroups, which differ by the choice of Borel subalgebra in $\mathfrak{g}$, are not isomorphic as Hopf superalgebras.
The diagrams in (4.18) are related to the twisted RE in the form of (3.4). The basic representation intertwines the matrix supertransposition $A\mapsto A^{t}$, $A^{t}_{ij}=A_{ji}(-1)^{(|i|+|j|)|i|}$ with a superalgebra anti-automorphphism $\sigma\colon U_{q}(\mathfrak{g})\to U_{q}(\mathfrak{g})$ acting by
$e_{\alpha}\mapsto a_{\alpha}f_{\alpha}q^{h_{\alpha}},\quad f_{\alpha}\mapsto(-1)^{|\alpha|}a_{\alpha}^{-1}q^{-h_{\alpha}}e_{\alpha},\quad q^{\pm h_{\alpha}}\mapsto q^{\pm h_{\alpha}},$
with an appropriate choice of $a_{\alpha}\in\mathbb{C}^{\times}$. The K-matrix for the right diagrams in (4.18) corresponds to $\mathfrak{g}=\mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$ and is believed to be
$K$ $=$ $\sum_{i\leqslant\mathbf{m}\atop i=1\!\!\!\!\mod 2}z_{i}(e_{i,i+1}-qe_{i+1,i})+z_{i^{\prime}-1}(e_{i^{\prime}-1,i^{\prime}}-qe_{i^{\prime},i^{\prime}-1})+\sum_{\mathbf{m}<i<\mathbf{m}^{\prime}}y_{i}e_{ii},$ (4.31)
assuming $\mathbf{m}\in 2\mathbb{N}$.
The left diagrams in (4.18) for $\mathfrak{g}=\mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$ imply that $N\in 2\mathbb{N}$. The K-matrix for it is conjectured to be
$K$ $=$ $\sum_{i\leqslant\mathbf{m}}y_{i}e_{ii}+y_{i^{\prime}}e_{i^{\prime},i^{\prime}}+\sum_{\mathbf{m}<i<\mathbf{m}^{\prime}\atop i-\mathbf{m}=1\!\!\!\!\mod 2}z_{i}(e_{i,i+1}-q^{-1}e_{i+1,i}).$ (4.32)
The parameters $y_{i}$ and $z_{i}$ in (4.31) and (4.32) are arbitrary.
The hypothetical K-matrices presented in this section cover all graded Satake diagrams of type II. They have been checked in low dimensions, and we expect that they satisfy the appropriate RE in general.
4.4 Spherical pairs and Reflection Equation
Consider matrices $A,B,C$ from Theorem 3.1, and the invertible matrix $A$ from (3.6). Denote their classical limit ($q\rightarrow 1$) by $A_{0}$, $B_{0}$, and $C_{0}$ ($=C$ as it is independent of $q$). The subalgebra $\mathfrak{k}$ is centralizing this classical limit. As in the non-graded case, we describe it with the map $\theta=-w_{\mathfrak{l}}\circ\tau$, which takes $\alpha\in\bar{\Pi}_{\mathfrak{l}}$ to $-\tilde{\alpha}\in\mathrm{R}^{-}_{\mathfrak{g}}$.
The action of $\theta$ on the basis of $\mathfrak{h}^{*}$ is as follows (recall that $n$ stands for the rank of $\mathfrak{g}$). In the case of matrix $A_{0}$ it is
$\theta(\zeta_{i})=\begin{cases}\zeta_{i^{\prime}},&\mathrm{if}\ i\leqslant m,\\ \zeta_{i},&\mathrm{if}\ m<i,\\ \end{cases},\quad\zeta_{i}=\delta_{i}\ \mathrm{or}\ \varepsilon_{i},$
for $\mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$ and
$\theta(\delta_{i})=\begin{cases}-\delta_{i},&\mathrm{if}\ i\leqslant m,\\ \delta_{i},&\mathrm{if}\ i>m,\\ \end{cases},\quad\theta(\varepsilon_{i})=\varepsilon_{i}.$
for ortho-symplectic $\mathfrak{g}$. In the case of matrix $B_{0}$ for ortho-symplectic $\mathfrak{g}$ it is
$\theta(\delta_{i})=\begin{cases}-\delta_{i+1},&\mathrm{if}\ i<m\ \mathrm{is\ odd},\\ -\delta_{i-1},&\mathrm{if}\ i\leqslant m\ \mathrm{is\ even},\\ \delta_{i},&\mathrm{if}\ i>m,\\ \end{cases}\quad\theta(\varepsilon_{i})=\varepsilon_{i}.$
In the case of matrix $C_{0}$ for ortho-symplectic $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|4\mathbf{m})$ it is
$\theta(\delta_{i})=\begin{cases}-\delta_{i+1},&\mathrm{if}\ i<n\ \mathrm{is\ odd},\\ -\delta_{i-1},&\mathrm{if}\ i\leqslant n-1\ \mathrm{is\ even},\\ \end{cases}\quad\theta(\varepsilon_{1})=-\varepsilon_{1}.$
The root basis of the subalgebra $\mathfrak{l}$ is explicitly
$\Pi_{\mathfrak{l}}=\begin{cases}\{\alpha_{i}\}_{i=m+1}^{n-m},&\mbox{for }\mathfrak{g}\mathfrak{l}(N|2\mathbf{m}),\\ \{\alpha_{i}\}_{i=m+1}^{n},&\mbox{for }\mathfrak{o}\mathfrak{s}\mathfrak{p}(N|2\mathbf{m}),\ \mbox{and}\ \mathfrak{s}\mathfrak{p}\mathfrak{o}(N|2\mathbf{m})\ \mbox{related\ to}\ A_{0},\\ \{\alpha_{2i+1}\}_{i=0}^{\frac{m}{2}-1}\cup\{\alpha_{i}\}_{i=m+1}^{n},&\mbox{for }\mathfrak{o}\mathfrak{s}\mathfrak{p}(N|2\mathbf{m}),\ \mbox{and}\ \mathfrak{s}\mathfrak{p}\mathfrak{o}(N|2\mathbf{m})\ \mbox{related\ to}\ B_{0}.\\ \{\alpha_{2i+1}\}_{i=0}^{\frac{n-3}{2}},&\mbox{for }\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|4\mathbf{m}),\ \mbox{related\ to}\ C_{0}.\\ \end{cases}$
For simple roots $\alpha\in\bar{\Pi}_{\mathfrak{l}}$ in the case of $A_{0}$ the roots $\tilde{\alpha}$ are given by
$\tilde{\alpha}_{i}=\alpha_{i^{\prime}-1},\quad i=1,\ldots,m-1,\quad i=m=\frac{N+2\mathbf{m}}{2},\quad\tilde{\alpha}_{i}=\alpha_{i^{\prime}-1},\quad i=m^{\prime}+1,\ldots,n,$
$\tilde{\alpha}_{m}=\sum_{l=m+1}^{n-m+1}\alpha_{l},\quad\tilde{\alpha}_{n-m+1}=\sum_{l=m}^{n-m}\alpha_{l},$
for $\mathfrak{g}=\mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$,
$\tilde{\alpha}_{i}=\alpha_{i},\quad i=1,\ldots,m-1,\quad\tilde{\alpha}_{m}=\alpha_{m}+2\sum_{l=m+1}^{n}\alpha_{l},$
for $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m})$,
$\tilde{\alpha}_{i}=\alpha_{i},\quad i=1,\ldots,m-1,\quad\tilde{\alpha}_{m}=\alpha_{m}+2\sum_{l=m+1}^{n-2}\alpha_{l}+\alpha_{n-1}+\alpha_{n},\quad m<n-1,$
$\tilde{\alpha}_{n-1}=\alpha_{n},\quad\tilde{\alpha}_{n}=\alpha_{n-1},\quad m=n-1,$
for $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m})$, and
$\tilde{\alpha}_{i}=\alpha_{i},\quad i=1,\ldots,m-1,\quad\tilde{\alpha}_{m}=\alpha_{m}+2\sum_{l=m+1}^{n-1}\alpha_{l}+\alpha_{n},\quad m<n-1,$
$\tilde{\alpha}_{m}=\alpha_{n-1}+\alpha_{n},\quad m=n-1,$
for $\mathfrak{g}=\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m})$.
For simple roots $\alpha\in\bar{\Pi}_{\mathfrak{l}}$, in the case of $B_{0}$, the roots $\tilde{\alpha}$ are given by
$\tilde{\alpha}_{2i}=\sum_{l=2i-1}^{2i+1}\alpha_{l},\quad i=1,\ldots,\frac{m}{2}-1,\quad\tilde{\alpha}_{m}=\alpha_{m-1}+\alpha_{m}+2\sum_{l=m+1}^{n}\alpha_{l},$
for $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m})$,
$\tilde{\alpha}_{2i}=\sum_{l=2i-1}^{2i+1}\alpha_{l},\quad i=1,\ldots,\frac{m}{2}-1,\quad\tilde{\alpha}_{m}=\alpha_{m-1}+\alpha_{m}+2\sum_{l=m+1}^{n-2}\alpha_{l}+\alpha_{n-1}+\alpha_{n},$
$\tilde{\alpha}_{n-1}=\alpha_{n-2}+\alpha_{n},\quad\tilde{\alpha}_{n}=\alpha_{n-2}+\alpha_{n-1},\quad m=n-1,$
for $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m})$,
$\tilde{\alpha}_{2i}=\sum_{l=2i-1}^{2i+1}\alpha_{l},\quad i=1,\ldots,\frac{m}{2}-1,\quad\tilde{\alpha}_{m}=\alpha_{m-1}+\alpha_{m}+2\sum_{l=m+1}^{n-1}\alpha_{l}+\alpha_{n},$
for $\mathfrak{g}=\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m})$.
For simple roots $\alpha\in\bar{\Pi}_{\mathfrak{l}}$, in the case of $C_{0}$, the roots $\tilde{\alpha}$ are given by
$\tilde{\alpha}_{2i}=\sum_{l=2i-1}^{2i+1}\alpha_{l},\quad i=1,\ldots,\frac{n-3}{2},$
$\tilde{\alpha}_{n-1}=\alpha_{n-2}+\alpha_{n-1},\quad\tilde{\alpha}_{n}=\alpha_{n-2}+\alpha_{n},$
for $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|4\mathbf{m})$.
The Lie superalgebra $\mathfrak{k}$ is generated by $\mathfrak{l}$ and additional elements $x_{\alpha}=e_{\alpha}+c_{\alpha}f_{\tilde{\alpha}}+\grave{c}_{\alpha}u_{\alpha}$, $h_{\tilde{\alpha}}-h_{\alpha}$, where $\alpha\in\bar{\Pi}_{\mathfrak{l}}$, and $c_{\alpha}\in\mathbb{C}^{\times}$, $\grave{c}_{\alpha}\in\mathbb{C}$. The scalar $\grave{c}_{\alpha}\neq 0$ only if $\mathfrak{g}=\mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$ and $\alpha=\alpha_{\frac{N+2\mathbf{m}}{2}}$; then $u_{\alpha}=h_{\alpha}$. The mixture parameters $c_{\alpha}$ are determined by the K-matrices.
5 Coideal subalgebras and K-matrices
Recall that a total order on the set of positive roots $\mathrm{R}^{+}\supset\Pi$ of $\mathfrak{g}$ is called normal if every sum $\alpha+\beta\in\mathrm{R}^{+}$ with $\alpha,\beta\in\mathrm{R}^{+}$ is between $\alpha$ and $\beta$.
Choose a normal order $\beta^{1},\beta^{2},\ldots$ on $\mathrm{R}^{+}$ and extend $F_{\beta}=q^{h_{\beta}}f_{\beta}$ with simple $\beta$ to all $\beta\in\mathrm{R}^{+}$ as it is done for $e_{-\beta}$ in [25]. Denote by $\mathcal{U}_{m}^{-}$ the subalgebra in $\mathcal{U}^{-}$ generated by $F_{\beta^{i}}$ with $i\leqslant m$.
Lemma 5.1.
For each $m=1,\ldots,|\mathrm{R}^{+}|$, $\mathcal{U}_{m}^{-}$ is a left coideal subalgebra in $U_{q}(\mathfrak{g})$. Furthermore,
$\Delta(F_{\beta^{m}})=(F_{\beta^{m}}\otimes 1+q^{h_{\beta^{m}}}\otimes F_{\beta^{m}})+U_{q}(\mathfrak{g})\otimes\mathcal{U}^{-}_{m-1}.$
Proof.
This readily follows from [25], Prop. 8.3. upon the assignment $F_{\alpha}\mapsto e_{-\alpha}$ on simple root vectors, extended as an algebra automorphism to $U_{q}(\mathfrak{b}_{-})$ (and a coalgebra anti-isomorphism). ∎
We apply this fact to quantize $U(\mathfrak{k})$. Pick $\alpha\in\bar{\Pi}_{\mathfrak{l}}$, put $\alpha^{\prime}=\tau(\alpha)\in\bar{\Pi}_{\mathfrak{l}}$ and $\tilde{\alpha}=w_{\mathfrak{l}}(\alpha^{\prime})\in\mathrm{R}^{+}$. Consider the root system generated by $\alpha^{\prime}$ and $\Pi_{\mathfrak{l}}$; denote by $\Pi_{\mathfrak{k}_{\tilde{\alpha}}}$ its connected component containing $\alpha^{\prime}$, and by $\mathfrak{k}_{\tilde{\alpha}}$ the subalgebra generated by the corresponding simple root vectors. Denote also $\Pi_{\mathfrak{l}_{\tilde{\alpha}}}=\Pi_{\mathfrak{k}_{\tilde{\alpha}}}\cap\Pi_{\mathfrak{l}}$ and by $\mathfrak{l}_{\tilde{\alpha}}$ the corresponding subalgebra in $\mathfrak{l}$.
Proposition 5.2.
For each $\alpha\in\Pi_{\alpha}$ there is a normal order on $\mathrm{R}^{+}_{\mathfrak{k}_{\tilde{\alpha}}}$ such that
  • $\alpha^{\prime}$ is in the rightmost position and all $\mathrm{R}^{+}_{\mathfrak{l}_{\tilde{\alpha}}}$ are on the left.
  • the root $\tilde{\alpha}$ is next to the right after $\mathrm{R}^{+}_{\mathfrak{l}_{\tilde{\alpha}}}$.
Proof.
In the non-graded case this is derived from the properties of the longest element of the Weyl group. In our case this is checked by a direct examination. ∎
We can conclude now, by Lemma 5.1, that for each $\alpha\in\bar{\Pi}_{\mathfrak{l}}$ the elements $F_{\tilde{\alpha}}$ and $F_{\mu}$ with $\mu\in\Pi_{\mathfrak{l}_{\tilde{\alpha}}}$ form a left coideal subalgebra in $U_{q}(\mathfrak{g})$.
Theorem 5.3.
The subalgebra $U_{q}(\mathfrak{k})\subset U_{q}(\mathfrak{g})$ generated by $e_{\alpha},f_{\alpha},q^{\pm h_{\alpha}}$ with $\alpha\in\Pi_{\mathfrak{l}}$, and by $X_{\alpha}=q^{h_{\tilde{\alpha}}-h_{\alpha}}e_{\alpha}+c_{\alpha}F_{\tilde{\alpha}}+\grave{c}_{\alpha}(q^{u_{\alpha}}-1)$, $q^{\pm(h_{\tilde{\alpha}}-h_{\alpha})}$ with $\alpha\in\bar{\Pi}_{\mathfrak{l}}$, $u_{\alpha}\in\mathfrak{c}$, is a left coideal in $U_{q}(\mathfrak{g})$.
Proof.
First set $\grave{c}_{\alpha}=0$. For each $\alpha\in\bar{\Pi}_{\mathfrak{l}}$, we find, using Lemma 5.1:
$\Delta(F_{\tilde{\alpha}})\in(q^{h_{\tilde{\alpha}}}\otimes F_{\tilde{\alpha}}+F_{\tilde{\alpha}}\otimes 1)+U_{q}(\mathfrak{g})\otimes U_{q}(\mathfrak{l}_{\tilde{\alpha}}),$
$\Delta(q^{h_{\tilde{\alpha}}-h_{\alpha}}e_{\alpha})=q^{h_{\tilde{\alpha}}}\otimes q^{h_{\tilde{\alpha}}-h_{\alpha}}e_{\alpha}+q^{h_{\tilde{\alpha}}-h_{\alpha}}e_{\alpha}\otimes q^{h_{\tilde{\alpha}}-h_{\alpha}}.$
Adding the lines together we arrive at
$\Delta(X_{\tilde{\alpha}})\in q^{h_{\tilde{\alpha}}}\otimes X_{\alpha}+F_{\tilde{\alpha}}\otimes 1+q^{h_{\tilde{\alpha}}-h_{\alpha}}e_{\alpha}\otimes q^{h_{\alpha}-h_{\tilde{\alpha}}}+U_{q}(\mathfrak{g})\otimes U_{q}(\mathfrak{l}_{\tilde{\alpha}}),$
as required. It is straightforward to see that the coproduct of the term with $\grave{c}_{\alpha}\neq 0$ is in $U_{q}(\mathfrak{g})\otimes U_{q}(\mathfrak{k})$ too. This completes the proof. ∎
In the next sections we relate the RE matrices presented in Theorem 3.1 with coideal subalgebras. We require that such a matrix commutes with $\pi\bigl{(}U_{q}(\mathfrak{l})\bigr{)}$ and all $\pi(q^{h_{\tilde{\alpha}}-h_{\alpha}})$, where $\alpha\in\bar{\Pi}_{\mathfrak{l}}$. Th e values of $c_{\alpha}$ and $\grave{c}_{\alpha}$ are determined from this requirement. The constant $\grave{c}_{\alpha}$ is distinct from zero only if $\mathfrak{g}=\mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$ with $\alpha=\tilde{\alpha}_{m}=\alpha_{m},\ m=\frac{N+2\mathbf{m}}{2}$, and $\mu\neq-\lambda$. Expressions for $F_{\tilde{\alpha}}$ with non-simple $\tilde{\alpha}$ will be provided explicitly in terms of q-commutator defined as
$[F_{\alpha},F_{\beta}]_{q^{\pm i}}=F_{\alpha}F_{\beta}-(-1)^{(|F_{\alpha}||F_{\beta}|)}q^{\mp\lceil\frac{i}{2}\rceil(\alpha,\beta)}F_{\beta}F_{\alpha},\quad\forall\alpha,\beta\in\mathrm{R}^{+}.$
Here $i=0,1,2$ and $i\mapsto\lceil\frac{i}{2}\rceil$ sends it to $0,1,1$.
5.1 K-matrices of type $A$
It is found that $\tilde{\alpha}_{i}=\alpha_{i}$ for all $i<m$, and $c_{\alpha_{i}}=-\frac{y_{i+1}}{qy_{i}}$, for ortho-symplectic $\mathfrak{g}$. The four types of $\mathfrak{g}$ will be further treated separately.
$\bullet\quad\mathfrak{g}=\mathfrak{g}\mathfrak{l}(N|2\mathbf{m})$
The Satake diagrams fall into two classes:
[…]
with the mixture parameters
$c_{\alpha_{i}}=(-1)^{\delta_{i}^{\mathbf{m}^{\prime}}}\frac{y_{i+1}}{y_{i}},\quad\mathrm{if}\quad i<m\quad\mathrm{or}\quad i>N+2\mathbf{m}-m.$
The roots $\tilde{\alpha}_{m}$ and $\tilde{\alpha}_{N+2\mathbf{m}-m}$ are not simple when $m<\frac{N+2\mathbf{m}}{2}$. The corresponding root vectors can be defined as
$F_{\tilde{\alpha}_{m}}=[\dots[F_{m+1},F_{m+2}]_{q},\dots F_{n-m+1}]_{q},\quad F_{\tilde{\alpha}_{n-m+1}}=[\dots[F_{m},F_{m+1}]_{\bar{q}},\dots F_{n-m}]_{\bar{q}}.$
The mixture parameters are expressed by
$c_{\alpha_{m}}=\frac{(-1)^{N}\mu}{y_{m}},\quad c_{\alpha_{N+2\mathbf{m}-m}}=\begin{cases}\frac{(-1)^{N+1+\delta_{m}^{\mathbf{m}}}q^{2(N+2\mathbf{m})-4m-3}\lambda}{y_{m}},&\mbox{if }\ \mathbf{m}\leq m,\\ \frac{(-1)^{N+1}q^{2(N-2\mathbf{m})+4m+3}\lambda}{y_{m}},&\mbox{if }\ m<\mathbf{m}.\end{cases}$
The case $m=\frac{N+2\mathbf{m}}{2}$ with $\mathbf{m}<m$ is described by the diagram on the right. The simple root $\alpha_{m}=\alpha$ corresponds to the following mixed vector
$X_{\alpha}=e_{\alpha}+c_{\alpha}F_{\alpha}+\grave{c}_{\alpha}(q^{h_{\alpha}}-1),$
where
$c_{\alpha}=\frac{-\lambda\mu q}{y_{m}^{2}},\quad\grave{c}_{\alpha}=\frac{(\mu+\lambda)q}{(q^{2}-1)y_{m}}.$
Remark that $\mathfrak{l}=\{0\}$, and $\mathfrak{c}=\mathfrak{h}$ in this case. The element $u_{\alpha}\in\mathfrak{c}$ is taken equal to $h_{\alpha}$.
$\bullet\quad\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m})$
There are two Satake diagrams leading the K-matries of type $A$:
Note that the diagram on the right is admissible. It does not fall under conditions of Lemma 4.19 because the odd root $\alpha_{m}$ is not grey. The composite root vectors for the left diagram are
$F_{\tilde{\alpha}_{m}}=[\dots[[[\dots[[F_{m},F_{m+1}]_{q},F_{m+2}]_{q},\dots F_{n}]_{q},F_{n}],F_{n-1}]_{q},\dots F_{m+1}]_{q},\ m<n-1,$
$F_{\tilde{\alpha}_{m}}=[[F_{n-1},F_{n}]_{q},F_{n}],\quad m=n-1,$
with mixture coefficient
$c_{\alpha_{m}}=\frac{(-1)^{(n-m-\delta_{m}^{\mathbf{m}})}\bar{q}^{2\delta_{m}^{\mathbf{m}}}\lambda}{y_{m}}.$
In the scenario with $\mathbf{m}=m=n$ relative to the second diagram, the mixture constant is
$c_{\alpha_{m}}=-\frac{\lambda}{y_{m}q}.$
$\bullet\quad\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m})$
The case $m\leqslant n-2$ includes two diagrams
$\ldots$ $\ldots$ $\alpha_{1}$$\alpha_{m}$$\alpha_{n-1}$$\alpha_{n}$$\mathbf{n}\neq 1$
$\ldots$ $\ldots$   $\alpha_{1}$$\alpha_{m}$$\alpha_{n-1}$$\alpha_{n}$$\mathbf{n}=1$
The root vectors are
$F_{\tilde{\alpha}_{m}}=[\dots[[\dots[[F_{m},F_{m+1}]_{q},F_{m+2}]_{q},\dots F_{n}]_{q},F_{n-2}]_{q},\dots F_{m+1}]_{q},\quad\mbox{for}\quad m<n-2,$
$F_{\tilde{\alpha}_{m}}=[[F_{n-2},F_{n-1}]_{q},F_{n}]_{q}\quad\mbox{for}\quad m=n-2.$
In both cases, the parameters are
$c_{\alpha_{m}}=\frac{(-1)^{(n-m+1+\delta_{m}^{\mathbf{m}})}\bar{q}^{2\delta_{m}^{\mathbf{m}}}\lambda}{y_{m}}.$
The case $m=n-1$ corresponds to the diagram
[…]
$\bullet\quad\mathfrak{g}=\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m})$
The graded Satake diagram is
$\ldots$ $\ldots$ $<$$\alpha_{1}$$\alpha_{m}$$\alpha_{n}$
with coefficients
$c_{\alpha_{m}}=(-1)^{n-m+\delta_{m}^{\mathbf{m}}}\frac{\bar{q}^{2\delta_{m}^{\mathbf{m}}}\lambda}{y_{m}},$
and root vectors
$F_{\tilde{\alpha}_{m}}=[\dots[F_{m},F_{m+1}]_{q},F_{m+2}]_{q},\dots F_{n}]_{q^{2}},F_{n-1}]_{q},\dots F_{m+1}]_{q},\quad m<n-1,$
$F_{\tilde{\alpha}_{m}}=[F_{n-1},F_{n}]_{q^{2}},\quad m=n-1.$
This completes the description of the diagrams for this type of K-matrix.
5.2 K-matrices of type $B$
For all cases, we find
$F_{\tilde{\alpha}_{2i}}=[[F_{2i},F_{2i+1}]_{q},F_{2i-1}]_{q}\ \mathrm{with}\ i=1,\ldots,\frac{m}{2}-1,\quad\mbox{and }$
[…]
$\bullet\quad\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}+1|2\mathbf{m})$
The Satake diagram is
$\ldots$ $\ldots$ $>$ $\alpha_{1}$$\alpha_{m}$$\alpha_{n}$
For even $m=2i$, the non-simple root vectors are
$F_{\tilde{\alpha}_{m}}=[\dots[[[\dots[[F_{m},F_{m+1}]_{q},F_{m+2}]_{q},\dots F_{n}]_{q},F_{n}],F_{n-1}]_{q},\dots F_{m+1}]_{q},F_{m-1}]_{q},\quad m<n,$
with
$c_{\alpha_{m}}=(-1)^{n+\delta_{m}^{\mathbf{m}}}\frac{q^{-2\delta_{m}^{\mathbf{m}}-1}\lambda}{z_{m-1}},$
and
$F_{\tilde{\alpha}_{m}}=[F_{n-1},F_{n}]_{q^{-1}},\quad m=n,$
with
$c_{\alpha_{m}}=\frac{\lambda}{z_{n-1}q^{3}}.$
$\bullet\quad\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2\mathbf{n}|2\mathbf{m})$
For $m<n-1$, and $\mathbf{n}\neq 1$, we have the diagram
$\ldots$  $\alpha_{1}$$\alpha_{m}$$\alpha_{n-1}$$\alpha_{n}$$\ldots$
When $m<n-1$, and $\mathbf{n}=1$, there is another diagram
$\ldots$    $\alpha_{1}$$\alpha_{m}$$\alpha_{n-1}$$\alpha_{n}$$\ldots$
The mixture parameter is
$c_{\alpha_{m}}=(-1)^{n+1+\delta_{m}^{\mathbf{m}}}\frac{q^{-2\delta_{m}^{\mathbf{m}}-1}\lambda}{z_{m-1}},\quad m<n-1$
and root vector
$F_{\tilde{\alpha}_{m}}=[[\dots[[\dots[[F_{m},F_{m+1}]_{q},F_{m+2}]_{q},\dots F_{n}]_{q},F_{n-2}]_{q},\dots F_{m+1}]_{q},F_{m-1}]_{q}.$
In the case of $m=n-1$, it corresponds to the diagram
[…]
Composite root vectors are defined as
$F_{\tilde{\alpha}_{n-1}}=[F_{n},F_{n-2}]_{q},\quad F_{\tilde{\alpha}_{n}}=[F_{n-1},F_{n-2}]_{q},\quad\mbox{and}\quad c_{\alpha_{n-1}}=c_{\alpha_{n}}=-\frac{\lambda}{q^{3}z_{n-1}}.$
$\bullet\quad\mathfrak{g}=\mathfrak{s}\mathfrak{p}\mathfrak{o}(2\mathbf{n}|2\mathbf{m})$
The diagram is
$\ldots$ $\ldots$ $<$$\alpha_{1}$$\alpha_{m}$$\alpha_{n}$
Depending on the value of $m$ we define the root vectors as
$F_{\tilde{\alpha}_{m}}=[[\dots[[\dots[[F_{m},F_{m+1}]_{q},F_{m+2}]_{q},\dots F_{n}]_{q^{2}},F_{n-1}]_{q},\dots F_{m+1}]_{q},F_{m-1}]_{q},\quad\mathrm{if}\ m<n-1,$
$F_{\tilde{\alpha}_{m}}=[[F_{n-1},F_{n}]_{q^{2}},F_{n-2}]_{q},\quad\mathrm{if}\ m=n-1\in 2\mathbb{Z},$
with the mixture parameter
$c_{\alpha_{m}}=(-1)^{n+\delta_{m}^{\mathbf{m}}}\frac{q^{-2\delta_{m}^{\mathbf{m}}-1}\lambda}{z_{m-1}},\quad m\leq\mathbf{m}.$
This completes our description of type I coideal subalgebras and their K-matrices.
5.3 K-matrix of type $C$
The matrix $C$ is related with the algebra $\mathfrak{g}=\mathfrak{o}\mathfrak{s}\mathfrak{p}(2|4\mathbf{m})$ and its Satake diagram
[…]
The composite root vectors of roots $\tilde{\alpha}$, where $\alpha\in\bar{\Pi}_{\mathfrak{l}}$, are
$F_{\tilde{\alpha}_{2i}}=[[F_{2i},F_{2i+1}]_{q},F_{2i-1}]_{q}\ \mathrm{with}\ c_{\alpha_{2i}}=-\frac{x_{2i+1}}{qx_{2i-1}},\ \mathrm{where}\ i=1,\ldots,\frac{n-3}{2},$
for even $\alpha$. Also, there are two odd root vectors
$F_{\tilde{\alpha}_{n}}=[F_{n},F_{n-2}]_{q},\quad F_{\tilde{\alpha}_{n-1}}=[F_{n-1},F_{n-2}]_{q},$
with mixture parameters
$c_{\alpha_{n}}=-\frac{x_{n+2}}{x_{n}q},\quad c_{\alpha_{n-1}}=-\frac{x_{n}}{x_{n-2}q}.$
Acknowledgement
This work is done at the Center of Pure Mathematics MIPT. It is financially supported by Russian Science Foundation grant 23-21-00282.
The first author (D. Algethami) is thankful to the Deanship of Graduate Studies and Scientific Research at University of Bisha for the financial support through the Scholars Research Support Program of the University.
A. Mudrov and V. Stukopin are grateful to V. Zhgoon for valuable discussions of spherical manifolds.
The authors are profoundly indebted to the anonymous reviewer for careful reading of the manuscript and for extremely valuable comments that served to improvement of its first version.
Data Availability.
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Declarations
Competing interests.
The authors have no competing interests to declare that are relevant to the content of this article.